### v1

# Time-optimal problem in the space of probabilities measures

Time-optimal problem in the space of probabilities measures
Yurii Averboukh HSE University, Moscow, Russia
& Krasovskii Institute of Mathematics and Mechanics,
Yekaterinburg, Russia
averboukh@gmail.com
   Ekaterina Kolpakova Krasovskii Institute of Mathematics and Mechanics,
& Yekaterinburg, Russia
eakolpakova@gmail.com
Аннотация
This paper focuses on the value function in the time-optimal problem for a continuity equation in the space of probability measures. We derive the dynamic programming principle for this problem. In particular, we prove that the Kruzhkov transform of the value function coincides with a unique discontinuous viscosity solution to the corresponding Dirichlet problem for the Hamilton–Jacobi equation. Finally, we establish the $\Gamma$-convergence of the value function in a perturbed problem to the value function in the unperturbed problem.
MSC Classification (2020): 49J20, 49J52, 49L25, 49N70, 91A23, 49K20, 82C22.
keywords:
time-optimal problem, Hamilton—Jacobi equation, controlled continuity equation, viscosity solution, Dirichlet problem, nonsmooth analysis in the Wasserstein space.
1 Introduction
In this paper, we study a control system in the space of probability measures, where the dynamics are governed by a nonlocal continuity equation driven by an external force. This equation models the evolution of a system comprising infinitely many identical particles, with each particle influenced by the collective field generated by all other particles.
The nonlocal continuity equation traces its origins to the pioneering work of Vlasov on plasma dynamics. Nowadays, it finds widespread application in various fields, including opinion dynamics models [16, 32, 38], crowd behavior analysis [21, 23, 26, 34] and swarm dynamics [24, 31].
The objective of the control is to steer the system to a target in minimal time. The target here is assumed to be an arbitrary closed set of probability measures. In this work, we develop the dynamic programming approach and investigate the properties of the value function.
The main results of the paper are as follows. First, we establish the existence of an optimal relaxed control. Next, we derive the dynamic programming principle for the value function in this setting. We then implement a framework based on the Dirichlet problem for the Hamilton-Jacobi equation, which, in this context, is a partial differential equation in the space of measures. We demonstrate that the Kruzhkov transform of the value function satisfies the aforementioned Dirichlet problem in the viscosity sense. Furthermore, by deriving a comparison principle, we prove the existence and uniqueness of a solution to the Dirichlet problem associated with the time-optimal problem for a nonlocal continuity equation. In general, the value function is discontinuous; however, we identify a sufficient condition that ensures its continuity. Finally, we show that the values of perturbed time-optimal problems $\Gamma$-converge to the solution of the limiting problem.
Now, let us give a brief overview of the existing literature and compare our results with what is already known. The study of control problems for nonlocal continuity began with papers [13, 36]. Necessary optimality conditions in the form of the Pontryagin maximum principle for finite-horizon problems were derived in [8, 9, 10, 11, 35]. We should also highlight the paper [5], where the authors develop the dynamic programming principle for the optimal control problem of the continuity equation on a finite time interval. In that work, it is shown that the value function restricted to the space of compactly supported measures is a solution to the Cauchy problem for the Hamilton-Jacobi equation in the space of probability measures. Moreover, the uniqueness of the solution is also established on the set of compactly supported probability measures.
Notice that the problem studied in this paper assumes the system is influenced by an external force. Consequently, the velocity field is a continuous function of a point. An alternative approach considers the possibility of a discontinuous velocity field. This framework known as the mean field type control theory is particularly relevant for systems involving intellectual agents striving to achieve a common goal. For the current state of research in this direction, we refer to the works of [3, 27, 28]. Time-optimal problems for mean field type control systems were investigated in [19, 17, 18], where the target is assumed to be a hyperplane in the space of probability measures. In those papers, the value function was characterized as a solution to a Dirichlet problem for the Hamilton-Jacobi equation in the space of measures. Furthermore, sufficient conditions for the regularity of the value function were established. However, these works do not provide a uniqueness result.
As previously mentioned, our approach to analyzing time-optimal problems relies on viscosity solutions to the Hamilton–Jacobi equation. This methodology was initially developed for finite-dimensional equations (see [6, 22, 37]).
Nowadays, the concept of viscosity solutions has been extended to spaces of probability measures (see [1, 5, 19, 18, 25, 27, 33]). It is worth to notice that the definition of a viscosity solution hinges on the concepts of subdifferentials and superdifferentials. In the context of probability measure spaces, there are multiple approaches to defining these objects. The central question revolves around determining which variations are admissible and what constitutes sub- and superdifferentials. In our framework, we assume that elements of the sub- and superdifferentials are square-integrable functions, and we permit variations defined by directions that are also square-integrable functions. This approach was first introduced in [15]. Alternatively, one can consider varying a measure near an element specified by a given direction. This method was employed in [5] for finite-horizon optimal control problems involving nonlocal continuity equations and their corresponding Hamilton-Jacobi equations. There are also approaches that rely on directions defined by plans (see [1, 19, 18, 25, 27]). Moreover, in [2] elements of subdifferentials are assumed to be distributions. It is important to note that the aforementioned approaches are intrinsic, as they utilize only measures and directions within the space of measures. However, a probability measure with finite second moment can also be represented as a push-forward measure via some square-integrable random variable. This leads to an external approach to nonsmooth analysis in the space of probability measures (see [14, 25, 33]). Connections between the external approach to nonsmooth analysis and certain definitions of sub- and superdifferentials within the intrinsic framework are explored in [14, 25, 28]. Additionally, the papers [27, 28] offer a comparative analysis of various intrinsic approaches.
In our view, the approach to nonsmooth analysis employed in this paper is particularly well-suited for the control problem associated with the nonlocal continuity equation. However, in Section 9, we discuss an alternative approach that involves varying a measure near an element specified by a given direction, as introduced in [5], alongside definitions based on directions provided by plans (see [19, 18, 27]). We show that these approaches yield the same notion of a viscosity solution to the Dirichlet problem when the Hamiltonian arises from the control problem for the nonlocal continuity equation.
The paper is organized as follows. In Section 2, we introduce the general notation used throughout the paper. The formulation of the time-optimal control problem for the nonlocal continuity equation is presented in Section 3. In this section, we also prove the existence of an optimal control and show that the value function is lower semicontinuous. Section 4 introduces the statement of the Dirichlet problem and defines the notion of a viscosity solution corresponding to the examined problem. The comparison principle for this problem is established in Section 5, which, in turn, implies the uniqueness of the viscosity solution. In Section 6, we prove that the Kruzhkov transform of the value function is a viscosity solution to the aforementioned Dirichlet problem. Additionally, we derive a sufficient condition ensuring the continuity of the value function and show that the values of perturbed time-optimal control problems for the nonlocal continuity equation $\Gamma$-converge to the original problem (see Section 7). An example of a time-optimal control problem for the controlled continuity equation is provided in Section 8. Finally, in Section 9, we discuss equivalent definitions of the viscosity solution in the context of the examined case.
2 Preliminaries
If $X_{1},\ldots,X_{n}$ are sets, $i_{1}\ldots,i_{k}\in\{1,\ldots,n\}$, then $\operatorname{p}^{i_{1},\ldots,i_{k}}$ denotes a natural projection of $X_{1}\times\ldots\times X_{n}$ onto $X_{i_{1}}\times\ldots\times X_{i_{k}}$.
If $(\Omega^{\prime},\mathcal{F}^{\prime})$ and $(\Omega^{\prime\prime},\mathcal{F}^{\prime\prime})$ are measurable spaces, $m$ is a measure on $\mathcal{F}^{\prime}$, a map $h:\Omega^{\prime}\rightarrow\Omega^{\prime\prime}$ is $\mathcal{F}^{\prime}/\mathcal{F}^{\prime\prime}$-measurable, then we denote by $h\sharp m$ the push-forward measure [2] defined by the rule: for each $\Gamma\in\mathcal{F}^{\prime\prime}$
$(h\sharp m)(\Gamma)=m(h^{-1}(\Gamma)).$
If $(X,\rho_{X})$ is a Polish space, $x\in X$, $r>0$, then $\mathbb{B}_{r}(x)$ stands for the ball of radius $r$ centered at $x$.
If $(X,\rho)$ is a Polish space, then $\mathcal{M}(X)$ denotes the space of finite Borel measures on $X$, while $\mathcal{P}(X)$ stands for the space of Borel probabilities on $X$. Recall that $\mathcal{P}(X)=\{m\in\mathcal{M}(X):\,m(X)=1\}$.
We consider on $\mathcal{M}(X)$ the narrow convergence. This means that $\{m_{n}\}_{n=1}^{\infty}\subset\mathcal{M}(X)$ converges to $m\in\mathcal{M}(X)$ if, for every $\phi\in C_{b}(X)$,
$\int_{X}\phi(x)m_{n}(dx)\rightarrow\int_{X}\phi(x)m(dx)\text{ as }n\rightarrow\infty.$
Obviously, $\mathcal{P}(X)$ is closed w.r.t. the narrow convergence.
We denote the space of probability measures with the finite second moment by $\mathcal{P}^{2}(X)$, i.e., $m\in\mathcal{P}(X)$ belongs to $\mathcal{P}^{2}(X)$ provided that, for some $x_{*}\in X$, $\int_{X}\rho^{2}(x,x_{*})m(dx)<\infty$. If $X$ is a Banach space, then we denote
$\varsigma(m)\triangleq\Bigg{[}\int_{X}\|x\|^{2}m(dx)\Bigg{]}^{1/2}.$
We endow $\mathcal{P}^{2}(X)$ with the second Wasserstein distance defined by the rule:
$W_{2}(m_{1},m_{2})=\Bigg{[}\inf\limits_{\pi\in\Pi(m_{1},m_{2})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\rho(x_{1},x_{2})^{2}\pi(d(x_{1},x_{2}))\Bigg{]}^{1/2},$
where $\Pi(m_{1},m_{2})$ stands for the set of probabilities $\pi\in\mathcal{P}^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ such that $\operatorname{p}^{i}\sharp\pi=m_{i}$.
If $(X,\rho)$ is a Polish space, $Y$ is a Banach space, $m$ is a measure on $X$, then $L^{2}(X,m;Y)$ denotes the space of functions $\varphi:X\mapsto Y$ such that
$\int_{X}\|\varphi(x)\|^{2}_{Y}m(dx)<\infty.$
If $X=Y=\mathbb{R}^{d}$, we will shorten the notation and write $L^{2}(m)$ instead of $L^{2}(\mathbb{R}^{d},m;\mathbb{R}^{d})$. Given $s_{1},s_{2}\in L^{2}(m)$, the inner product of these functions in $L^{2}(m)$ is denoted by $\langle s_{1},s_{2}\rangle_{m}$, i.e.,
$\langle s_{1},s_{2}\rangle_{m}\triangleq\int_{\mathbb{R}^{d}}\langle s_{1}(x),s_{2}(x)\rangle m(dx).$
Additionally, $L^{0}(X;Y)$ denotes the set of all measurable functions from $X$ to $Y$.
If $(X,\rho_{X})$, $(Y,\rho_{Y})$ are Polish spaces, $m$ is a finite measure on $X$, then we denote by $\Lambda(X,m;Y)$ the set of measures $\alpha$ on $X\times Y$ such that $\operatorname{p}^{1}\sharp\alpha=m$. Notice that due to the disintegration theorem [7, Theorem 10.4.14], given $\alpha\in\Lambda(X,m;Y)$ one can find a family of probability measures $\{\alpha(\cdot|x)\}_{x\in X}\subset\mathcal{P}(Y)$ such that, for every function $\phi\in C_{b}(X\times Y)$,
$\int_{X\times Y}\phi(x,y)\alpha(d(x,y))=\int_{X}\int_{Y}\phi(x,y)\alpha(dy|x)m(dx).$ (1)
Moreover, this family is $m$-a.e. unique. If $\{\alpha(\cdot|x)\}_{x\in X}$ is a family of probabilities on $X$ that is weakly measurable, one can uniquely construct a measure $\alpha\in\Lambda(X,m;Y)$ such that (1) holds true.
Notice that each Borel function $h:X\rightarrow Y$ produces a weakly measurable family of probabilities by the rule
$\alpha_{h}(\cdot|x)\triangleq\delta_{h(x)}.$ (2)
Hereinafter, $\delta_{z}$ stands for the Dirac measure concentrated at $z$. So, (2) produces the embedding of the space of $L^{1}(X,m;Y)$ into $\Lambda(X,m;Y)$. Notice that, if $Y$ is a compact, then the set $\{\alpha_{h}:\,h\in L^{0}(X;Y)\}$ is dense in $\Lambda(X,m;Y)$.
In the following, given a metric compact $U$, we will consider measures on $[0,+\infty)\times U$ those marginal distributions on $[0,+\infty)$ are equal to the Lebesgue measure on $[0,+\infty)$. The set of such measures is denoted by $\mathcal{U}$. Elements of $\mathcal{U}$ are regarded as a relaxed (generalized) controls. If $\xi\in\mathcal{U}$, $T>0$, then $\xi|_{T}$ stands for its restriction on $[0,T]\times U$. Notice that $\xi|_{T}$ lies in $\mathcal{U}_{T}\triangleq\Lambda([0,T],\lambda;U)$, where $\lambda$ denotes the standard Lebesgue measures.
Definition 2.1.
We say that a sequence $\{\xi_{n}\}_{n=1}^{\infty}\subset\mathcal{U}$ converges to $\xi\in\mathcal{U}$ provided that, for each $T>0$, $\{\xi_{n}|_{T}\}_{n=1}^{\infty}$ converges to $\{\xi|_{T}\}$ narrowly.
Notice that in this definition it suffices to consider only natural $T$. Thus, the space $\mathcal{U}$ is compact. Moreover, each measurable function $u:[0,+\infty)\rightarrow U$ generates an element of $\mathcal{U}$ such that $\xi(\cdot|t)\triangleq\delta_{u(t)}$. The set of such elements is dense in $\mathcal{U}$.
3 Statement of the problem
The main object of the paper is the following controlled nonlocal continuity equation
$\partial_{t}m(t)+\operatorname{div}(f(x,m(t),u(t))m(t))=0,\ \ t\in[0,+\infty).$ (3)
Here, $x\in\mathbb{R}^{d}$, $m(t)\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $u(t)\in U$, the set $U$ is a metric compact set. We interpret $u(t)$ is an instantaneous control. Thus, $U$ is the space of admissible controls.
Equation (3) describes an evolution of the distribution of infinitely many similar particles obeying the following equation
$\frac{d}{dt}x(t)=f(x(t),m(t),u(t)),\ x\in\mathbb{R}^{d},\ u\in U.$
Here, the distribution of agents at the instant $t$ is denoted $m(t)\in\mathcal{P}^{2}(\mathbb{R}^{d})$, while $u(t)$ stands for an external control that influences the whole system.
We put the following assumptions.
Hypothesis 3.1.
The function $f$ is continuous in all arguments.
Hypothesis 3.2.
There exists a constant $L>0$ such that, for every $x_{1},x_{2}\in\mathbb{R}^{d}$, $m_{1},m_{2}\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $u\in U$, one has the inequality
$\|f(x_{1},m_{1},u)-f(x_{2},m_{2},u)\|\leq L(\|x_{1}-x_{2}\|+W_{2}(m_{1},m_{2})).$
Within the paper, we use relaxed controls. As we mentioned above, each relaxed control is a measure on $[0,+\infty)\times U$ with the marginal distribution on $[0,+\infty)$ coinciding with the Lebesgue measure. Furthermore, given $\xi\in\mathcal{U}$, there exists its disintegration w.r.t. the Lebesgue measure denoted by $\xi(\cdot|t)$. Formally, the usage of relaxed controls means that we replace continuity equation (3) with
$\partial_{t}m(t)+\operatorname{div}\Bigg{(}\int_{U}f(x,m(t),u)\xi(du|t)\cdot m(t)\Bigg{)}=0,\ \ t\in[0,+\infty).$ (4)
We will consider this equation in the distributional sense.
Definition 3.3.
Let $T>0$, $\xi\in\mathcal{U}_{T}$. A measure-valued function $m(\cdot):[0,T]\rightarrow\mathcal{P}^{2}(\mathbb{R}^{d})$ is called a solution of equation (4) on $[0,T]$ if, for every $\varphi\in C_{0}^{\infty}((0,T);\mathbb{R}^{d})$, one has that
$\int_{[0,T]\times U}\int_{\mathbb{R}^{d}}\partial_{t}\varphi(t,x)+\left\langle\nabla\varphi(t,x),f(x,m(t),u)\right\rangle m(t,dx)\xi(d(t,u))=0.$
If $\xi\in\mathcal{U}$, then a measure-valued function $m(\cdot):[0,+\infty)\rightarrow\mathcal{P}^{2}(\mathbb{R}^{d})$ is a solution of (4) on $[0,+\infty)$ if, for every $T>0$, the restriction of $m(\cdot)$ on $[0,T]$ is a solution of (4) on $[0,T]$.
Given $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $\xi\in\mathcal{U}$, we denote by $m(\cdot;\mu,\xi)$ the unique solution of (4) on $[0,+\infty)$ satisfying $m(0)=\mu$. If $\xi\in\mathcal{U}_{T}$, then without loss of generality, we will keep the designation for the solution of (4) satisfying the initial condition $m(\cdot)=\mu$.
An alternative definition of the solution to continuity equation (4) relies on the particle interpretation. To introduce it, for given $y\in\mathbb{R}^{d}$, a continuous function $m(\cdot):[0,T]\rightarrow\mathcal{P}^{2}(\mathbb{R}^{d})$, $\xi\in\mathcal{U}_{T}$, we denote by $X(\cdot;y,m(\cdot),\xi)$ a solution of the initial value problem
$\frac{d}{dt}x(t)=\int_{U}f(x(t),m(t),u)\xi(du|t),\ x(0)=y.$
The function $X(\cdot;y,m(\cdot),\xi)$ describes the motion of a particle on the time interval $[0,T]$ in the case where the distribution of all particles is given by $m(\cdot)$, while the system is affected by a relaxed control $\xi$. The motion on $[0,+\infty)$ can be defined in the same way.
Furthermore, for $T>0$, $t\in[0,T]$, the evaluation operator $e_{t}$ assigns to $x(\cdot)\in C([0,T];\mathbb{R}^{d})$ the value $x(t)$.
Proposition 3.4.
Given $T>0$, $\xi\in\mathcal{U}_{T}$, a measure-valued function $m(\cdot):[0,T]\rightarrow\mathcal{P}^{2}(\mathbb{R}^{d})$ solves (4) if and only if there exists a probability $\chi\in\mathcal{P}(C([0,T];\mathbb{R}^{d}))$ such that the following properties holds true:
  1. (M1)
    $m(t)\triangleq e_{t}\sharp\chi$;
  2. (M2)
    $\chi$-a.e. curves $x(\cdot)$ are such that
    $x(\cdot)=X(\cdot;x(0),m(\cdot),\xi).$
This statement directly follows from the seminal superposition principle (see [2, Theorem 8.2.1]. Furthermore, we have the following equivalent definition of the solution of (4) on $[0,+\infty)$.
Corollary 3.5.
Assume that $\xi\in\mathcal{U}$. Then, a measure-valued function $m(\cdot):[0,T]\rightarrow\mathcal{P}^{2}(\mathbb{R}^{d})$ solves (4) if and only if there exists a probability $\chi\in\mathcal{P}(C([0,+\infty);\mathbb{R}^{d}))$ such that conditions (M1)(M2) hold true.
Now, let us describe the time-optimal problem examined in the paper. We assume that we are given with a closed set $M\subset\mathcal{P}^{2}(\mathbb{R}^{d})$. Furthermore, we denote $G\triangleq\mathcal{P}^{2}(\mathbb{R}^{d})\setminus M$. Additionally, $\partial G$ stands for the border of $G$. Let us consider the functional $\tau$ defined on $C([0,+\infty);\mathcal{P}^{2}(\mathbb{R}^{d}))$ such that
$\tau(m(\cdot))\triangleq\inf\{t\in[0;+\infty):\ m(t)\in M\}.$ (5)
Notice that, if, for every $t>0$, $m(t)\not\in M$, then $\tau(m(\cdot))=+\infty$. The time-optimal problem means that, given an initial distribution $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, one wishes to minimize the quantity $\tau(m(\cdot;\mu,\xi))$. Thus, the value function of time-optimal problem (3), (5) is defined by the rule:
$\operatorname{Val}(\mu)=\inf\{\tau(m(\cdot;\mu,\xi)):\ \ \xi\in\mathcal{U}\}.$ (6)
3.1 Properties of the value function
The properties of the value function in a time-optimal problem in $\mathbb{R}^{d}$ are well studied [6]. It is proved that the value function is lower semicontinuous and satisfies the dynamical programming principle. In this section, we derive the analogous properties for the examined time–optimal control in the space of probability measures.
First, let us recall the continuity of the trajectories on each finite interval proved in [4].
Proposition 3.6.
Assume that, for each $n$, $f_{n}$ is a function from $\mathbb{R}^{d}\times\mathcal{P}^{2}(\mathbb{R}^{d})\times U$ to $\mathbb{R}^{d}$; $\mu_{n}\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $\xi_{n}\in\mathcal{U}$ are such that
  • each function $f_{n}$ is continuous w.r.t. all variables and Lipschitz continuous w.r.t. $x$ and $m$ for the constant $L$;
  • there exists a continuous function $f:\mathbb{R}^{d}\times\mathcal{P}^{2}(\mathbb{R}^{d})\times U\rightarrow\mathbb{R}^{d}$ that is Lipschitz continuous w.r.t. $x$ and $m$ such that, for each $c>0$,
    $\begin{split}\sup\Big{\{}\|f_{n}(x,&m,u)-f(x,m,u)\|:\\ &x\in\mathbb{R}^{d},\,m\in\mathcal{P}^{2}(\mathbb{R}^{d}),\,\varsigma(m)\leq c,\,u\in U\Big{\}}\rightarrow 0\text{ as }n\rightarrow\infty;\end{split}$
  • the sequences $\{\mu_{n}\}_{n=1}^{\infty}$ and $\{\xi_{n}\}_{n=1}^{\infty}$ converge to $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$ and $\xi\in\mathcal{U}$ respectively;
  • for each natural $n$, $m_{n}(\cdot)$ is the solution of (4) for the dynamics equal to $f_{n}$, relaxed control $\xi_{n}$ such that $m_{n}(0)=\mu_{n}$.
Then, for every $T>0$, the sequence of $\{m_{n}(\cdot)\}_{n=1}^{\infty}$ converges to $m(\cdot;\mu,\xi)$ in $C([0,T];\mathcal{P}^{2}(\mathbb{R}^{d}))$.
Remark 3.7.
This proposition and the fact that the set of relaxed controls corresponding to usual ones is dense in $\mathcal{U}$ imply that, given $\xi\in\mathcal{U}$ and $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, there exists a sequence $\{m_{n}\}_{n=1}^{\infty}$ such that, for each $n$, $m_{n}(\cdot)$ satisfies
$\partial_{t}m_{n}(t)+\operatorname{div}(f(x,m_{n}(t),u_{n}(t))m_{n}(t))=0,\ \ m_{n}(0)=\mu$
whereas $m_{n}(\cdot)\rightarrow m(\cdot;\mu,\xi)$ as $n\rightarrow\infty$. In particular,
$\begin{split}\operatorname{Val}(\mu)=\inf\Big{\{}\tau(m(\cdot):\,m(\cdot)\text{ satisfies }\partial_{t}m(t)+\operatorname{div}(f(&x,m(t),u(t))m(t))=0,\\ m(0)=&\mu,\ \ u(\cdot)\in L^{0}([0,+\infty);U)\Big{\}}.\end{split}$
Now, we establish the existence of an optimal control whenever the value function is finite.
Theorem 3.8.
Assume that $\mu\in\operatorname{cl}G$ satisfies $\operatorname{Val}(\mu)<+\infty$. Then, there exists a relaxed control $\xi_{*}\in\mathcal{U}$ such that
$\operatorname{Val}(\mu)=\tau(m(\cdot;\mu,\xi_{*})).$
Remark 3.9.
We say that a relaxed control $\xi_{*}\in\mathcal{U}$ is optimal if
$\operatorname{Val}(\mu)=\tau(m(\cdot;\mu,\xi_{*})).$
Proof of Theorem 3.8.
By the definition of the value function, there exists a sequence $\{\xi_{n}\}_{n=1}^{\infty}\subset\mathcal{U}$ such that
$\operatorname{Val}(\mu)=\tau_{*}\triangleq\lim_{n\rightarrow\infty}\tau(m(\cdot;\mu,\xi_{n})).$
Furthermore, we denote $\tau_{n}\triangleq\tau(m(\cdot;\mu,\xi_{n})).$ Since $\mathcal{U}$ is compact, without loss of generality, one may assume that $\{\xi_{n}\}_{n=1}^{\infty}$ converges to some control $\xi_{*}\in\mathcal{U}$. Proposition 3.6 implies that $\{m(\cdot;\mu,\xi_{n})\}_{n=1}^{\infty}$ converges to $m(\cdot;\mu,\xi_{*})$. Since $m(\tau_{n};\mu,\xi_{n})\in M$, passing to the limit and using the closeness of the target set $M$, we conclude that
$m(\tau_{*};\mu,\xi_{*})\in M.$
Therefore, $\operatorname{Val}(\mu)=\tau(m(\cdot;\mu,\xi_{*}))$. ∎
Proposition 3.10.
The value function for problem (3), (5) is lower semicontinuous.
Доказательство.
Let $\{\mu_{k}\}_{k=1}^{\infty}\subset\operatorname{cl}G$ be a sequence of initial conditions for problem (3) converging to $\mu_{0}\in\operatorname{cl}G$. Denote $T_{k}\triangleq\operatorname{Val}(\mu_{k})$. First, we consider the case where $\{T_{k}\}_{k=1}^{\infty}$ is bounded. Without loss of generality, we assume that the sequence $\{T_{k}\}_{k=1}^{\infty}$ converges to some $T_{0}\in[0,+\infty)$. Furthermore, due to Theorem 3.8, for each $k$ there exists a relaxed control $\xi_{k}$ such that, $T_{k}=\tau(m(\cdot;\mu_{k},\xi_{k}))$. Again, without loss of generality, we assume that the whole sequence $\{\xi_{k}\}_{k=1}^{\infty}$ converges to $\xi_{0}\in\mathcal{U}$. Proposition 3.6 gives that the probabilities $m(T_{k};\mu_{k},\xi_{k})$ tend to $m(T_{0};\mu_{0},\xi_{0})$ as $k\rightarrow\infty$. Since the set $M$ is closed, while $m(T_{k};\mu_{k},\xi_{k})\in M$, we have that $m(T_{0};\mu_{0},\xi_{0})\in M$. Thus, using the choice of $T_{k}$, we conclude that
$\lim_{k\rightarrow\infty}\operatorname{Val}(\mu_{k})=T_{0}\geq\tau(m(\cdot;\mu_{0},\xi_{0}))\geq\operatorname{Val}(\mu_{0}).$
If
$\lim_{k\rightarrow\infty}T_{k}=+\infty,$
one obviously has that
$\lim_{k\rightarrow\infty}\operatorname{Val}(\mu_{k})\geq\operatorname{Val}(\mu_{0}).$
3.2 Dynamic programming principle
Theorem 3.11.
The value function for problem (3), (5) satisfies the dynamical programming principle on $\operatorname{cl}G$, i.e., for every $\mu\in\operatorname{cl}G$ and $h\in[0,\operatorname{Val}(\mu)]$,
$\operatorname{Val}(\mu)-h=\inf_{\xi\in\mathcal{U}_{h}}\operatorname{Val}(m(h;\mu,\xi)).$
Доказательство.
If $\xi\in\mathcal{U}_{h}$, $\eta\in\mathcal{U}$, then we denote by $\xi\diamond_{h}\eta$ the relaxed control determined by its disintegration as follows:
$(\xi\diamond_{h}\eta)(\cdot|t)\triangleq\begin{cases}\xi(\cdot|t),&t\in[0,h),\\ \eta(\cdot|t-h),&t\in[h,+\infty).\end{cases}$
We have that
$\begin{split}\inf\limits_{\xi\in\mathcal{U}}\operatorname{Val}(m(h;\mu,\xi))&=\inf\limits_{\xi\in\mathcal{U}_{h}}\inf\limits_{\eta\in\mathcal{U}}\tau(m(\cdot;m(h;\mu,\xi),\eta))\\ =&\inf\limits_{\xi\in\mathcal{U}_{h}}\inf\limits_{\eta\in\mathcal{U}}\tau(m(\cdot;\mu,\xi\diamond_{h}\eta)-h\\ =&\inf\limits_{\xi\in\mathcal{U}}\tau(m(\cdot;\mu,\xi)-h=\operatorname{Val}(\mu)-h.\end{split}$
4 The Dirichlet problem for the Hamilton—Jacobi equation
First, we define the Hamiltonian by the following rule: for every $m\in\mathcal{P}^{2}(\mathbb{R}^{d})$ and $s(\cdot)\in L^{2}(m)$
$H(m,s)=\inf\limits_{u\in U}\int_{\mathbb{R}^{d}}\left\langle s(x),f(x,m,u)\right\rangle m(dx).$
The value function of problem (3), (5) can have infinite value. So, we apply the Kruzhkov transform and define the function $\phi:\operatorname{cl}G\rightarrow\mathbb{R}$ by the rule:
$\phi(m)=1-e^{-\operatorname{Val}(m)}.$ (7)
Notice that $\phi$ takes values in $[0,1]$. Thus, we arrive at the following Dirichlet problem for the Hamilton—Jacobi equation and the transformed value function:
$H(m,D\phi(m))+1-\phi(m)=0,\ m\in G;\phi(m)=0,m\in\partial G.$ (8)
Here, we formally use $D\phi(m)$ to designate the derivative of the function $\phi$. Below, we will develop the viscosity solution approach for this problem that comes back to papers [5, 20]. To this end, we need some concepts of the nonsmooth analysis in the Wasserstein space.
Let $m\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $r>0$, and let $\psi$ be a real-valued function defined on $\mathbb{B}_{r}(m)$.
Definition 4.1.
A subdifferential of the function $\psi$ at the measure $m$ is defined by the rule:
$\begin{split}\partial^{-}\psi(m)\triangleq\Big{\{}s\in L^{2}&(m):\,\text{for every }F\in L_{2}(m)\text{ and }h>0,\\ &\psi((\operatorname{Id}+hF)\sharp m)-\psi(m)\geq h\langle s,F\rangle_{m}+o(h)\Big{\}}.\end{split}$ (9)
Similarly, the superdifferential of the $\psi$ at the measure $m$ is equal to
$\begin{split}\partial^{+}\psi(m)\triangleq\Big{\{}s\in L^{2}&(m):\,\text{for every }F\in L_{2}(m)\text{ and }h>0,\\ &\psi((\operatorname{Id}+hF)\sharp m)-\psi(m)\leq h\langle s,F\rangle_{m}+o(h)\Big{\}}.\end{split}$ (10)
Remark 4.2.
An alternative definition of the sub-/superdifferentials involves the concept of Hadamard directional derivatives. Below, we endow directions determined by functions from $L^{2}(m)$. Let a measure $m$, a radius $r$ and a function $\psi$ be as above. A Hadamard lower directional derivative of the function $\psi$ at the probability $m$ in a direction $F\in L^{2}(m)$ is
$d^{-}_{H}\psi(m;F)\triangleq\liminf\limits_{\begin{subarray}{c}h\rightarrow 0\\ \|F^{\prime}-F\|_{L^{2}(m)}=O(h)\end{subarray}}\frac{\psi((\operatorname{Id}+hF^{\prime})\sharp m)-\psi(m)}{h}.$
Similarly, a Hadamard upper directional derivative of the function $\psi$ at a probability $m\in G$ in a direction $F\in L^{2}(m)$ is equal to
$d^{+}_{H}\psi(m;F)\triangleq\limsup\limits_{\begin{subarray}{c}h\rightarrow 0,\\ \|F^{\prime}-F\|_{L^{2}(m)}=O(h)\end{subarray}}\frac{\psi((\operatorname{\operatorname{Id}}+hF^{\prime})\sharp m)-\psi(m)}{h}.$
One can easily show that
$\partial^{-}\psi(m)=\big{\{}F\in L^{2}(m):\,\left\langle s,F\right\rangle_{m}\leq d^{-}_{H}\psi(m;F)\big{\}},$
$\partial^{+}\psi(m)=\big{\{}F\in L^{2}(m):\,\left\langle s,F\right\rangle_{m}\geq d^{+}_{H}\psi(m;F)\big{\}}.$
Notice that from the very definition of the subdifferential (see  (9)) it directly follows that, if $\psi_{1},\psi_{2}\in\mathbb{B}_{r}(m)\rightarrow\mathbb{R}$ for some positive number $r$, then
$\partial^{-}\psi_{1}(m)+\partial^{-}\psi_{2}(m)\subset\partial^{-}(\psi_{1}+\psi_{2})(m).$ (11)
Analogously, (10) implies the inclusion
$\partial^{+}\psi_{1}(m)+\partial^{+}\psi_{2}(m)\subset\partial^{+}(\psi_{1}+\psi_{2})(m).$
Our approach to the definition of the viscosity solution to the Dirichlet problem is close to one proposed in [37, Definition 18.3, Theorem 18.6] and [6, §IV.3] for the finite dimensional case. In particular, we postulate that the viscosity solution coincides with the supersolution. Notice that in [6] this notion is called an envelope viscosity solution, while in [37] the words ‘minimax solution’ are used.
Definition 4.3.
A bounded lower semicontinuous function $\phi_{1}:\operatorname{cl}G\rightarrow\mathbb{R}$ is called a viscosity supersolution of problem (8) provided that
  • for each $m\in\partial G$, $\phi_{1}(m)=0$;
  • for every $m\in G$ and $s\in\partial^{-}\phi_{1}(m)$
    $H(m,s)+1-\phi_{1}(m)\leq 0.$
Definition 4.4.
A bounded upper semicontinuous function $\phi_{2}:\operatorname{cl}G\rightarrow\mathbb{R}$ is called a viscosity subsolution of problem (8) if
  • for each $m\in\partial G$, $\phi_{2}(m)=0$;
  • $\phi_{2}$ is continuous on $\partial G$;
  • for every $m\in G$ and $s\in\partial^{+}\phi_{2}(m)$ one has that
    $H(m,s)+1-\phi_{2}(m)\geq 0.$
Definition 4.5.
A bounded lower semicontinuos function $\phi:\operatorname{cl}G\rightarrow\mathbb{R}$ is called a viscosity solution of problem (8) if
  • $\phi$ is a supersolution of problem (8);
  • there exists a sequence of functions $\{\phi_{k}\}_{k=1}^{\infty}$ defined on $\operatorname{cl}G$ with values in $\mathbb{R}$ such that each function $\phi_{k}$ is a subsolution of (8) while, for every $m\in G$, one has that
    $\phi(m)=\lim_{k\rightarrow\infty}\phi_{k}(m).$
5 Comparison principle
The proof of the comparison principle relies on the differentiability properties of the squared Wasserstein distance [2, Theorem 10.2.2]. It states that, given $\mu,\mu^{\prime},\nu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $\pi\in\Pi(\mu,\mu^{\prime})$, $\vartheta\in\Pi(\mu,\nu)$ that is an optimal plan between $\mu$ and $\nu$, and $\varpi\in\mathcal{P}((\mathbb{R}^{d})^{3})$ such that $\operatorname{p}^{1,2}\sharp\varpi=\pi$, while $\operatorname{p}^{1,3}\sharp\varpi=\vartheta$, one has that
$\begin{split}\frac{1}{2}W^{2}_{2}(\mu^{\prime},\nu)-\frac{1}{2}W^{2}_{2}(\mu,\nu)-\int_{(\mathbb{R}^{d})^{3}}\langle x-y,x^{\prime}-&x\rangle\varpi(d(x,x^{\prime},y))\\ &\leq\int_{(\mathbb{R}^{d})^{2}}\|x^{\prime}-x\|^{2}\pi(d(x,x^{\prime})).\end{split}$
Letting in this inequality $\mu^{\prime}=(\operatorname{Id}+hF)\sharp\mu$, where $F\in L^{2}(\mu)$, we have that
$\frac{1}{2}W^{2}_{2}((\operatorname{Id}+hF)\sharp\mu,\nu)-\frac{1}{2}W^{2}_{2}(\mu,\nu)-h\int_{\mathbb{R}^{d}}\langle\hat{\vartheta},F\rangle_{\mu}\leq h^{2}\|F\|^{2}_{L^{2}(\mu)}.$ (12)
Here, $\hat{\vartheta}\in L^{2}(\mu)$ is a barycenter of the plan $\vartheta$ defined by the rule:
$\hat{\vartheta}(x)\triangleq\int_{\mathbb{R}^{d}}(x-y)\vartheta(dy|x).$ (13)
Theorem 5.1.
Let $\phi_{1}:\operatorname{cl}G\rightarrow[0,1]$ be a supersolution of problem (8) and let $\phi_{2}:\operatorname{cl}G\rightarrow[0,1]$ be a subsolution of problem (8). Then, for each $m\in\operatorname{cl}G$ $\phi_{1}(m)\geq\phi_{2}(m).$
Доказательство.
Below, we duplicate the space and work with $G^{2}$. The distance on this space between two pairs $(m_{1},m_{2})$ and $(\mu_{1},\mu_{2})$ is assumed to be equal to
$d((m_{1},m_{2}),(\mu_{1},\mu_{2}))\triangleq W_{2}(m_{1},\mu_{1})+W_{2}(m_{2},\mu_{2}).$
Additionally, without loss of generality, we assume that the Lipschitz constant of the function $f$ is greater than $1/4$.
Furthermore, we introduce the functions: $\Phi:G^{2}\rightarrow\mathbb{R}$ and $\rho:(G^{2})^{2}\rightarrow\mathbb{R}$ by the following rules:
$\Phi(m_{1},m_{2})\triangleq\phi_{2}(m_{2})-\phi_{1}(m_{1})-\frac{1}{2}W_{2}^{2}(m_{1},m_{2}),$
$\rho(m_{1},m_{2};\mu_{1},\mu_{2})\triangleq(4L-1)\big{[}W_{2}^{2}(m_{1},\mu_{1})+W_{2}^{2}(m_{2},\mu_{2})\big{]}.$
Notice that $\rho$ is a gauge-type function in the sense of [30, Definition 2], i.e.,
  • $\rho(m_{1},m_{2};m_{1},m_{1})=0$ for every $m_{1},m_{2}\in\mathcal{P}^{2}(\mathbb{R}^{d})$;
  • given $\epsilon>0$ there exists $\beta>0$ such that, for all $(m_{1},m_{2})\in G^{2}$, $(\mu_{1},\mu_{2})\in G^{2}$ inequality $\rho(m_{1},m_{2};\mu_{1},\mu_{2})\leq\beta$ implies $d((m_{1},m_{2}),(\mu_{1},\mu_{2}))\leq\epsilon$.
Let $\hat{m}$ be a point in $G$. By Borwein—Preiss Variational Principle [30, Corollary 5], there exists a pair $(m_{1}^{*},m_{2}^{*})\in G^{2}$ such that
$\Phi(m_{1}^{*},m_{2}^{*})-\rho(\hat{m},\hat{m};m_{1}^{*},m_{2}^{*})\geq\Phi(\hat{m},\hat{m}),$ (14)
and, for every pair $(m_{1},m_{2})\in G^{2}$,
$\Phi(m_{1}^{*},m_{2}^{*})\geq\Phi(m_{1},m_{2})-\rho(m_{1},m_{2};m_{1}^{*},m_{2}^{*}).$ (15)
Now, let us introduce the following functions:
$\Psi_{1}(m_{1})=\phi_{1}(m_{1})-\phi_{2}(m_{2}^{*})+\frac{1}{2}W_{2}^{2}(m_{1},m_{2}^{*}),$
$\Psi_{2}(m_{2})=\phi_{2}(m_{2})-\phi_{1}(m_{1}^{*})-\frac{1}{2}W_{2}^{2}(m_{1}^{*},m_{2}).$
We will show that $\mathbf{0}(\cdot)\in\partial^{-}\Psi_{1}(m_{1}^{*})$ and $\mathbf{0}(\cdot)\in\partial^{+}\Psi_{2}(m_{2}^{*})$. Hereinafter, $\mathbf{0}(\cdot)$ is a function that is equal to zero everywhere. To this end, given $F\in L^{2}(m_{2}^{*})$ and $h>0$, we compute
$\begin{split}\Psi_{1}((\operatorname{Id}+h&F)\sharp m_{1}^{*})-\Psi_{1}(m_{1}^{*})=-\Phi((\operatorname{Id}+hF)\sharp m_{1}^{*},m_{2}^{*})+\Phi(m_{1}^{*},m_{2}^{*})\\ &\geq-\rho((\operatorname{Id}+hF)\sharp m_{1}^{*},m_{2}^{*};m_{1}^{*},m_{2}^{*})\\ &=-(4L-1)W_{2}^{2}((\operatorname{Id}+hF)\sharp m_{1}^{*},m_{1}^{*})\geq-(4L-1)h^{2}\|F\|^{2}_{L_{2}(m_{1}^{*})}.\end{split}$ (16)
Here, we first used the definition of the function $\Psi_{2}$ and, then, (15). Inequality (16) and the equivalent definition of the subdifferential (see (9)) imply that $\mathbf{0}(\cdot)\in\partial^{-}\Psi_{1}(m_{1}^{*})$. The inclusion $\mathbf{0}(\cdot)\in\partial^{+}\Psi_{2}(m_{2}^{*})$ is proved in the similar way.
Now, let $\vartheta$ be an optimal plan between $m_{1}^{*}$ and $m_{1}^{*}$. We put
$s_{1}(x_{1})\triangleq\int_{\mathbb{R}^{d}}(x_{1}-x_{2})\vartheta(dx_{2}|x_{1}),\ \ s_{2}(x_{2})\triangleq\int_{\mathbb{R}^{d}}(x_{2}-x_{1})\vartheta(dx_{1}|x_{2}),$ (17)
where $(\vartheta(\cdot|x_{1}))_{x_{1}\in\mathbb{R}^{d}}$ (respectively, $(\vartheta(\cdot|x_{2}))_{x_{2}\in\mathbb{R}^{d}}$) denotes the disintegration of $\vartheta$ w.r.t. $m_{1}^{*}$ (respectively, w.r.t. $m_{2}^{*}$). If
$w_{1}(m)\triangleq\frac{1}{2}W_{2}(m,m_{2}^{*}),\ \ w_{2}(m)\triangleq\frac{1}{2}W_{2}(m_{1}^{*},m),$ (18)
then inequality (12) implies that
$s_{1}\in\partial^{+}w_{1}(m_{1}^{*}),\ \ s_{2}\in\partial^{+}w_{2}(m_{2}^{*}).$ (19)
Thanks to equality $\phi_{1}(m_{1})=\Psi_{1}(m_{1})+\phi_{2}(m_{2}^{*})-w_{1}(m_{1})$, inclusions (11), (19) and the fact that $\mathbf{0}(\cdot)\in\partial^{-}\Psi_{1}(m_{1}^{*})$ give that
$-s_{1}\in\partial^{-}\phi_{1}(m_{1}^{*}).$
Similarly, since $\mathbf{0}(\cdot)\in\partial^{+}\Psi_{2}(m_{2}^{*})$, we conclude that
$s_{2}\in\partial^{+}\phi_{2}(m_{2}^{*}).$
Now, let us recall that $\phi_{1}$ is a supersolution of (8). By substituting the measure $m_{1}^{*}$ and the element of the subdifferential $-s_{1}$ into the definition of the supersolution, we conclude that
$H(m_{1}^{*},-s_{1})+1-\phi_{1}(m_{1}^{*})\leq 0.$
Analogously, since $\phi_{2}$ is a subsolution of (8) and $s_{2}\in\partial^{+}\phi_{2}(m_{2}^{*})$, one has
$H(m_{2}^{*},s_{2})+1-\phi_{2}(m_{2}^{*})\geq 0.$
Thus,
$H(m_{1}^{*},-s_{1})-H(m_{2}^{*},s_{2})-\phi_{1}(m_{1}^{*})+\phi_{2}(m_{2}^{*})\leq 0.$
By the definition of the Hamiltonian, we have
$\begin{split}\inf\limits_{u\in U}\int_{\mathbb{R}^{d}}\big{\langle}-s_{1}(x),f(x,m_{1}^{*},u)\big{\rangle}m_{1}^{*}(dx)-\inf\limits_{u\in U}\int_{\mathbb{R}^{d}}\big{\langle}s_{2}(y),f(&y,m_{2}^{*},u)\big{\rangle}m_{2}^{*}(dy)\\ &-\phi_{1}(m_{1}^{*})+\phi_{2}(m_{2}^{*})\leq 0.\end{split}$
Let $u_{1}^{*}$ be such that
$\begin{split}\int_{\mathbb{R}^{d}}\big{\langle}-s_{1}(x),f(x,m_{1}^{*},u_{1}^{*})\big{\rangle}m_{1}^{*}(dx)=\inf\limits_{u\in U}\int_{\mathbb{R}^{d}}\big{\langle}-s_{1}(x),f(x,m_{1}^{*},u)\big{\rangle}m_{1}^{*}(dx).\end{split}$
Using this, the definitions of the functions $s_{1}$ and $s_{2}$ (see (17)) and the fact that $\vartheta$ is a plan between $m_{1}^{*}$ and $m_{1}^{*}$, we conclude the following:
$\begin{split}\phi_{2}(&m_{2}^{*})-\phi_{1}(m_{1}^{*})\\ \leq&\int_{(\mathbb{R}^{d})^{2}}\langle x_{2}-x_{1},f(x_{2},m_{2}^{*},u_{1}^{*})-f(x_{1},m_{1}^{*},u_{1}^{*})\rangle\vartheta(d(x_{1},x_{2})).\end{split}$ (20)
Since $f$ is Lipschitz continuous with the constant $L$, we have that
$\begin{split}\int_{(\mathbb{R}^{d})^{2}}\langle&x_{2}-x_{1},f(x_{2},m_{2}^{*},u_{1}^{*})-f(x_{1},m_{1}^{*},u_{1}^{*})\rangle\vartheta(d(x_{1},x_{2}))\\ &\leq\int_{(\mathbb{R}^{d})^{2}}L\big{(}\|x_{1}-x_{2}\|^{2}+\|x_{1}-x_{2}\|\cdot W_{2}(m_{1}^{*},m_{2}^{*})\big{)}\vartheta(d(x_{1},x_{2})\\ &\leq 2LW_{2}^{2}(m_{1}^{*},m_{2}^{*}).\end{split}$
Here, we used the choice of $\vartheta$ as an optimal plan between $m_{1}^{*}$ and $m_{1}^{*}$. Therefore, we estimate the right-hand side in (20) and obtain
$\phi_{2}(m_{2}^{*})-\phi_{1}(m_{1}^{*})\leq 2LW_{2}^{2}(m_{1}^{*},m_{2}^{*}).$
This and the definition of the function $\Phi$ yield that
$\Phi(m_{1}^{*},m_{2}^{*})+\frac{1}{2}W_{2}^{2}(m_{1}^{*},m_{2}^{*})\leq 2LW_{2}^{2}(m_{1}^{*},m_{2}^{*}).$
From this and (14), it follows that
$\Phi(\hat{m},\hat{m})+\rho(\hat{m},\hat{m};m_{1}^{*},m_{2}^{*})+\frac{1}{2}W_{2}^{2}(m_{1}^{*},m_{2}^{*})\leq 2LW_{2}^{2}(m_{1}^{*},m_{2}^{*}).$ (21)
Now we use the definition of gauge function $\rho$ and notice that
$W_{2}^{2}(m_{1}^{*},m_{2}^{*})\leq\big{[}W_{2}(\hat{m},m_{1}^{*})+W_{2}(\hat{m},m_{2}^{*})\big{]}^{2}\leq\frac{2}{4L-1}\rho(\hat{m},\hat{m};m_{1}^{*},m_{2}^{*}).$
Therefore, (21) directly implies the following inequality:
$\phi_{2}(\hat{m})-\phi_{1}(\hat{m})=\Phi(\hat{m},\hat{m})\leq 0.$
Since $\hat{m}$ was arbitrarily, we have that $\phi_{2}\leq\phi_{1}$ on the whole domain $G$. ∎
Theorem 5.1 and Definition 4.5 imply the following.
Corollary 5.2.
There exists at most one solution of Dirichlet problem (8).
6 Characterization of the value function
The aim of this section is to show that the Kruzhkov transform of the value function defined by (7) is a viscosity solution of (8). This, in particular, gives the existence result for  (8).
Our proof will use the relaxations of the time-optimal problem in the space of measures. To define them, for every $\varepsilon>0$, we put
$M^{\varepsilon}\triangleq\operatorname{cl}\Big{(}\big{\{}m\in\mathcal{P}^{2}(\mathbb{R}^{d}):\,\operatorname{dist}(m,M)\leq\varepsilon\big{\}}\Big{)}.$
Here,
$\operatorname{dist}(m,M)\triangleq\inf\big{\{}W_{2}(m,\mu):\,\mu\in M\big{\}}.$
Notice that
$M^{\varepsilon_{1}}\subset M^{\varepsilon_{2}}\text{ whenever }\varepsilon_{1}\leq\varepsilon_{2}$ (22)
and
$M=\bigcap_{\varepsilon>0}M^{\varepsilon}.$
We now consider the time-optimal problem for the target set $M^{\varepsilon}$. As above, given $m(\cdot)\in C([0,+\infty);\mathcal{P}^{2}(\mathbb{R}^{d}))$, we set
$\tau^{\varepsilon}(m(\cdot))\triangleq\inf\big{\{}t\in[0,+\infty):\,m(t)\in M^{\varepsilon}\big{\}}.$
By (22), we have that, if $\varepsilon_{1}\leq\varepsilon_{2}$,
$\tau^{\varepsilon_{1}}(m(\cdot))\geq\tau^{\varepsilon_{2}}(m(\cdot)).$ (23)
We define the value function for the time-optimal problem with the target set $M^{\varepsilon}$ by the rule:
$\operatorname{Val}^{\varepsilon}(\mu)\triangleq\inf\Big{\{}\tau^{\varepsilon}(m(\cdot;\mu,\xi):\,\xi\in\mathcal{U}\Big{\}}.$
As for the original value function we have the following properties.
Proposition 6.1.
The function $\operatorname{Val}^{\varepsilon}$ is lower semicontinuous. Moreover, $\operatorname{Val}^{\varepsilon}(\mu)=0$ whenever $\mu\in M^{\varepsilon}$.
Proposition 6.2.
If $\varepsilon>0$, $\mu\in G$ is such that $\operatorname{Val}^{\varepsilon}(\mu)<+\infty$. Then, there exists $\xi_{\varepsilon,\mu}$ satisfying
$\operatorname{Val}^{\varepsilon}(\mu)=\tau^{\varepsilon}(m(\cdot;\mu,\xi_{\varepsilon,\mu})).$
Proposition 6.3.
For every $\mu\in\operatorname{cl}G$ and $h\in[0,\operatorname{Val}^{\varepsilon}(\mu)]$,
$\operatorname{Val}^{\varepsilon}(\mu)-h=\inf_{\xi\in\mathcal{U}_{h}}\operatorname{Val}(m(h;\mu,\xi)).$
Proposition 6.4.
If $\varepsilon_{1}\leq\varepsilon_{2}$, then
$\operatorname{Val}^{\varepsilon_{1}}(\mu)\geq\operatorname{Val}^{\varepsilon_{2}}(\mu).$
Доказательство.
If $\operatorname{Val}^{\varepsilon}(\mu)=+\infty$, there is nothing to prove. We will consider the case where $\operatorname{Val}^{\varepsilon}(\mu)$ is finite. Let $\xi_{\varepsilon_{1},\mu}$ be such that $\operatorname{Val}^{\varepsilon_{1}}(\mu)=\tau^{\varepsilon_{1}}(m(\cdot;\mu,\xi_{\varepsilon_{1},\mu})).$ By (23), we have that
$\tau^{\varepsilon_{1}}(m(\cdot;\mu,\xi_{\varepsilon_{1},\mu}))\geq\tau^{\varepsilon_{2}}(m(\cdot;\mu,\xi_{\varepsilon_{1},\mu})).$
This, together with the definition of the function $\operatorname{Val}$ yield the statement of the proposition. ∎
Recall that above we introduced the function $\phi$ that is the Kruzhkov transform of the value function $\operatorname{Val}$ (see (7)).
Theorem 6.5.
The function $\phi$ is a viscosity solution of Dirichlet problem for Hamilton-Jacobi equation (8).
Доказательство.
We split the proof into four steps.
Step 1. Here, we prove that the function $\phi$ is a supersolution to (8). First, notice that $\phi$ is bounded and lower semicontinuous, while $\phi(\mu)=0$ when $\mu\in\partial G$.
Furthermore, we choose a probability $\mu\in G$ and a function $s\in\partial^{-}\phi^{*}(\mu)$. By Remark 4.2, we have that, for every $F\in L^{2}(\mu)$,
$\left\langle s,F\right\rangle_{\mu}\leq d^{-}_{H}\phi(\mu;F).$ (24)
The dynamic programmin principle (see Theorem 3.11) can be rewritten as follows: for each $h\in[0,-\ln(1-\phi(x)]$,
$(1-\phi(\mu))e^{h}=\sup_{\xi\in\mathcal{U}_{h}}(1-\phi(m(h;\mu,\xi))).$
Due to Theorem 3.8, there exists $\xi_{h}\in\mathcal{U}_{h}$ such that
$(1-\phi(\mu))e^{h}=(1-\phi(m(h;\mu,\xi_{h}))).$
Let
[…]
where $X(\cdot;y,\mu,\xi_{h})$ is a trajectory such that
$\frac{d}{dt}X(t;y,\mu,\xi_{h})=\int_{U}f(X(t;y,\mu,\xi),m(t;\mu,\xi_{h}),u)\xi_{h}(du|t),\ \ X(0;y,\mu,\xi)=y.$
Notice that $m(h;\mu,\xi_{h})=(\operatorname{Id}+hF_{h})\sharp\mu$. Thus, we have that
$\phi((\operatorname{Id}+hF_{h})\sharp\mu)-\phi(\mu)=(\phi(\mu)-1)\cdot(e^{h}-1).$ (25)
Furthermore, Proposition A.3 gives the existence of a sequence $\{h_{k}\}_{k=1}^{\infty}$ and a probability $\zeta\in\mathcal{P}(U)$ such that $h_{k}\rightarrow 0$, $\|F_{h_{k}}-F\|_{L^{2}(\mu)}\rightarrow 0$ as $k\rightarrow\infty$ for $F(x)\triangleq\int_{U}f(x,\mu,u)\zeta(du)$. Therefore, from (25), we conclude that
$d^{-}_{H}\phi(\mu;F)\leq\phi(\mu)-1.$
This and (24) imply that
$\int_{U}\int_{\mathbb{R}^{d}}\left\langle s(x),f(x,\mu,u)\right\rangle\mu(dx)\zeta(du)\leq\phi(\mu)-1.$
Using the definition of the Hamiltonian, we obtain the following inequality, for every $\mu\in G$ and $s\in\partial^{-}\phi(\mu)$,
$H(\mu,s)+1-\phi(\mu)\leq 0.$
Thus, $\phi$ is a supersolution of problem (8).
Step 2. Let $\psi^{\varepsilon}$ be the Kruzhkov transform of the function $\operatorname{Val}^{\varepsilon}$, i.e.,
$\psi^{\varepsilon}(\mu)\triangleq 1-e^{\operatorname{Val}^{\varepsilon}(\mu)}.$
Moreover, we put $\phi^{\varepsilon}$ to be an upper envelope of the function $\psi^{\varepsilon}$. This means that, for every $\mu\in G$,
$\phi^{\varepsilon}(\mu)\triangleq\limsup_{\mu^{\prime}\rightarrow\mu,\ \ \mu^{\prime}\in\operatorname{cl}G}\psi^{\varepsilon}(\mu^{\prime})=\lim_{\delta\downarrow 0}\sup\big{\{}\psi^{\varepsilon}(\mu^{\prime}):\,W_{2}(\mu,\mu^{\prime})\leq\delta,\ \ \mu^{\prime}\in G\setminus M^{\varepsilon}\big{\}}.$
Step 3. Here we prove that each function $\phi^{\varepsilon}$ is a subsolution of (8). Notice that, by the construction, $\psi^{\varepsilon}(\mu)=0$ whenever $\mu\in\operatorname{int}M^{\varepsilon}$. Thus, $\phi^{\varepsilon}=0$ on $\partial M$ and $\phi^{\varepsilon}$ is continuous on $\partial M$.
Moreover, the very definition of the function $\phi^{\varepsilon}$ gives that it is bounded and upper semicontinuous.
Furthermore, let us show that the function $\phi^{\varepsilon}$ satisfies the following condition: for every $\mu\in G$ and $s\in\partial^{+}\phi^{\varepsilon}(\mu)$,
$H(\mu,s)+1-\phi^{\varepsilon}(\mu)\geq 0$ (26)
First, we consider the case when $\operatorname{Val}^{\varepsilon}(\mu)>0$.
By the construction of the function $\phi^{\varepsilon}$, there exists a $\{\mu_{k}\}_{k=1}^{\infty}\subset G\setminus M^{\varepsilon}$ converging to $\mu$, such that
$\lim_{k\rightarrow\infty}\psi^{\varepsilon}(\mu_{k})=\phi^{\varepsilon}(\mu).$
Without loss of generality, we assume that $\psi^{\varepsilon}(\mu_{k})\geq\hat{h}\triangleq\phi^{\varepsilon}(\mu)/2>0$. By the definition of the superdifferential, for every $F\in L^{2}(\mu)$, we have
$\left\langle s,F\right\rangle_{\mu}\geq d^{+}_{H}\phi^{\varepsilon}(\mu;F).$
We choose $f^{0}(x)\triangleq f(x,\mu,u_{0})$, where $u_{0}$ is such that
$H(\mu,s)=\left\langle s,f(\cdot,\mu,u_{0})\right\rangle_{\mu}.$ (27)
Hence,
$H(\mu,s)=\left\langle s,f^{0}\right\rangle_{\mu}\geq d^{+}_{H}\phi^{\varepsilon}(\mu;f^{0}).$ (28)
Furthermore, we evaluate $d^{+}_{H}\phi^{\varepsilon}(\mu;f^{0})$ in the following way.
Since $\psi^{\varepsilon}$ is the Kruzhkov transform of the function $\operatorname{Val}^{\varepsilon}$, the dynamical programming principle for problem with the target $M^{\varepsilon}$ (see Proposition 6.3) implies that, for $h\in[0,-\ln(1-\psi^{\varepsilon}(\mu_{k}))]$,
$(1-\psi^{\varepsilon}(\mu_{k}))e^{h}=\sup_{\xi\in\mathcal{U}_{h}}(1-\psi^{\varepsilon}(m(h;\mu_{k},\xi))).$ (29)
We consider $h\in(0,\hat{h})$ and the constant relaxed control on $\mathcal{U}_{h}$ defined by the rule $\xi^{0}(\cdot|t)\triangleq\delta_{u^{0}}$. Inequality (29) gives that
$(1-\psi^{\varepsilon}(\mu_{k}))e^{h}\geq 1-\psi^{\varepsilon}(m(h;\mu_{k},\xi^{0})).$
Using the definition of the function $\phi^{\varepsilon}$, the choice of the sequence $\{\mu_{k}\}_{k=1}^{\infty}$ and Proposition 3.6, we conclude that
$(1-\phi^{\varepsilon}(\mu))e^{h}\geq 1-\phi^{\varepsilon}(m(h;\mu,\xi^{0})).$
Now, as above, let $X(\cdot;y,\mu,\xi^{0})$ satisfy
$\frac{d}{dt}X(t;y,\mu,\xi^{0})=f(X(t;y,\mu,\xi^{0}),m(t;\mu,\xi^{0}),u^{0}),\ \ X(0;y,\mu,\xi^{0})=y.$
Set
[…]
Proposition A.3 yields that
$\|F_{h}-f^{0}\|_{L^{2}(\mu)}\rightarrow 0\text{ as }h\rightarrow 0.$
Additionally, notice that $X(h;y,\mu,\xi^{0})=y+F_{h}(y)\cdot h.$ Thus,
$m(h;\mu,\xi^{0})=(\operatorname{Id}+hF_{h})\sharp\mu.$
Therefore,
$\phi^{\varepsilon}((\operatorname{Id}+hF_{h})\sharp\mu)-\phi^{\varepsilon}(\mu)\geq(\phi(\mu)-1)\cdot(e^{h}-1).$
Dividing both parts on $h$ and using the definition of the Hadamard upper derivative, we conclude that
$d^{+}_{H}\phi^{\varepsilon}(\mu,f_{0})\geq\phi^{*}(\mu)-1.$
This and (28) imply (26) for each $\mu\in G$ satisfying $\phi^{\varepsilon}(\mu)>0$ and every $s\in\partial^{+}\phi^{\varepsilon}(\mu)$.
Now, let $\phi^{\varepsilon}(\mu)=0$. As above, we choose $s\in\partial^{+}\phi^{\varepsilon}(\mu)$. Since $\phi^{\varepsilon}(\mu)=0$, we have that $d^{+}_{H}\phi^{\varepsilon}(\mu;f^{0})\geq 0$. Here, $f^{0}(\cdot)=f(\cdot,\mu,u^{0})$ with $u^{0}$ satisfying (27). Thus,
$H(\mu,s)\geq d^{+}_{H}\phi^{\varepsilon}(\mu;f^{0})\geq 0.$
This implies (26) for each $\mu$ satisfying $\phi^{\varepsilon}(\mu)=0$ and every $s\in\partial^{+}\phi^{\varepsilon}(\mu)$.
Step 4. We claim that, for each $\mu\in G$,
$\lim_{\varepsilon\downarrow 0}\phi^{\varepsilon}(\mu)=\phi(\mu).$ (30)
This equality directly follows from the inequality
$\lim_{\varepsilon\downarrow 0}\operatorname{Val}^{\varepsilon}(\mu)\geq\operatorname{Val}(\mu).$ (31)
the monotonicity of the Kruzhkov transform and the comparison principle (see Theorem 5.1). Equality (30) and Steps 1–3 give that the function $\phi$ is a viscosity solution of (8).
So, it remains to prove (31). First, recall that Proposition 6.4 means that $\varepsilon\mapsto\operatorname{Val}^{\varepsilon}(\mu)$ decreases. Therefore, there exists a limit $\lim_{\varepsilon\downarrow 0}\operatorname{Val}^{\varepsilon}(\mu)$.
Now, we consider the case where
$T\triangleq\lim_{\varepsilon\downarrow 0}\operatorname{Val}^{\varepsilon}(\mu)<\infty.$ (32)
By Proposition 6.2, for each $\varepsilon>0$, there exists a relaxed control $\xi_{\varepsilon,\mu}$ such that
$\operatorname{Val}^{\varepsilon}(\mu)=\tau^{\varepsilon}(m(\cdot;\mu,\xi_{\varepsilon,\mu}))$
and $m(\operatorname{Val}^{\varepsilon}(\mu);\mu,\xi_{\varepsilon,\mu})\in M^{\varepsilon}$. Furthermore, the fact that $\mathcal{U}$ is compact implies that there exist a sequence $\{\varepsilon_{k}\}_{k=1}^{\infty}$ converging to zero and a relaxed control $\xi$ such that $\xi_{\varepsilon_{k},\mu}$ tend to $\xi$. Thus, by Proposition 3.6, we have that $\{m(\cdot;\mu,\xi_{\varepsilon_{k},\mu})\}_{k=1}^{\infty}$ converges to $m(\cdot;\mu,\xi)$. Using the definition of the sets $M^{\varepsilon}$, the functions $\operatorname{Val}^{\varepsilon}$ and the convergence (32), we conclude that $m(T;\mu,\xi)\in M.$ Hence,
$\operatorname{Val}(\mu)\leq T=\lim_{\varepsilon\downarrow 0}\operatorname{Val}^{\varepsilon}(\mu).$
The case where
$\lim_{\varepsilon\downarrow 0}\operatorname{Val}^{\varepsilon}(\mu)=\infty$
is obvious. ∎
7 Topological properties of the value function
First, we consider a sufficient condition assuring the continuity of the value function.
Theorem 7.1.
Assume that the viscosity solution of Dirichlet problem (8) is continuous on $\partial G$. Then, it is continuous at every point of $\operatorname{cl}G$.
Remark 7.2.
Theorem 7.1 means that, if the value function is continuous on $\partial G$, then it is continuous everywhere on $\operatorname{cl}G$. Moreover, the continuity of the value function on $\partial G$ directly follows from the small time local attainability property introduced in [29]. Indeed, for the examined case, it takes the following form: given $\varepsilon>0$, one can find $\delta>0$ such that $\operatorname{Val}(\mu)\leq\varepsilon$ whenever $\mu$ satisfies $\inf_{\mu^{\prime}\in M}W_{2}(\mu,\mu^{\prime})<\delta$. Obviously, under the small time local attainability assumption, the value function is continuous at every point of $\partial G$. If, additionally, $M$ is compact, the small time local attainability property is equivalent to the continuity of the value function on $\partial G$. For the space of probability measures, this property was considered in [17, 20]. In these studies, the regularity of the value function for the time-optimal problem in the space of probability measures was investigated under the assumption that the problem satisfies the small-time local attainability property, with the target being a hyperplane in the space of measures.
Proof of Theorem 7.1.
Let $\phi$ denote the aforementioned viscosity solution. For every $\mu\in\operatorname{cl}G$, we put
$\phi^{*}(\mu)\triangleq\limsup_{\mu^{\prime}\rightarrow\mu}\phi(\mu).$
Notice that the function $\phi^{*}:\operatorname{cl}G\rightarrow\mathbb{R}$ is upper semicontinuous. By following the same reasoning as in Step 3 of the proof of Theorem 6.5, one can show that
$H(\mu,s)+1-\phi^{*}(\mu)\leq 0$
for every $\mu\in G$ and $s\in\partial^{+}\phi^{*}(\mu)$. Moreover, the continuity of the function $\phi$ on $\partial G$ gives that $\phi^{*}(\mu)=0$ for every $\mu\in\partial G$. Thus, $\phi^{*}$ is a subsolution of (8). The comparison principle (see Theorem 5.1) gives that $\phi^{*}(\mu)\leq\phi(\mu)$ on $\operatorname{cl}G$. Since the opposite inequality directly follows from the definition of the function $\phi^{*}$, we conclude that $\phi=\phi^{*}$ and the function $\phi$ is continuous. ∎
Now, let us consider the perturbation of the dynamics of the time-optimal problem. Since the value function is generally only lower semicontinuous, we need a concept of convergence. In our opinion, the most natural convergence here is the $\Gamma$-convergence defined as follows (see [12, Definition 1.5]).
Definition 7.3.
For each natural $n$, let $F_{n}$ be a functional on defined on a topological space $X$ with values in {$\mathbb{R}\cup\{+\infty\}$. The sequence $\{F_{n}\}_{n=1}^{\infty}$ is said to $\Gamma$-converge to a functional $F:X\rightarrow\mathbb{R}\cup\{+\infty\}$ if following two conditions hold:
  1. 1.
    for every sequence $\{x_{n}\}_{n=1}^{\infty}\in X$ satisfying $\lim\limits_{n\rightarrow\infty}x_{n}=x$,
    $F(x)\leq\liminf\limits_{n\rightarrow\infty}F_{n}(x_{n});$
  2. 2.
    for each $x\in X$, there is a sequence $\{x_{n}\}_{n=1}^{\infty}\in X$ such that
    $\lim\limits_{n\rightarrow\infty}x_{n}=x\text{ and }F(x)\geq\limsup\limits_{n\rightarrow\infty}F_{n}(x_{n}).$
If $\{F_{n}\}_{n=1}^{\infty}$ $\Gamma$-converges to $F$, then we write
$F=\Gamma\mbox{-}\lim\limits_{n\rightarrow\infty}F_{n}.$
Now we consider a perturbation of the dynamics of (3). For each natural $n$, let a function $f_{n}:\mathbb{R}^{d}\times\mathcal{P}^{2}(\mathbb{R}^{d})\times U\rightarrow\mathbb{R}^{d}$, $n\in\mathbb{N}$ be continuous and Lipschitz continuous w.r.t. $x$ and $m$. If $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $\xi\in\mathcal{U}$, then there exists a unique distributional solution to
$\partial_{t}m_{n}(t)+\operatorname{div}\Bigg{(}\int_{U}f_{n}(x,m(t),u)m(t)\xi(du|t)\Bigg{)}=0,\ \ m(0)=\mu.$
We denote this solution by $m(\cdot;\mu,\xi)$. The value function of the perturbed time-optimal is defined by the rule:
$\operatorname{Val}_{n}(\mu)\triangleq\inf\big{\{}\tau(m_{n}(\cdot;\mu,\xi):\,\xi\in\mathcal{U}\big{\}}.$
Theorem 7.4.
Assume the following convergence of the dynamics $f_{n}$ to $f$: for each $c>0$,
$\begin{split}\sup\Big{\{}\|f_{n}(x,&m,u)-f(x,m,u)\|:\\ &x\in\mathbb{R}^{d},\,m\in\mathcal{P}^{2}(\mathbb{R}^{d}),\,\varsigma(m)\leq c,\,u\in U\Big{\}}\rightarrow 0\text{ as }n\rightarrow\infty.\end{split}$
Then,
$\Gamma\mbox{-}\lim\limits_{n\rightarrow\infty}\operatorname{Val}_{n}=\operatorname{Val}.$
Доказательство.
Let us prove the first property in Definition 7.3. Let $\{\mu_{n}\}_{n=1}^{\infty}\in\mathcal{P}^{2}(\mathbb{R}^{d})$ be a sequence converging to $\mu$. If $\operatorname{Val}_{n}(\mu_{n})\rightarrow+\infty$ as $n\rightarrow\infty$, then, obviously,
$\operatorname{Val}(\mu)\leq\lim_{n\rightarrow\infty}\operatorname{Val}_{n}(\mu_{n}).$
If the limit is finite, we, without loss of generality, can assume that the whole sequence $\{\operatorname{Val}_{n}(\mu_{n})\}_{n=1}^{\infty}$ converges to some number $\theta$. By Theorem 3.8, for each natural $n$, there exists a control $\xi_{n}\in\mathcal{U}$ such that
$\theta_{n}\triangleq\tau(m_{n}(\cdot;\mu_{n},\xi_{n}))=\operatorname{Val}_{n}(\mu_{n}).$
Additionally, for sufficiently large $n$ one has that $\tau(m_{n}(\cdot;\mu_{n},\xi_{n}))<2\theta$. Since $\mathcal{U}$ is compact, we passing, if necessary, to a subsequence, have that $\xi_{n}\rightarrow\xi$ as $n\rightarrow\infty$. Thus, Proposition 3.6 gives that $\{m_{n}(\cdot;\mu_{n},\xi_{n})\}_{n=1}^{\infty}$ converges to $m(\cdot;\mu,\xi)$ on $C([0,2\theta],\mathcal{P}^{2}(\mathbb{R}^{d}))$. Since, $m(\theta_{n};\mu_{n},\xi_{n})\in M,$ whereas $M$ is closed, we have that $m(\theta;\mu,\xi)\in M$. Hence,
$\operatorname{Val}(\mu)\leq\tau(m(\cdot;\mu,\xi))\leq\theta=\liminf_{n\rightarrow\infty}\operatorname{Val}_{n}(\mu_{n}).$
This proves the first property of the $\Gamma$-convergence.
To prove the second property, we choose a measure $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$. First, we assume the case when $\operatorname{Val}(\mu)<\infty$. As above, by Theorem 3.8, one can find a control $\xi\in\mathcal{U}$, such that $\operatorname{Val}(\mu)=\tau(m(\cdot;\mu,\xi))$. For shortness, we put $\theta\triangleq\operatorname{Val}(\mu)$, $\nu\triangleq m(\theta,\mu,\xi)$. Notice that $\nu\in M$. Now, let $m^{\prime}_{n}(\cdot):[0,\theta]\rightarrow\mathcal{P}^{2}(\mathbb{R}^{d})$ satisfy
$\partial_{t}m_{n}^{\prime}(t)+\operatorname{div}\Bigg{(}\int_{U}f_{n}(\cdot,m_{n}^{\prime}(t),u)\xi(du|t)m_{n}^{\prime}(t)\Bigg{)}=0,\ \ m_{n}^{\prime}(\theta)=\nu.$
One can use the time reversing and prove the existence of such motion. Additionally, Proposition 3.6 applied in the reverse time gives that
$\sup_{t\in[0,\theta]}W_{2}(m_{n}^{\prime}(t),m(t,\mu,\xi))\rightarrow 0\text{ as }n\rightarrow\infty.$
We put $\mu_{n}\triangleq m_{n}^{\prime}(0)$. By construction, $m_{n}(\cdot,\mu_{n},\xi)=m_{n}^{\prime}(\cdot)$. Moreover, $W_{2}(\mu_{n},\mu)$ tends to 00 when $n\rightarrow\infty$. In particular, $\mu_{n}\in G$ for sufficiently large $n$. Thus, due to the fact that $m_{n}^{\prime}(\theta)=\nu\in M$, $\tau(m_{n}(\cdot,\mu_{n},\xi))\leq\theta.$ Hence,
$\operatorname{Val}_{n}(\mu_{n})\leq\tau(m_{n}(\cdot;\mu_{n},\xi))\leq\tau(m(\cdot;\mu,\xi))=\operatorname{Val}(\mu).$
This means that $\limsup\limits_{n\rightarrow\infty}\operatorname{Val}_{n}(\mu_{n})\leq\operatorname{Val}(\mu).$
If $\operatorname{Val}(\mu)=+\infty$, then, for an arbitrary sequence $\{\mu_{n}\}_{n=1}^{\infty}\in G$ converging to $\mu\in G$, one has that $\limsup\limits_{n\rightarrow\infty}\operatorname{Val}_{n}(\mu_{n})\leq\operatorname{Val}(\mu).$
8 An example
We consider a time-optimal problem for the nonlocal continuity equation assuming that $d=1$, $U=[-1,0]$,
$f(x,m,u)=u-\int_{\mathbb{R}}ym(dy),$
while the target set is
$M=\Bigg{\{}m\in\mathcal{P}^{2}(\mathbb{R}):\ \int_{\mathbb{R}}ym(dy)=0\Bigg{\}}.$
Therefore, nonlocal continuity equation (3) takes the form:
$\partial_{t}m(t)+\mbox{div}\Bigg{(}(u(t)-\int_{\mathbb{R}}ym(t,dy))m(t)\Bigg{)}=0,\ m(0)=\mu\in\mathcal{P}^{2}(\mathbb{R}).$
The mean of the measure $m(t)$ defined by the rule:
$\bar{m}(t)\triangleq\int_{\mathbb{R}}ym(t,dy)$
obeys the following equation:
$\frac{d}{dt}\bar{m}(t)=u-\bar{m}(t),\ \ \bar{m}(0)=\bar{\mu}.$ (33)
Here, given a measure $\mu$, we put $\bar{\mu}\triangleq\int_{\mathbb{R}}y\mu(dy)$. Notice that the target set also is determined by the mean, i.e, the aim of the control is to find the first time $\tau$ such that $\bar{m}(\tau)=0$. Given an open-loop control $u(\cdot)$, we have that the corresponding evaluation of the mean is
[…]
Thus, the optimal control is as follows:
  1. 1.
    $u(\cdot)\equiv-1$ if $\bar{\mu}\geq 0$;
  2. 2.
    no optimal control if $\bar{\mu}<0$;
whereas the value function is
$\operatorname{Val}(\mu)=\left\{\begin{array}[]{ll}\ln(\bar{\mu}+1),&\bar{\mu}\geq 0,\\ +\infty,&\text{otherwise.}\end{array}\right.$
Notice that the value function is only lower semicontinuous.
Now, let us write down the Hamiltonian corresponding to the examined problem. It has the form:
$\begin{split}H(m,s)&=\inf\limits_{u\in[-1;0]}\int_{\mathbb{R}}s(x)\Bigl{(}u-\int_{\mathbb{R}}ym(dy)\Bigr{)}m(dx)\\ &=\begin{cases}\Bigl{(}-1-\bar{m}\Bigr{)}\int_{\mathbb{R}}s(x)m(dx),&\int_{\mathbb{R}}s(x)m(dx)>0,\\ -\bar{m}\int_{\mathbb{R}}s(x)m(dx),&\int_{\mathbb{R}}s(x)m(dx)\leq 0.\end{cases}\end{split}$
Thus, the Dirichlet problem is
$H(m,\nabla\phi(m))+1-\phi(m)=0,\ \ \int_{\mathbb{R}}ym(dy)\neq 0,\ \ \phi(m)=0,\text{ when }\int_{\mathbb{R}}ym(dy)=0.$
Due to the Kruzhkov transform of the function $\operatorname{Val}$, the viscosity solution is equal to
$\phi(\mu)=\begin{cases}1-e^{-\ln(1+\bar{\mu})},&\bar{\mu}\geq 0,\\ 1,&\bar{\mu}<0.\end{cases}$
9 Equivalent definitions of the viscosity solution
As we mentioned in the Introduction, there are several approaches to the notion of a viscosity solution of the Hamilton-Jacobi equation in the space of probability measures. In this paper, we do not touch the approaches relying on the $L^{2}$ representation and restrict our attention only to some concepts involving intrinsic sub- and superdifferentials. We consider two approaches coming back to papers [5] and [27].
9.1 Measure-varying viscosity solution
This approach relies on definitions proposed in [5]. There, the Hadamard lower derivative of a function $\psi:\mathbb{B}_{r}(m)\rightarrow\mathbb{R}$ at $m$ in the direction $F$ is defined by the rule:
$\tilde{d}_{H}^{-}\psi(m;F)\triangleq\liminf_{h\downarrow 0,W_{2}(\mu,(\operatorname{Id}+hF)=o(h)}\frac{1}{h}[\psi(\mu)-\psi(m)].$
The corresponding subdifferential is defined by the rule
~$\tilde{\partial}^{-}\psi(m)\triangleq\Big{\{}s\in L^{2}(m):\,\langle s,F\rangle_{m}\leq\tilde{d}^{-}_{H}\psi(m;F)\text{ for every }F\in L^{2}(m)\Big{\}}.$
Below, we will call these subdifferentials a measure-varying subdifferential (MV-subdifferential for shortness). Notice that
$\tilde{d}^{-}_{H}\psi(m;F)\leq d^{-}_{H}\psi(m;F).$
Thus,
~$\tilde{\partial}^{-}\psi(m)\subset\partial^{-}\psi(m).$ (34)
One can also define the MV-superdifferential letting
~$\tilde{\partial}^{+}\psi(m)\triangleq-\tilde{\partial}^{-}(-\psi)(m).$
Analogously,
~$\tilde{\partial}^{+}\psi(m)\subset\partial^{+}\psi(m).$ (35)
Notice that, for every $\psi_{1},\psi_{2}:\mathbb{B}_{r}(m)\rightarrow\mathbb{R}$, one has that
~$\tilde{\partial}^{-}\psi_{1}(m)+\tilde{\partial}^{-}\psi_{2}(m)\subset\tilde{\partial}^{-}(\psi_{1}+\psi_{2})(m).$ (36)
Indeed, given $s_{1}\in\tilde{\partial}^{-}\psi_{1}(m)$, $s_{2}\in\tilde{\partial}^{-}\psi_{2}(m)$, $F\in L_{2}(m)$, let $\{h_{k}\}_{k=1}^{\infty}\subset(0,+\infty)$ converge to zero and let $\{\mu_{k}^{\prime}\}_{k=1}^{\infty}\subset\mathbb{B}_{r}(m)$ be such that $W_{2}(\mu_{k}^{\prime},(\operatorname{Id}+h_{k}F)\sharp m)=o(h_{k})$ while
$\tilde{d}^{-}_{H}(\psi_{1}+\psi_{2})(m;F)=\lim_{k\rightarrow\infty}\frac{1}{h_{k}}[(\psi_{1}(\mu_{k}^{\prime})+\psi_{2}(\mu_{k}^{\prime}))-(\psi_{1}(m)-\psi_{2}(m))].$
By construction, for $i=1,2$,
$\langle s_{i},F\rangle_{m}\leq\liminf_{k\rightarrow\infty}\frac{1}{h_{k}}[\psi_{i}(\mu_{k}^{\prime})-\psi_{i}(m)].$
Threfore,
$\langle s_{1}+s_{2},F\rangle_{m}\leq\tilde{d}^{-}_{H}(\psi_{1}+\psi_{2})(m).$
Since $F$ was an arbitrary element of $L_{2}(m)$, $s_{1}+s_{2}\in\tilde{\partial}(\psi_{1}+\psi_{2})(m)$.
Analogously,
~$\tilde{\partial}^{+}\psi_{1}(m)+\tilde{\partial}^{+}\psi_{2}(m)\subset\tilde{\partial}^{+}(\psi_{1}+\psi_{2})(m).$
Replacing the sub- and superdifferentials in the definitions of super- and subsolutions with MV-sup- and MV-superdifferentials, we arrive at the concepts of MV-viscosity supersolutions, subsolutions and MV-viscosity solutions of problem (8).
Proposition 9.1.
Each viscosity solution of (8) is a MV-viscosity solution. Moreover, the $MV$-viscosity solution of (8) is unique.
Доказательство.
The fact that each viscosity solution of (8) is a MV-viscosity solution directly follows from (34) and (35).
To show the uniqueness of the MV-viscosity solution, one should prove the comparison principle for MV-viscosity sub- and supersolution. It mimics the proof of Theorem 5.1 with two differences.
First, we are to prove that $\mathbf{0}(\cdot)\in\tilde{\partial}^{-}\Psi_{1}(m_{1}^{*})$. To this end, we choose $F\in L^{2}(m_{1}^{*})$, $h>0$ and $m_{1}^{\prime}\in G$ such that $W_{2}((\operatorname{Id}+hF)\sharp m_{1}^{*},m_{1}^{\prime})=o(h)$. We have that
$\begin{split}\Psi_{1}(m_{1}^{\prime})-\Psi_{1}(m_{1}^{*})&=-\Phi(m_{1}^{\prime},m_{2}^{*})+\Phi(m_{1}^{*},m_{2}^{*})\\ &\geq-\rho((\operatorname{Id}+hF)\sharp m_{1}^{*},m_{2}^{*};m_{1}^{*},m_{2}^{*})\\ &=-(4L-1)W_{2}^{2}(m_{1}^{\prime},m_{1}^{*})\\ &\geq-(4L-1)[h^{2}\|F\|^{2}_{L_{2}(m_{1}^{*})}+W_{2}(m_{1}^{\prime},(\operatorname{Id}+hF)\sharp m_{1}^{*}).\end{split}$
Thus, $\tilde{d}_{H}^{-}\phi(m;F)\geq 0$ for every $F\in L^{2}(m)$. This as in the proof of Theorem 5.1 gives that $\mathbf{0}(\cdot)\in\tilde{\partial}\Psi_{1}(m_{1}^{*})$.
Second difference is to show that the functions $s_{1}$ and $s_{2}$ defined by (17) lie in ~$\tilde{\partial}^{+}w_{1}(m_{1}^{*})$ and ~$\tilde{\partial}^{+}w_{2}(m_{2}^{*})$ respectively (here the functions $w_{1}$ and $w_{2}$ are given by (18)). This directly follows from the definition of the MV-superdifferential and (12). ∎
9.2 Plan determined optimal displacement viscosity solution
Notice that the definition of the sub-/superdifferentials used in the paper and the MV-sub- superdifferentails rely directions given by a square integrable functions. One can also consider more general directions those are determined by plans. This approach is natural for mean field type optimal control problems where the agents are affected by individual controls. The study of a viscosity solution for Hamilton-Jacobi equations corresponding to finite horizon mean filed type control problems is presented in [27]. We also refer to papers [19, 18] where the viscosity solutions for Dirichlet problem corresponding to time-optimal mean field type control problem were introduced. Additionally, the approach of [19, 18, 27] implies that the elements of the subdifferential are optimal displacements. Recall that a function $s\in L^{2}(m)$ is an optimal displacement from $m$ provided that $s=\operatorname{Id}-T$, where $T$ is an optimal transportion map between $m$ and $T\sharp m$. We denote the set of all optimal displacement from $m$ by $\operatorname{Dis}(m)$. Equivalent definitions of optimal displacement are derived in [27, Lemmas 4 and 5]. Now let us introduce the concept of sub- and superdifferentials given in [27]. We will call them PDOD-sub- and superdifferentials due to the fact that this concept uses directions determined by plans and optimal displacements. Let $\phi:\mathbb{B}_{r}(m)\rightarrow\mathbb{R}$.
Definition 9.2.
Let $\delta\geq 0$, a $\delta$-PDOD-subdifferential is a set ^$\hat{\partial}^{-}_{\delta}\phi(m)$ consisting of all $s\in L^{2}(m)$ such that
  • $-s\in\operatorname{Dis}(m)$;
  • for every $\mu\in\mathbb{B}_{r}(m),$ and $\pi\in\Pi(m,\mu)$, one has that
    $\begin{split}\phi(\mu)-\phi(m)\leq\int_{\mathbb{R}^{d}}\langle s(x),y-x\rangle\pi(d(x,y))+&\delta\bigg{[}\int_{\mathbb{R}^{d}}\|y-x\|^{2}\pi(d(x,y)\bigg{]}^{1/2}\\ +o\Bigg{(}&\bigg{[}\int_{\mathbb{R}^{d}}\|y-x\|^{2}\pi(d(x,y)\bigg{]}^{1/2}\Bigg{)}.\end{split}$
A $\delta$-PDOD-subdifferential is a set
^$\hat{\partial}^{+}_{\delta}\phi(m)\triangleq-\hat{\partial}^{-}_{\delta}(-\phi)(m).$
If $\delta=0$, then we will use words PDOD-subdifferentials and PDOD-superdifferentials.
Notice that
^$\hat{\partial}^{-}_{0}\phi(m)=\bigcap_{\delta>0}\hat{\partial}^{-}_{\delta}\phi(m),\ \ \hat{\partial}^{+}_{0}\phi(m)=\bigcap_{\delta>0}\hat{\partial}^{+}_{\delta}\phi(m).$ (37)
Moreover, the definitions 4.19.2 yield that
^$\hat{\partial}_{0}^{-}\phi(m)\subset\partial^{-}\phi(m).$ (38)
Similarly,
^$\hat{\partial}_{0}^{+}\phi(m)\subset\partial^{+}\phi(m).$ (39)
Using notions of $\delta$-PDOD-sub- and superdifferentials, one can define analogs of the viscosity solution as follows.
Definition 9.3.
A bounded lower semicontinuous function $\phi_{1}:\operatorname{cl}G\rightarrow\mathbb{R}$ is called a relaxed PDOD-viscosity supersolution of problem (8) provided that
  • for each $m\in\partial G$, $\phi_{1}(m)=0$;
  • for every $\delta>0$, $m\in G$ and $s\in\hat{\partial}^{-}_{\delta}\phi_{1}(m)$
    $H(m,s)+1-\phi_{1}(m)\leq C\delta,$
    where $C$ is a constant depending only on $H$ and $G$.
Definition 9.4.
A bounded upper semicontinuous function $\phi_{2}:\operatorname{cl}G\rightarrow\mathbb{R}$ is called a relaxed PDOD-viscosity subsolution of problem (8) if
  • for each $m\in\partial G$, $\phi_{2}(m)=0$;
  • $\phi_{2}$ is continuous on $\partial G$;
  • for every $\delta>0$, $m\in G$ and $s\in\partial^{+}\phi_{2}(m)$ one has that
    $H(m,s)+1-\phi_{2}(m)\geq-C\delta,$
    where $C$ is a constant depending only $H$ and $G$.
Definition 9.5.
A bounded function $\phi:\operatorname{cl}G\rightarrow\mathbb{R}$ is called a PDOD-viscosity solution if
  • $\phi$ is a PDOD-supersolution of (8);
  • there exists a sequence $\{\phi_{k}\}_{k=1}^{\infty}$, where, for each $k$, the function $\phi_{k}:\operatorname{cl}G\rightarrow\mathbb{R}$ is a PDOD-subsolution of (8), while
    $\phi(m)=\lim_{k\rightarrow\infty}\phi_{k}(m).$
Furthermore, we consider the limiting case where only pure PDOD-sub- and PDOD-superdifferentials are used. The replacement of words viscosity sub- and superdifferentials with the words PDOD-sub- and PDOD-superdifferentials leads to the notion of PDOD-viscosity solution. The link between the introduced definitions is as follows.
Proposition 9.6.
Each relaxed PDOD-viscosity solution is a PDOD-viscosity solution.
Доказательство.
The fact that each relaxed PDOD-viscosity subsolution is a PDOD-viscosity subsolution directly follows from equality (37) and the definitions of these notions. Analogously, every relaxed PDOD-viscosity supersolution is a PDOD-viscosity supersolution. These two statements prove the proposition. ∎
Now we discuss the link between PDOD-viscosity solutions and viscosity solutions in the sense of Definition 4.5.
Proposition 9.7.
Each viscosity solution in the sense of Definition 4.5 is a PDOD-viscosity solution. A PDOD-viscosity solution is unique.
Доказательство.
First notice that due to (38), one has that each viscosity subsolution is a PDOD-viscosity subsolution. Analogously, due to (39), we have that each viscosity supersolution is a PDOD-viscosity supersolution. This proves the first statement of the proposition.
To obtain the uniquenss of the PDOD-viscosity solution, it suffices to prove the comparison principle for PDOD-viscosity sub- and supersolutions. It mimics the proof of Theorem 5.1 with two differences those concern the sub- and superdifferentials. First, we are to show that $\mathbf{0}(\cdot)\in\hat{\partial}^{-}_{0}\Psi_{1}(m_{1}^{*})$. To this end, consider $m^{\prime}\in G$ and a plan $\pi\in\Pi(m^{\prime},m_{1}^{*})$ and replace (16) with the following estimate
$\begin{split}\Psi_{1}(m^{\prime})-&\Psi_{1}(m_{1}^{*})=-\Phi(m^{\prime},m_{2}^{*})+\Phi(m_{1}^{*},m_{2}^{*})\\ &\geq-\rho(m^{\prime},m_{2}^{*};m_{1}^{*},m_{2}^{*})=-(4L-1)W_{2}^{2}(m^{\prime},m_{1}^{*})\\ &\geq-(4L-1)\int_{(\mathbb{R}^{d})^{2}}\|x^{\prime}-x\|^{2}\pi(d(x^{\prime},x).\end{split}$
This and the fact that $\mathbf{0}(\cdot)$ is an optimal displacement yield the inclusion $\mathbf{0}(\cdot)\in\hat{\partial}^{-}_{0}\Psi_{1}(m_{1}^{*})$. Analogously, $\mathbf{0}(\cdot)\in\hat{\partial}^{+}_{0}\Psi_{2}(m_{2}^{*})$.
The second difference as in the case of MV-viscosity solution is that $s_{1}\in\hat{\partial}^{+}_{0}w_{1}(m_{1}^{*})$, $s_{2}\in\hat{\partial}_{0}^{-}w_{2}(m_{2}^{*})$. This directly follows from (12) and the fact that $s_{1}$ (respectively, $s_{2}$) is an optimal displacement from $m_{1}^{*}$ (respectively, from $m_{2}^{*})$. Furthermore, we use the definition of PDOD-sub and PDOD-superdifferentials and conclude that $-s_{1}\in\hat{\partial}_{0}^{-}\phi_{1}(m_{1}^{*})$, while $s_{2}\in\hat{\partial}_{0}^{+}\phi_{1}(m_{2}^{*})$. Inclusion 36 gives that $-s_{1}\in\hat{\partial}^{-}_{0}\phi_{1}(m_{1}^{*})$ and $s_{2}\in\hat{\partial}^{+}_{0}\phi_{2}(m_{1}^{*})$. ∎
Propositions 9.19.69.7 imply the following.
Theorem 9.8.
The notion of the viscosity solution presented in Definition 4.5 is equivalent to the notions of MV-, relaxed PDOD-, and PDOD-viscosity solutions.
Приложение A Properties of relaxed controls
Recall that, if $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$,
$\varsigma(\mu)\triangleq\Bigg{[}\int_{\mathbb{R}^{d}}\|x\|^{2}\mu(dx)\Bigg{]}^{1/2}.$
Furthermore, for given $T>0$, $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $\xi\in\mathcal{U}_{T}$ and $m(\cdot)\triangleq m(\cdot;\mu,\xi)$, we denote by $X(\cdot;y,\mu,\xi)$ a solution of the following initial value problem:
$\frac{d}{dt}x(t)=\int_{U}f(x(t),m(t),u)\xi(du|t),\ \ x(0)=y.$ (40)
Proposition A.1.
The following estimates hold true, for every $t\in[0,T]$, $y\in\mathbb{R}^{d}$ and some constants dependent only on $\varsigma(\mu)$ and $T$:
  1. 1.
    $\varsigma(m(t))\leq c_{0}$;
  2. 2.
    $W_{2}(m(t),\mu)\leq c_{1}t$;
  3. 3.
    $\|X(t;y,\mu,\xi)\|\leq c_{3}(1+\|y\|)$;
  4. 4.
    $\|X(t;y,\mu,\xi)-y\|\leq c_{4}(1+\|y\|)t$.
Доказательство.
The proof relies on the integral representation of (40)
[…]
and the upper bound on $f$ that directly follows from Hypothesis 3.2
$\|f(x,m,u)\|\leq C^{\prime}(1+\|x\|+\varsigma(m)).$ (42)
Here $C^{\prime}$ is a constant.
Using this, the fact that $\varsigma(m(t))=\|X(t;\cdot,\mu,\xi)\|_{L^{2}(\mu)}$ and the Minkowski inequalities, we conclude that
$\varsigma(m(t))\leq\varsigma(\mu)+C^{\prime}\int_{0}^{t}(1+2\varsigma(m(t^{\prime}))dt^{\prime}.$
This and the Gronwall’s inequality imply the first statement of the proposition.
To show the second statement, it suffices to notice that $W_{2}(m(t),\mu)\leq\|X(t;\cdot,\mu,\xi)-\operatorname{Id}\|_{L^{2}(\mu)}$, use estimate (42) and the first statement.
The third and the fourth statements directly follow from integral representation of solution (41), estimate (42), the first statement of this proposition and the Gronwall’s inequality. ∎
Below, we will prove that the set of averaged velocities over the time interval $[0,h]$ is precompact in $L^{2}$. This result is based on the following.
Lemma A.2.
Let
  • $\{\zeta_{k}\}_{k=1}^{\infty}\subset\mathcal{P}(U)$ narrowly converges to $\zeta\in\mathcal{P}(U)$;
  • a function $g:\mathbb{R}^{d}\times U\rightarrow\mathbb{R}^{d}$ be continuous and satisfies the sublinear growth conditions: $\|g(x,u)\|\leq C_{1}(1+\|x\|)$;
  • $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$;
  • for each $k$, $G_{k}\in L^{2}(\mu)$ be defined by the rule: $G_{k}(x)\triangleq\int_{U}g(x,u)\zeta_{k}(du)$;
  • $G\in L_{2}(\mu)$ be such that $G(x)\triangleq\int_{U}g(x,u)\zeta(du)$.
Then the sequence $\{G_{k}\}_{k=1}^{\infty}$ converges to $G$ in $L^{2}(\mu)$.
Доказательство.
Let $\varepsilon>0$. For each $R>0$, we have that
$\begin{split}\int_{\mathbb{R}^{d}}\|G_{k}(x)-G(x)\|^{2}\mu(dx)=\int_{\|x\|\leq R}\|G_{k}&(x)-G(x)\|^{2}\mu(dx)\\ +&\int_{\|x\|>R}\|G_{k}(x)-G(x)\|^{2}\mu(dx).\end{split}$ (43)
Using the definitions of the functions $G_{k}$, $G$ and the sublinear growth condition we have that
$\int_{\|x\|>R}\|G_{k}(x)-G(x)\|^{2}\mu(dx)\leq 8\int_{\|x\|>R}C_{1}^{2}(1+\|x\|^{2})\mu(dx).$ (44)
Since $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, one can choose $R$ such that
$8\int_{\|x\|>R}C_{1}^{2}(1+\|x\|^{2})\mu(dx)\leq\varepsilon.$
Since the function $g$ is continuous, there exists a finite set $\{y_{j}\}_{j=1}^{J}\subset\mathbb{B}_{R}(0)$ such that
$\max_{\|x\|\leq R}\min_{j=1,\ldots,J}\max_{u\in U}\|g(x,u)-g(y_{j},u)\|\leq\sqrt{\varepsilon}/3.$
Furthermore, we use the assumption that $\{\zeta_{k}\}_{k=1}^{\infty}$ narrowly converges to $\zeta$. Thus, there exists a natural number $K$ such that for every $k\geq K$ and $j=1,\ldots,J$,
$\Bigg{|}\int_{U}g(y_{j},u)\zeta_{k}(du)-\int_{U}g(y_{j},u)\zeta(du)\Bigg{|}\leq\sqrt{\varepsilon}/3.$
Moreover, for every $x$ satisfying $\|x\|\leq R$, there exists a number $j\in\{1,\ldots,J\}$ such that
$\max_{u\in U}\|g(x,u)-g(y_{j},u)\|\leq\sqrt{\varepsilon}/3.$
Hence, if $\|x\|\leq R$ and $k\geq K$, then
$\begin{split}|G_{k}(x)-G(x)|&=\Bigg{|}\int_{U}g(x,u)\zeta_{k}(du)-\int_{U}g(x,u)\zeta(du)\Bigg{|}\\ &\leq\Bigg{|}\int_{U}g(y_{j},u)\zeta_{k}(du)-\int_{U}g(y_{j},u)\zeta(du)\Bigg{|}+2\sqrt{\varepsilon}/3\leq\sqrt{\varepsilon}.\end{split}$ (45)
From (45) it directly follows that
$\int_{\|x\|\leq R}\|G_{k}(x)-G(x)\|^{2}\mu(dx)\leq\varepsilon.$
Evaluating the right-hand side of (43) using this estimate and (44), we have that, there exists a natural number $K$ such that, for every $k\geq K$, one has that
$\int_{\mathbb{R}^{d}}\|G_{k}(x)-G(x)\|^{2}\mu(dx)\leq 2\varepsilon.$
Proposition A.3.
Let
  • $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$;
  • $T>0$;
  • for each $h\in[0,T]$ $\xi_{h}\in\mathcal{U}_{h}$;
  • $F_{h}(y)\triangleq h^{-1}\int_{[0,h]\times U}f(X(t;y,\mu,\xi_{h}),m(t;\mu,\xi_{h}),u)\xi(d(t,u))$.
Then, there exists a sequence $\{h_{k}\}_{k=1}^{\infty}$ such that $h_{k}\rightarrow 0$, $\|F_{h_{k}}-F\|_{L^{2}(\mu)}\rightarrow 0$ as $k\rightarrow\infty$, where $F\in L^{2}(\mu)$ with $F(y)\triangleq\int_{U}f(y,\mu,u)\zeta(du)$ for some $\zeta\in\mathcal{P}(U)$. If, additionally, one can find a probability $\zeta_{0}\in\mathcal{P}(U)$ such that $\xi_{h}(\cdot|t)=\zeta_{0}$, then $\zeta=\zeta_{0}$ and $\|F_{h}-F\|_{L^{2}(\mu)}\rightarrow 0$ as $h\rightarrow 0$.
Доказательство.
Let $F^{\prime}_{h}(y)\triangleq h^{-1}\int_{[0,h]\times U}f(y,\mu,u)\xi_{h}(d(t,u))$. From Proposition A.1, we have that
$\|F^{\prime}_{h}-F_{h}\|\leq C_{2}h.$ (46)
Furthermore, we define the measure $\zeta_{h}\in\mathcal{P}(U)$ by the rule: for every Borel set $\Gamma\subset U$,
$\zeta_{h}(\Gamma)\triangleq h^{-1}\xi_{h}([0,h]\times\Gamma).$
Thus, $F^{\prime}_{h}(y)=\int_{U}f(y,\mu,u)\zeta_{h}(du)$. Since $U$ is compact, there exists a sequence $\{h_{k}\}_{k=1}^{\infty}$ converging to zero such that $\{\zeta_{h_{k}}\}_{k=1}^{\infty}$ narrowly converges to some $\zeta\in\mathcal{P}(U)$. Letting $F(y)\triangleq\int_{U}f(y,\mu,u)\zeta(du)$ and using Proposition A.2, we conclude that $\|F^{\prime}_{h_{k}}-F\|_{L^{2}(\mu)}\rightarrow 0$ as $k\rightarrow\infty$. This and (46) imply the convergence $\|F_{h_{k}}-F\|_{L^{2}(\mu)}\rightarrow 0$ as $k\rightarrow\infty$.
In the case where $\xi_{h}(\cdot|t)=\zeta_{0}$, we have that $F^{\prime}_{h}(y)=\int_{U}f(y,\mu,u)\zeta_{0}(du)=F(y).$ Thus, $F_{h}$ converges itself to $F$ in $L^{2}(\mu)$. ∎
Список литературы
  • [1] L. Ambrosio and W. Gangbo. Hamiltonian ODEs in the Wasserstein space of probability measures. Comm. Pure Appl. Math., 61(1):18–53, 2008.
  • [2] L. Ambrosio, N. Gigli, and G. Savaré. Gradient flows in metric spaces and in the space of probability measures. Lectures in Mathematics ETH Zürich. Birkhäuser, 2. ed edition, 2008.
  • [3] Y. Averboukh. Krasovskii-Subbotin approach to mean field type differential games. Dyn. Games Appl., 9:573–593, 2019.
  • [4] Y. Averboukh. Stability analysis of mean field type control system with major agent. COT, 2023:1–22, 2023.
  • [5] Z. Badreddine and H. Frankowska. Hamilton-Jacobi inequalities on a metric space. J. Differ. Equ., 271:1058–1091, 2021.
  • [6] M. Bardi and I. Capuzzo-Dolcetta. Optimal control and viscosity solutions of Hamilton-Jacobi-Bellman equations. Birkhäuser Boston, Boston, 1997.
  • [7] V. I. Bogachev. Measure theory. Vol. I and II. Berlin: Springer, 2007.
  • [8] M. Bongini, M. Fornasier, F. Rossi, and F. Solombrino. Mean-field Pontryagin maximum principle. J. Optim. Theory Appl., 175(1):1–38, 2017.
  • [9] B. Bonnet. A Pontryagin maximum principle in Wasserstein spaces for constrained optimal control problems. ESAIM: COCV, 25:38, 2019. Id/No 52.
  • [10] B. Bonnet and H. Frankowska. Necessary optimality conditions for optimal control problems in Wasserstein spaces. Appl. Math. Optim., 84:1281–1330, 2021.
  • [11] B. Bonnet and F. Rossi. The Pontryagin maximum principle in the Wasserstein space. Calc. Var. Partial Differ. Equ., 58(1):36, 2019. Id/No 11.
  • [12] A. Braides. Gamma-convergence for beginners. Oxford University Press, 2002.
  • [13] R. W. Brockett. Notes on the control of the Liouville equation. In Control of Partial Differential Equations, pages 101–129, 2012.
  • [14] P. Cardaliaguet, F. Delarue, J.-M. Lasry, and P.-L. Lions. The master equation and the convergence problem in mean field games, 2019.
  • [15] P. Cardaliaguet and M. Quincampoix. Deterministic differential games under probability knowledge of initial condition. Int. Game Theory Rev., 10(01):1–16, 2008.
  • [16] J. A. Carrillo, M. Fornasier, G. Toscani, and F. Vecil. Particle, kinetic, and hydrodynamic models of swarming. In Mathematical modeling of collective behavior in socio-economic and life sciences, pages 297–336. Boston, MA: Birkhäuser, 2010.
  • [17] G. Cavagnari. Regularity results for a time-optimal control problem in the space of probability measures. Math. Control Relat. Fields, 7(2):213–233, 2017.
  • [18] G. Cavagnari and A. Marigonda. Time-pptimal control problem in the space of probability measures, 2015.
  • [19] G. Cavagnari and A. Marigonda. Attainability property for a probabilistic target in Wasserstein spaces. Discrete Continuous Dyn. Syst. - A, 41(2):777–812, 2021.
  • [20] G. Cavagnari, A. Marigonda, and M. Quincampoix. Compatibility of state constraints and dynamics for multiagent control systems. J. Evol. Equ., 21(4):4491–4537, 2021.
  • [21] R. M. Colombo, M. Garavello, M. Lécureux-Mercier, and N. Pogodaev. Conservation laws in the modeling of moving crowds. Preprint, arXiv:1211.0408 [math.AP] (2012), 2012.
  • [22] M. G. Crandall and P.-L. Lions. Viscosity solutions of Hamilton-Jacobi equations. Trans. Amer. Math. Soc., 277(1):1, 1983.
  • [23] C. Dogbe. On the modelling of crowd dynamics by generalized kinetic models. J. Math. Anal. Appl., 387(2):512–532, 2012.
  • [24] R. C. Fetecau, Y. Huang, and T. Kolokolnikov. Swarm dynamics and equilibria for a nonlocal aggregation model. Nonlinearity, 24(10):2681–2716, 2011.
  • [25] W. Gangbo and A. Tudorascu. On differentiability in the Wasserstein space and well-posedness for Hamilton-Jacobi equations. J. Math. Pures Appl., 125:119–174, 2019.
  • [26] R. L. Hughes. A continuum theory for the flow of pedestrians. Transp. Res. Part B, 36(6):507–535, 2002.
  • [27] C. Jimenez, A. Marigonda, and M. Quincampoix. Optimal control of multiagent systems in the Wasserstein space. Calc. Var. Partial Differ. Equ., 59(2):45, 2020. Id/No 58.
  • [28] C. Jimenez, A. Marigonda, and M. Quincampoix. Dynamical systems and Hamilton–Jacobi–Bellman equations on the Wasserstein space and their l<sup>2</sup> representations. SIAM J. Math. Anal., 55(5):5919–5966, 2023.
  • [29] M. Krastanov and M. Quincampoix. Local small time controllability and attainability of a set for nonlinear control system. ESAIM Control Optim. Calc. Var., 6:499–516, 2001.
  • [30] A. Kruger, S. Plubtieng, and T. Seangwattana. Borwein–Preiss variational principle revisited. J. Math. Anal. Appl., 435:1183–1193, 2016.
  • [31] A. Mogilner and L. Edelstein-Keshet. A non-local model for a swarm. J. Math. Biol., 38(6):534–570, 1999.
  • [32] G. Naldi, L. Pareschi, and G. Toscani, editors. Mathematical modeling of collective behavior in socio-economic and life sciences. Model. Simul. Sci. Eng. Technol. Boston, MA: Birkhäuser, 2010.
  • [33] H. Pham and X. Wei. Bellman equation and viscosity solutions for mean-field stochastic control problem. ESAIM Control Optim. Calc. Var., 24(1):437–461, 2018.
  • [34] B. Piccoli and F. Rossi. Measure-theoretic models for crowd dynamics. In Crowd dynamics, Volume 1. Theory, models, and safety problems, pages 137–165. Cham: Birkhäuser, 2018.
  • [35] N. Pogodaev. Optimal control of continuity equations. Nonlinear Differ. Equ. Appl., 23(2):24, 2016. Id/No 21.
  • [36] M. H. Rinaldo M. Colombo and M. Mercier. Control of the continuity equation with a non local flow. ESAIM Control Optim. Calc. Var., 17(2):353–379, 2011.
  • [37] A. I. Subbotin. Generalized solutions of first-order PDEs: The dynamical pptimization perspective. Birkhäuser Boston, Boston, 1995.
  • [38] B.-C. Wang and Y. Liang. Robust mean field social control problems with applications in analysis of opinion dynamics. Int. J. Control, 95(12):3309–3325, 2022.