### v2

# arXiv:2502.20871v2[math.AP] 28 Feb 2026

arXiv:2502.20871v2[math.AP] 28 Feb 2026
Time-optimal problem in the space of probability measures
Yurii Averboukh HSE University, Moscow, Russia averboukh@gmail.com    Ekaterina Kolpakova Krasovskii Institute of Mathematics and Mechanics,
Yekaterinburg, Russia
eakolpakova@gmail.com
Аннотация
This paper focuses on the value function in the time-optimal problem for a continuity equation in the space of probability measures. We derive the dynamic programming principle for this problem. In particular, we prove that the Kruzhkov transform of the value function is a unique discontinuous viscosity solution to the corresponding Dirichlet problem for the Hamilton–Jacobi equation. Finally, we establish the $\Gamma$-convergence of the value function in a perturbed problem to the value function in the unperturbed problem.
MSC Classification (2020): 49J20, 49J52, 49L25, 49N70, 91A23, 49K20, 82C22.
keywords:
time-optimal problem, Hamilton-Jacobi equation, controlled continuity equation, viscosity solution, Dirichlet problem, nonsmooth analysis in the Wasserstein space
1 Introduction
In this paper, we study a control system in the space of probability measures, where the dynamics is governed by a nonlocal continuity equation driven by an external force. This equation models the evolution of a system comprising infinitely many identical particles, with each particle influenced by the collective field generated by all other particles.
The nonlocal continuity equation traces its origins to the pioneering work of Vlasov on plasma dynamics. Nowadays, it finds widespread application in various fields, including opinion dynamics models [16, 37, 42], crowd behavior analysis [21, 27, 31, 39] and swarm dynamics [28, 36].
The objective of the control is to steer the system to a target in minimal time. The target here is assumed to be an arbitrary closed set of probability measures. In this work, we develop the dynamic programming approach and investigate the properties of the value function.
The main results of the paper are as follows. First, we establish the existence of an optimal relaxed control. Next, we derive the dynamic programming principle for the value function in this setting. We then develop a framework based on the Dirichlet problem for the Hamilton-Jacobi equation, which, in this context, is a partial differential equation in the space of measures. We demonstrate that the Kruzhkov transform of the value function satisfies the aforementioned Dirichlet problem in the viscosity sense. Furthermore, by deriving a comparison principle, we prove the existence and uniqueness of a solution to the Dirichlet problem associated with the time-optimal problem for a nonlocal continuity equation. In general, the value function is discontinuous; however, we identify the sufficient condition that ensures its continuity. Finally, we show that the values of the perturbed time-optimal problems $\Gamma$-converge to the solution of the limiting problem.
Now, let us give a brief overview of the existing literature and compare our results with what is already known. The study of control problems for the nonlocal continuity equation began with papers [13, 22]. Necessary optimality conditions in the form of the Pontryagin maximum principle for finite-horizon problems were derived in [8, 9, 10, 11, 40]. We should also highlight paper [5], where the authors develop the dynamic programming principle for the optimal control problem of the continuity equation on a finite time interval. In that work, it is shown that the value function restricted to the space of compactly supported measures is a solution to the Cauchy problem for the Hamilton-Jacobi equation in the space of probability measures. Moreover, the uniqueness of the solution is also established on the set of compactly supported probability measures.
Notice that the problem studied in this paper assumes the system is influenced by an external force. Consequently, the velocity field is a continuous function of a point. An alternative approach considers the possibility of a discontinuous velocity field. This framework known as the mean field type control theory is particularly relevant for systems involving intellectual agents striving to achieve a common goal. For the current state of research in this direction, we refer to the works of [3, 33, 34]. Time-optimal problems for mean field type control systems were investigated in [19, 17, 18], where the target is assumed to be a hyperplane in the space of probability measures. In those papers, the value function was characterized as a solution to a Dirichlet problem for the Hamilton-Jacobi equation in the space of measures. Furthermore, sufficient conditions for the regularity of the value function were established. However, these works do not provide a uniqueness result.
As previously mentioned, our approach to analyzing time-optimal problems relies on viscosity solutions to the Hamilton–Jacobi equation. This methodology was initially developed for finite-dimensional equations (see [6, 24, 41]).
The concept of viscosity solutions has been extended in recent years to spaces of probability measures (see [5, 19, 18, 23, 29, 26, 33, 38]). The definition of a viscosity solution inherently relies on the notions of subdifferentials and superdifferentials, which can themselves be introduced in several ways in the context of probability measures. The central question here concerns the choice of admissible variations and the corresponding definition of sub- and superdifferentials. In our framework, admissible variations are Borel functions satisfying a sublinear growth condition. Consequently, elements of the sub- and superdifferentials belong to the dual class; specifically, they are functions whose product with any sublinearly growing variation is integrable. In general, these sub- and superdifferentials are larger than those in the approach originally proposed in [15], where both admissible variations and elements of the sub-/superdifferentials are required to be square-integrable. Consequently, our definition of viscosity solution is slightly stronger than the one based solely on square-integrable variations. Nevertheless, the nonsmooth analysis developed here is particularly well-suited to the Dirichlet problem arising from time-optimal control of a nonlocal continuity equation. In particular, our proof of the comparison principle (Theorem 5.1) essentially uses the fact that variations are only assumed to satisfy a sublinear growth condition.
We now briefly review alternative approaches to defining sub-/superdifferentials and viscosity solutions for Hamilton-Jacobi equations in the Wasserstein space. One approach, employed in [5] for finite-horizon optimal control problems with nonlocal continuity equations (see also [15]), considers variations of a measure along directions given by square-integrable functions. Another line of work uses directions defined by transport plans (see [1, 18, 29, 33]); this is primarily applied to Hamilton-Jacobi equations associated with deterministic mean-field control problems. In [2], elements of subdifferentials are even allowed to be distributions. A recent contribution [26] introduces strong viscosity solutions for semilinear second-order Hamilton-Jacobi equations by leveraging entropy regularization.
All the approaches mentioned above are intrinsic, as they rely only on measures and directions within the space of measures. Alternatively, a probability measure with finite second moment can be represented as the push-forward of a square-integrable random variable, leading to an external approach to nonsmooth analysis in the space of probability measures (see [14, 29, 38, 23]). Connections between the external viewpoint and certain intrinsic definitions of sub- and superdifferentials are discussed in [14, 29, 32, 34]. A comparative analysis of various intrinsic approaches can also be found in [33, 34].
The paper is organized as follows. In Section 2, we introduce the general notation used throughout the paper. The formulation of the time-optimal control problem for the nonlocal continuity equation is presented in Section 3, where we also prove the existence of an optimal control and show that the value function is lower semicontinuous. Section 4 introduces the Dirichlet problem for the Hamilton-Jacobi equation and defines the corresponding notion of a viscosity solution. The comparison principle for this problem is established in Section 5, which, in turn, implies the uniqueness of the viscosity solution. In Section 6, we prove that the Kruzhkov transform of the value function is a viscosity solution to the aforementioned Dirichlet problem. Additionally, we derive a sufficient condition ensuring the continuity of the value function and show that the values of perturbed time-optimal control problems $\Gamma$-converge to the original problem (see Section 7). Finally, Section 8 provides an example of a time-optimal control problem for the controlled continuity equation.
2 Preliminaries
If $X_{1},\ldots,X_{n}$ are sets, $i_{1},\ldots,i_{k}\in\{1,\ldots,n\}$, then $\operatorname{p}^{i_{1},\ldots,i_{k}}$ denotes a natural projection of $X_{1}\times\ldots\times X_{n}$ onto $X_{i_{1}}\times\ldots\times X_{i_{k}}$. Additionally, $\operatorname{Id}$ stands for the identity mapping.
If $(\Omega^{\prime},\mathcal{F}^{\prime})$ and $(\Omega^{\prime\prime},\mathcal{F}^{\prime\prime})$ are measurable spaces, $m$ is a measure on $\mathcal{F}^{\prime}$, and $h:\Omega^{\prime}\rightarrow\Omega^{\prime\prime}$ is a $\mathcal{F}^{\prime}/\mathcal{F}^{\prime\prime}$-measurable map, then we denote by $h\sharp m$ the push-forward measure defined by the rule: for each $\Gamma\in\mathcal{F}^{\prime\prime}$,
$(h\sharp m)(\Gamma)=m(h^{-1}(\Gamma)).$
If $(X,\rho_{X})$ is a Polish space, $x\in X$, $r>0$, then $\mathbb{B}_{r}(x)$ stands for the ball of radius $r$ centered at $x$.
Furthermore, $\mathcal{M}(X)$ denotes the space of positive finite Borel measures on $X$, while $\mathcal{P}(X)$ stands for the space of Borel probabilities on $X$, i.e., $\mathcal{P}(X)=\{m\in\mathcal{M}(X):\,m(X)=1\}$.
We consider the narrow convergence on $\mathcal{M}(X)$. This means that $\{m_{n}\}_{n=1}^{\infty}\subset\mathcal{M}(X)$ converges to $m\in\mathcal{M}(X)$ if, for every $\phi\in C_{b}(X)$,
$\int_{X}\phi(x)m_{n}(dx)\rightarrow\int_{X}\phi(x)m(dx)\text{ as }n\rightarrow\infty.$
Obviously, $\mathcal{P}(X)$ is closed w.r.t. the narrow convergence.
We denote the space of probability measures with the finite second moment by $\mathcal{P}^{2}(X)$, i.e., $m\in\mathcal{P}(X)$ belongs to $\mathcal{P}^{2}(X)$ provided that, for some $x_{*}\in X$, $\int_{X}\rho_{X}^{2}(x,x_{*})m(dx)<\infty$. If $X$ is a Banach space, then we denote
$\varsigma(m)\triangleq\Bigg[\int_{X}\|x\|^{2}m(dx)\Bigg]^{1/2}.$
We endow $\mathcal{P}^{2}(X)$ with the second Wasserstein distance defined by the rule:
$W_{2}(m_{1},m_{2})=\Bigg[\inf\limits_{\pi\in\Pi(m_{1},m_{2})}\int_{X\times X}\big(\rho_{X}(x_{1},x_{2})\big)^{2}\pi(d(x_{1},x_{2}))\Bigg]^{1/2},$
where $\Pi(m_{1},m_{2})$ stands for the set of probabilities $\pi\in\mathcal{P}^{2}(X\times X)$ such that $\operatorname{p}^{i}\sharp\pi=m_{i}$.
If, additionally, $Y$ is a Banach space, $m$ is a measure on $X$ and $p\geq 1$, then $L^{p}(X,m;Y)$ denotes the space of functions $\varphi:X\mapsto Y$ such that
$\int_{X}\|\varphi(x)\|^{p}_{Y}m(dx)<\infty.$
If $X=Y=\mathbb{R}^{d}$, $p=2$, we will shorten the notation and write $L^{2}(m)$ instead of $L^{2}(\mathbb{R}^{d},m;\mathbb{R}^{d})$. The norm on $L^{2}(m)$ is denoted by $\|\cdot\|_{L^{2}(m)}$. Additionally, $L^{0}(X;Y)$ denotes the set of all measurable functions from $X$ to $Y$.
If $s_{1},s_{2}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ are such that the mapping $x\mapsto\langle s_{1}(x),s_{2}(x)\rangle$ lies in $L^{1}(m)$, then we put
$\langle s_{1},s_{2}\rangle_{m}\triangleq\int_{\mathbb{R}^{d}}\langle s_{1}(x),s_{2}(x)\rangle m(dx).$
Notice that $\langle\cdot,\cdot\rangle_{m}$ is the inner product on $L^{2}(m)$.
If $m\in\mathcal{P}^{2}(\mathbb{R}^{d})$, then $\operatorname{SL}(m)$ stands for the set of Borel functions satisfying the sublinear growth condition, i.e., a Borel function $F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ lies in $\operatorname{SL}(m)$ if, for some constant $C>0$, and $m$-a.e. $x\in\mathbb{R}^{d}$,
$\|F(x)\|\leq C(1+\|x\|).$
Notice that $\operatorname{SL}(m)\subset L^{2}(m)$. We define the convergence on $\operatorname{SL}(m)$ in the following way: a sequence of functions $\{F_{n}\}_{n=1}^{\infty}\subset\operatorname{SL}(m)$ converges to $F\in\operatorname{SL}(m)$ provided that
  • there exists a constant $C$ such that
    $\|F_{n}(x)\|\leq C(1+\|x\|)\text{ for }m\text{-a.e. }x\in\mathbb{R}^{d};$
  • $F_{n}(x)\rightarrow F(x)$ as $n\rightarrow\infty$ for $m$-a.e. $x\in\mathbb{R}^{d}$.
An analogous definition is used when we consider the convergence of a family of functions $\{F_{h}\}_{h\in(0,\epsilon)}$.
The symbol $\operatorname{LD}(m)$ denotes the set of all functions $s:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ such that
$\int_{\mathbb{R}^{d}}\|s(x)\|(1+\|x\|)m(dx)<\infty.$
One can regard $\operatorname{LD}(m)$ as the dual space to $\operatorname{SL}(m)$. In particular, for every $F\in\operatorname{SL}(m)$, $s\in\operatorname{LD}(m)$, $\langle s,F\rangle_{m}$ is well-defined. Moreover, if $\{F_{n}\}_{n=1}^{\infty}\subset\operatorname{SL}(m)$ converges to some $F\in\operatorname{SL}(m)$, then $\langle s,F_{n}\rangle_{m}\rightarrow\langle s,F\rangle_{m}$.
Now let $(X,\rho_{X})$, $(Y,\rho_{Y})$ be Polish spaces. Assume that $m$ is a finite measure on $X$, while $\alpha\in\mathcal{M}(X\times Y)$ is such that $\operatorname{p}^{1}\sharp\alpha=m$. Notice that, due to the disintegration theorem [7, Theorem 10.4.14], one can find a family of probability measures $\{\alpha(\cdot|x)\}_{x\in X}\subset\mathcal{P}(Y)$ such that, for every function $\phi\in L^{1}(X\times Y,\alpha;\mathbb{R})$,
$\int_{X\times Y}\phi(x,y)\alpha(d(x,y))=\int_{X}\int_{Y}\phi(x,y)\alpha(dy|x)m(dx).$ (1)
Moreover, this family is $m$-a.e. unique. If $\{\alpha(\cdot|x)\}_{x\in X}$ is a family of probabilities on $Y$ that is weakly measurable, one can uniquely construct a measure $\alpha\in\mathcal{M}(X\times Y)$ such that $\operatorname{p}^{1}\sharp\alpha=m$ and (1) holds true.
In what follows, given a compact metric space $U$, we will consider measures on $0,+\infty)\times U$ whose marginal distributions on $0,+\infty)$ coincide with the Lebesgue measure. The set of such measures is denoted by $\mathcal{U}$. Elements of $\mathcal{U}$ are regarded as relaxed (generalized) controls. If $\xi\in\mathcal{U}$, $T>0$, then $\xi|_{T}$ stands for its restriction on $0,T]\times U$, i.e., $\xi|_{T}$ is a measure on $0,T]\times U$ such that its marginal distribution on $[0,T]$ is the Lebesgue measure. We denote the set of such measures by $\mathcal{U}_{T}$. The set $\mathcal{U}_{T}$ is closed w.r.t. the narrow convergence.
Definition 2.1.
We say that a sequence $\{\xi_{n}\}_{n=1}^{\infty}\subset\mathcal{U}$ converges to $\xi\in\mathcal{U}$ provided that, for each $T>0$, $\{\xi_{n}|_{T}\}_{n=1}^{\infty}$ converges to $\{\xi|_{T}\}$ narrowly.
Notice that in this definition it suffices to consider only natural numbers $T$. Consequently, the space $\mathcal{U}$ is compact. Moreover, each measurable function $u:[0,+\infty)\rightarrow U$ generates an element of $\mathcal{U}$ such that $\xi(\cdot|t)\triangleq\delta_{u(t)}(\cdot)$. The set of such elements is dense in $\mathcal{U}$ [43, Theorem IV.3.10].
3 Statement of the problem
The central object of this work is the controlled nonlocal continuity equation given by:
$\partial_{t}m(t)+\operatorname{div}(f(x,m(t),u(t))m(t))=0,\ \ t\in[0,+\infty).$ (2)
Here, $x\in\mathbb{R}^{d}$, $m(t)\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $u(t)\in U$, the set $U$ is a metric compact. We interpret $u(t)$ as an instantaneous control. Thus, $U$ is the space of admissible controls.
Equation (2) describes the evolution of the distribution of infinitely many similar particles obeying the following equation
$\frac{d}{dt}x(t)=f(x(t),m(t),u(t)),\ x\in\mathbb{R}^{d},\ u(t)\in U.$
Here, the distribution of agents at the instant $t$ is denoted $m(t)\in\mathcal{P}^{2}(\mathbb{R}^{d})$, while $u(t)$ stands for an external control that influences the whole system.
We impose the following assumptions.
Hypothesis 3.1.
The function $f$ is continuous in all arguments.
Hypothesis 3.2.
There exists a constant $L>0$ such that, for every $x_{1},x_{2}\in\mathbb{R}^{d}$, $m_{1},m_{2}\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $u\in U$, one has the inequality
$\|f(x_{1},m_{1},u)-f(x_{2},m_{2},u)\|\leq L(\|x_{1}-x_{2}\|+W_{2}(m_{1},m_{2})).$
Hypothesis 3.3.
There exists a constant $C_{f}>0$ such that, for every $x\in\mathbb{R}^{d}$, $m\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $u\in U$, one has the inequality
$\|f(x,m,u)\|\leq C_{f}\Bigg(\mathscr{a}(x)+\int_{\mathbb{R}^{d}}\mathscr{a}(x)m(dx)\Bigg),$
where
$\mathscr{a}(x)\triangleq(1+\|x\|^{2})^{1/2}.$ (3)
In the paper, we use relaxed controls. As mentioned above, each relaxed control is a measure on $0,+\infty)\times U$ with the marginal distribution on $0,+\infty)$ coinciding with the Lebesgue measure. Furthermore, given $\xi\in\mathcal{U}$, there exists its disintegration w.r.t. the Lebesgue measure denoted by $\xi(\cdot|t)$. Formally, the usage of relaxed controls means that we replace continuity equation (2) with
$\partial_{t}m(t)+\operatorname{div}\Bigg(\int_{U}f(x,m(t),u)\xi(du|t)\cdot m(t)\Bigg)=0,\ \ t\in[0,+\infty).$ (4)
We will consider this equation in the distributional sense.
Definition 3.4.
Let $T>0$, $\xi\in\mathcal{U}_{T}$. A measure-valued function $m(\cdot):[0,T]\rightarrow\mathcal{P}^{2}(\mathbb{R}^{d})$ is called a solution of equation (4) on $[0,T]$ if, for every $\varphi\in C_{0}^{1}((0,T)\times\mathbb{R}^{d})$, it holds that
$\int_{[0,T]\times U}\int_{\mathbb{R}^{d}}\partial_{t}\varphi(t,x)+\left\langle\nabla\varphi(t,x),f(x,m(t),u)\right\rangle m(t,dx)\xi(d(t,u))=0.$
If $\xi\in\mathcal{U}$, then a measure-valued function $m(\cdot):[0,+\infty)\rightarrow\mathcal{P}^{2}(\mathbb{R}^{d})$ is a solution of (4) on $0,+\infty)$ if, for every $T>0$, the restriction of $m(\cdot)$ on $[0,T]$ is a solution of (4) on $[0,T]$.
Given $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $\xi\in\mathcal{U}$, we denote by $m(\cdot;\mu,\xi)$ the unique solution of (4) on $0,+\infty)$ satisfying $m(0)=\mu$. If $\xi\in\mathcal{U}_{T}$, we will, by a slight abuse of notation, also denote the solution of (4) on $[0,T]$ satisfying $m(0)=\mu$ by $m(\cdot;\mu,\xi)$.
An alternative definition of the solution to continuity equation (4) relies on the particle interpretation. To introduce it, for given $y\in\mathbb{R}^{d}$, a continuous function $m(\cdot):[0,T]\rightarrow\mathcal{P}^{2}(\mathbb{R}^{d})$, $\xi\in\mathcal{U}_{T}$, we denote by $X(\cdot;y,m(\cdot),\xi)$ the solution of the initial value problem
$\frac{d}{dt}x(t)=\int_{U}f(x(t),m(t),u)\xi(du|t),\ x(0)=y.$
The function $X(\cdot;y,m(\cdot),\xi)$ describes the motion of a particle on the time interval $[0,T]$ in the case where the distribution of all particles is given by $m(\cdot)$, while the system is affected by a relaxed control $\xi$. The motion on $0,+\infty)$ can be defined in the same way.
From the superposition principle [2, Theorem 8.2.1], a measure-valued function $m(\cdot):[0,T]\rightarrow\mathcal{P}^{2}(\mathbb{R}^{d})$ solves (4) for some $\xi\in\mathcal{U}$ if and only if
$m(t)=X(t;\cdot,m(\cdot),\xi)\sharp m(0).$
Now, let us describe the time-optimal problem examined in the paper. We assume that we are provided with a closed set $M\subset\mathcal{P}^{2}(\mathbb{R}^{d})$. Furthermore, we denote $G\triangleq\mathcal{P}^{2}(\mathbb{R}^{d})\setminus M$. Additionally, $\partial G$ stands for the border of $G$. Let us consider the functional $\tau$ defined on $C([0,+\infty);\mathcal{P}^{2}(\mathbb{R}^{d}))$ such that
$\tau(m(\cdot))\triangleq\inf\{t\in[0;+\infty):\ m(t)\in M\}.$ (5)
Notice that, if, for every $t>0$, $m(t)\not\in M$, then $\tau(m(\cdot))=+\infty$. The time-optimal problem means that, given an initial distribution $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, one wishes to minimize the quantity $\tau(m(\cdot;\mu,\xi))$. Thus, the value function of time-optimal problem (2), (5) is defined by the rule:
$\operatorname{Val}(\mu)=\inf\{\tau(m(\cdot;\mu,\xi)):\ \ \xi\in\mathcal{U}\}.$
Notice that while this definition formally involves the relaxed continuity equation (4), the value function can be equivalently defined using usual controls and the dynamics (3), as shown in Remark 3.6 below.
3.1 Properties of the value function
The properties of the value function in a time-optimal problem in $\mathbb{R}^{d}$ are well studied (see [6, 41] for a comprehensive review). It is known that the value function is lower semicontinuous and satisfies the dynamic programming principle. In this section, we derive the analogous properties for the examined time-optimal control on the space of probability measures.
First, let us recall the stability of the trajectories on each finite interval proved in [4].
Proposition 3.5.
Assume that, for each $n$, $f_{n}$ is a function from $\mathbb{R}^{d}\times\mathcal{P}^{2}(\mathbb{R}^{d})\times U$ to $\mathbb{R}^{d}$; $\mu_{n}\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $\xi_{n}\in\mathcal{U}$ are such that
  • each function $f_{n}$ is continuous w.r.t. all variables and Lipschitz continuous w.r.t. $x$ and $m$ for the constant $L$;
  • there exists a continuous function $f:\mathbb{R}^{d}\times\mathcal{P}^{2}(\mathbb{R}^{d})\times U\rightarrow\mathbb{R}^{d}$ that is Lipschitz continuous w.r.t. $x$ and $m$ such that, for each $c>0$,
    $\begin{split}\sup\Big\{\|f_{n}(x,&m,u)-f(x,m,u)\|:\\ &x\in\mathbb{R}^{d},\,m\in\mathcal{P}^{2}(\mathbb{R}^{d}),\,\varsigma(m)\leq c,\,u\in U\Big\}\rightarrow 0\text{ as }n\rightarrow\infty;\end{split}$
  • the sequences $\{\mu_{n}\}_{n=1}^{\infty}$ and $\{\xi_{n}\}_{n=1}^{\infty}$ converge to $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$ and $\xi\in\mathcal{U}$ respectively;
  • for each natural $n$, $m_{n}(\cdot)$ is the solution of (4) for the dynamics equal to $f_{n}$, relaxed control $\xi_{n}$ such that $m_{n}(0)=\mu_{n}$.
Then, for every $T>0$, the sequence of $\{m_{n}(\cdot)\}_{n=1}^{\infty}$ converges to $m(\cdot;\mu,\xi)$ in $C([0,T];\mathcal{P}^{2}(\mathbb{R}^{d}))$.
Remark 3.6.
This proposition and the fact that the set of relaxed controls corresponding to usual ones is dense in $\mathcal{U}$ imply that, given $\xi\in\mathcal{U}$ and $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, there exists a sequence $\{m_{n}(\cdot)\}_{n=1}^{\infty}$ such that, for each $n$, $m_{n}(\cdot)$ satisfies
$\partial_{t}m_{n}(t)+\operatorname{div}\big(f(x,m_{n}(t),u_{n}(t))m_{n}(t)\big)=0,\ \ m_{n}(0)=\mu$
for some $u_{n}(\cdot)\in L^{0}([0,+\infty);U)$, whereas $m_{n}(\cdot)\rightarrow m(\cdot;\mu,\xi)$ as $n\rightarrow\infty$ uniformly on each time interval $[0,T]$. In particular,
$\begin{split}\operatorname{Val}(\mu)=\lim_{\varepsilon\downarrow 0}\inf\Big\{\tau^{\varepsilon}(m(\cdot)):\,m(\cdot)\text{ satisfies }\partial_{t}m(t)+\operatorname{div}(f(&x,m(t),u(t))m(t))=0,\\ m(0)=\mu&,\ \ u(\cdot)\in L^{0}([0,+\infty);U)\Big\},\end{split}$
where the mapping $\tau^{\varepsilon}(\cdot)$ assigns to a trajectory $m(\cdot)\in C([0,+\infty);\mathcal{P}^{2}(\mathbb{R}^{d}))$ the quantity $\inf\{t\in[0,+\infty):\,m(t)\in M+\mathbb{B}_{\varepsilon}\}$.
Now, we establish the existence of an optimal control whenever the value function is finite.
Theorem 3.7.
Assume that $\mu\in\operatorname{cl}G$ satisfies $\operatorname{Val}(\mu)<+\infty$. Then, there exists a relaxed control $\xi_{*}\in\mathcal{U}$ such that
$\operatorname{Val}(\mu)=\tau(m(\cdot;\mu,\xi_{*})).$
Remark 3.8.
We say that a relaxed control $\xi_{*}\in\mathcal{U}$ is optimal if
$\operatorname{Val}(\mu)=\tau(m(\cdot;\mu,\xi_{*})).$
Proof of Theorem 3.7.
By the definition of the value function, there exists a sequence $\{\xi_{n}\}_{n=1}^{\infty}\subset\mathcal{U}$ such that
$\operatorname{Val}(\mu)=\tau_{*}\triangleq\lim_{n\rightarrow\infty}\tau(m(\cdot;\mu,\xi_{n})).$
Furthermore, we denote $\tau_{n}\triangleq\tau(m(\cdot;\mu,\xi_{n})).$ Since $\mathcal{U}$ is compact, without loss of generality, one may assume that $\{\xi_{n}\}_{n=1}^{\infty}$ converges to some control $\xi_{*}\in\mathcal{U}$. Proposition 3.5 implies that $\{m(\cdot;\mu,\xi_{n})\}_{n=1}^{\infty}$ converges to $m(\cdot;\mu,\xi_{*})$. Since $m(\tau_{n};\mu,\xi_{n})\in M$, passing to the limit and using the closeness of the target set $M$, we conclude that
$m(\tau_{*};\mu,\xi_{*})\in M.$
Therefore, $\operatorname{Val}(\mu)=\tau(m(\cdot;\mu,\xi_{*}))$. ∎
Proposition 3.9.
The value function for problem (2), (5) is lower semicontinuous.
Доказательство.
Let $\{\mu_{k}\}_{k=1}^{\infty}\subset\operatorname{cl}G$ be a sequence of initial conditions for problem (2) converging to $\mu_{0}\in\operatorname{cl}G$. Denote $T_{k}\triangleq\operatorname{Val}(\mu_{k})$. First, we consider the case where $\{T_{k}\}_{k=1}^{\infty}$ is bounded. Without loss of generality, we assume that the sequence $\{T_{k}\}_{k=1}^{\infty}$ converges to some $T_{0}\in[0,+\infty)$. Furthermore, due to Theorem 3.7, for each $k$ there exists a relaxed control $\xi_{k}$ such that, $T_{k}=\tau(m(\cdot;\mu_{k},\xi_{k}))$. Again, without loss of generality, we assume that the whole sequence $\{\xi_{k}\}_{k=1}^{\infty}$ converges to $\xi_{0}\in\mathcal{U}$. Proposition 3.5 gives that the probabilities $m(T_{k};\mu_{k},\xi_{k})$ tend to $m(T_{0};\mu_{0},\xi_{0})$ as $k\rightarrow\infty$. Since the set $M$ is closed, while $m(T_{k};\mu_{k},\xi_{k})\in M$, we have that $m(T_{0};\mu_{0},\xi_{0})\in M$. Thus, using the choice of $T_{k}$, we conclude that
$\lim_{k\rightarrow\infty}\operatorname{Val}(\mu_{k})=T_{0}\geq\tau(m(\cdot;\mu_{0},\xi_{0}))\geq\operatorname{Val}(\mu_{0}).$
If
$\lim_{k\rightarrow\infty}T_{k}=+\infty,$
one obviously has that
$\lim_{k\rightarrow\infty}\operatorname{Val}(\mu_{k})\geq\operatorname{Val}(\mu_{0}).$
3.2 Dynamic programming principle
Theorem 3.10.
The value function for problem (2), (5) satisfies the dynamic programming principle on $\operatorname{cl}G$, i.e., for every $\mu\in\operatorname{cl}G$ and $h\in[0,\operatorname{Val}(\mu)]$,
$\operatorname{Val}(\mu)-h=\inf_{\xi\in\mathcal{U}_{h}}\operatorname{Val}(m(h;\mu,\xi)).$
Доказательство.
If $\xi\in\mathcal{U}_{h}$, $\eta\in\mathcal{U}$, then we denote by $\xi\diamond_{h}\eta$ the relaxed control determined by its disintegration as follows:
$(\xi\diamond_{h}\eta)(\cdot|t)\triangleq\begin{cases}\xi(\cdot|t),&t\in[0,h),\\ \eta(\cdot|t-h),&t\in[h,+\infty).\end{cases}$
We have that
$\begin{split}\inf\limits_{\xi\in\mathcal{U}_{h}}\operatorname{Val}(m(h;\mu,\xi))&=\inf\limits_{\xi\in\mathcal{U}_{h}}\inf\limits_{\eta\in\mathcal{U}}\tau(m(\cdot;m(h;\mu,\xi),\eta))\\ &=\inf\limits_{\xi\in\mathcal{U}_{h}}\inf\limits_{\eta\in\mathcal{U}}\tau(m(\cdot;\mu,\xi\diamond_{h}\eta))-h\\ &=\inf\limits_{\xi\in\mathcal{U}}\tau(m(\cdot;\mu,\xi))-h=\operatorname{Val}(\mu)-h.\end{split}$
4 The Dirichlet problem for the Hamilton-Jacobi equation
In this section, given $m\in\mathcal{P}^{2}(\mathbb{R}^{d})$ we use the set of functions satisfying sublinear growth condition $\operatorname{SL}(m)$ and its dual $\operatorname{LD}(m)$ introduced in Section 2.
We define the Hamiltonian by the following rule: for every $m\in\mathcal{P}^{2}(\mathbb{R}^{d})$ and $s\in\operatorname{LD}(m)$,
$H(m,s)=\inf\limits_{u\in U}\int_{\mathbb{R}^{d}}\left\langle s(x),f(x,m,u)\right\rangle m(dx).$
The value function of problem (2), (5) can have infinite value. So, we apply the Kruzhkov transform and define the function $\phi:\operatorname{cl}G\rightarrow\mathbb{R}$ by the rule:
$\phi(m)=1-e^{-\operatorname{Val}(m)}.$ (6)
Notice that $\phi$ takes values in $[0,1]$. Thus, we arrive at the following Dirichlet problem for the Hamilton-Jacobi equation and the transformed value function:
$H(m,\nabla\phi(m))+1-\phi(m)=0,\,m\in G;$ (7)
$\phi(m)=0,\,m\in\partial G.$ (8)
Here, we formally use $\nabla\phi(m)$ to designate the derivative of the function $\phi$. Below, we develop a viscosity solution approach based on nonsmooth analysis in the Wasserstein space, utilizing only sublinear variations.
Let $m\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $r>0$, and let $\psi$ be a function defined on $\mathbb{B}_{r}(m)$ with values in $\mathbb{R}\cup\{+\infty,-\infty\}$ such that $\psi(m)\in\mathbb{R}$.
Definition 4.1.
A subdifferential of the function $\psi$ at the measure $m$ consists of all functions $s\in\operatorname{LD}(m)$ such that, for every $F\in\operatorname{SL}(m)$,
$\liminf_{h\downarrow 0,F^{\prime}\in\operatorname{SL}(m),\,F^{\prime}\rightarrow F}\frac{\psi((\operatorname{Id}+hF^{\prime})\sharp m)-\psi(m)-h\langle s,F\rangle_{m}}{h}\geq 0.$ (9)
We denote the subdifferential of the function $\psi$ at the measure $m$ by $\partial^{-}\phi(m)$.
Similarly, the superdifferential of the $\psi$ at the measure $m$ denoted by $\partial^{+}\phi(m)$ contains all functions $s\in\operatorname{LD}(m)$ satisfying the following condition: given $F\in\operatorname{SL}(m)$,
$\limsup_{h\downarrow 0,F^{\prime}\in\operatorname{SL}(m),\,F^{\prime}\rightarrow F}\frac{\psi((\operatorname{Id}+hF^{\prime})\sharp m)-\psi(m)-h\langle s,F\rangle_{m}}{h}\leq 0.$ (10)
Remark 4.2.
An alternative definition of the sub-/superdifferentials involves the concept of Hadamard directional derivatives. Let a measure $m$, a radius $r$ and a function $\psi$ be as above. A Hadamard lower directional derivative of the function $\psi$ at the probability $m$ in a direction $F\in\operatorname{SL}(m)$ is
$d^{-}_{H}\psi(m;F)\triangleq\liminf\limits_{\begin{subarray}{c}h\downarrow 0\\ F^{\prime}\in\operatorname{SL}(m),F^{\prime}\rightarrow F\end{subarray}}\frac{\psi((\operatorname{Id}+hF^{\prime})\sharp m)-\psi(m)}{h}.$
Similarly, a Hadamard upper directional derivative of the function $\psi$ at a probability $m\in G$ in a direction $F\in\operatorname{SL}(m)$ is equal to
$d^{+}_{H}\psi(m;F)\triangleq\limsup\limits_{\begin{subarray}{c}h\downarrow 0,\\ F^{\prime}\in\operatorname{SL}(m),F^{\prime}\rightarrow F\end{subarray}}\frac{\psi((\operatorname{\operatorname{Id}}+hF^{\prime})\sharp m)-\psi(m)}{h}.$
One can easily show that
$\partial^{-}\psi(m)=\big\{s\in\operatorname{LD}(m):\,\left\langle s,F\right\rangle_{m}\leq d^{-}_{H}\psi(m;F)\big\},$
$\partial^{+}\psi(m)=\big\{s\in\operatorname{LD}(m):\,\left\langle s,F\right\rangle_{m}\geq d^{+}_{H}\psi(m;F)\big\}.$
Remark 4.3.
Notice that from the very definition of the subdifferential (see (9)) it directly follows that, if $\psi_{1},\psi_{2}:\mathbb{B}_{r}(m)\rightarrow\mathbb{R}\cup\{+\infty,-\infty\}$ for some positive number $r$ are such that $\psi_{1}(m)$ and $\psi_{2}(m)$ are finite, then
$\partial^{-}\psi_{1}(m)+\partial^{-}\psi_{2}(m)\subset\partial^{-}(\psi_{1}+\psi_{2})(m).$
Analogously, (10) implies the inclusion
$\partial^{+}\psi_{1}(m)+\partial^{+}\psi_{2}(m)\subset\partial^{+}(\psi_{1}+\psi_{2})(m).$
Remark 4.4.
As in the finite dimensional case, one has that, if $m\in\mathcal{P}^{2}(\mathbb{R}^{d})$ and $\psi:\mathbb{B}_{r}(m)\rightarrow\mathbb{R}\cup\{+\infty\}$ are such that $\psi(m)<\infty$ and $m$ is a local minimizer for the function $\psi$, then
$\mathbf{0}\in\partial^{-}\psi(m);$
hereinafter $\mathbf{0}$ stands for the function from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$ equal to zero everywhere.
Our approach to the definition of the viscosity solution to the Dirichlet problem is close to one proposed in [41, Definition 18.3, Theorem 18.6] and [6, §IV.3] for the finite dimensional case. In particular, we postulate that the viscosity solution coincides with the supersolution. Notice that in [6] this notion is called an envelope viscosity solution, while in [41] the words ‘minimax solution’ are used.
Definition 4.5.
A bounded lower semicontinuous function $\phi:\operatorname{cl}G\rightarrow\mathbb{R}$ is called a viscosity supersolution of Hamilton-Jacobi equation (7) provided that, for every $m\in G$ and $s\in\partial^{-}\phi(m)$,
$H(m,s)+1-\phi(m)\leq 0.$
We say that a bounded lower semicontinuous function $\phi:\operatorname{cl}G\rightarrow\mathbb{R}$ is called a viscosity supersolution of Dirichlet problem equation (7), (8) if it is a viscosity supersolution of (7) and satisfies (8).
Definition 4.6.
A bounded upper semicontinuous function $\phi:\operatorname{cl}G\rightarrow\mathbb{R}$ is called a viscosity subsolution of Hamilton-Jacobi equation (7) if, for every $m\in G$ and $s\in\partial^{+}\phi(m)$, one has that
$H(m,s)+1-\phi(m)\geq 0.$
We say that a bounded upper semicontinuous function $\phi:\operatorname{cl}G\rightarrow\mathbb{R}$ is called a viscosity subsolution of Dirichlet problem equation (7), (8) if it is a viscosity subsolution of (7) and satisfies (8).
Remark 4.7.
If $\phi$ is a nonnegative viscosity subsolution of (7), (8), then it is continuous at every point of $\partial G$.
Definition 4.8.
A bounded lower semicontinuos function $\phi:\operatorname{cl}G\rightarrow\mathbb{R}$ is called a viscosity solution of problem (7), (8) if
  • $\phi$ is a supersolution of problem (7), (8);
  • there exists a sequence of functions $\{\phi_{k}\}_{k=1}^{\infty}$ defined on $\operatorname{cl}G$ with values in $\mathbb{R}$ such that each function $\phi_{k}$ is a subsolution of (7), (8) while, for every $m\in G$, one has that
    $\phi(m)=\lim_{k\rightarrow\infty}\phi_{k}(m).$
Let us conclude this section with a discussion of the proposed definition of a viscosity solution. Our approach is based on sub- and superdifferentials. A key feature of our definition is that we restrict admissible variations to those satisfying a sublinear growth condition. Consequently, elements of the sub- and superdifferentials are permitted to grow faster than square-integrable functions. This differs from the frameworks in [2, Chapter 10] and [15, 5, 33, 34], where variations and elements of sub-/superdifferentials are required to be square-integrable (or, more generally, where variations belong to $L^{p}$ and sub-/superdifferentials to $L^{q}$).
Our motivation is twofold. First, in the proof of the comparison principle (Theorem 5.1), we rely heavily on a functional whose sub-/superdifferential contains an element that grows faster than an $L^{2}$ function. Second, when verifying that the Kruzhkov transform of the value function is a viscosity solution of (7), (8) (see Theorem 6.5), we use only variations of sublinear growth. Hence, if the controlled continuity equation satisfies the standard sublinear growth condition, the requirement of square-integrable variations is unnecessary.
The notion of viscosity solution given in Definition 4.8 appears stronger than those following [15, 5], where sub-/superdifferentials are defined via square-integrable variations. Whether one can define a viscosity solution to the Dirichlet problem (7), (8) while admitting general $L^{2}$ variations and still preserving the two main results (the comparison principle and the property that the Kruzhkov transform of the value function satisfies (7), (8)) remains an intriguing open question. Additionally, it would be interesting to develop an approach based on strong viscosity solutions in the spirit of [25, 33, 34].
5 Comparison principle
Theorem 5.1.
Let $\phi_{1}:\operatorname{cl}G\rightarrow\mathbb{R}$ be a subsolution of problem (7) and let $\phi_{2}:\operatorname{cl}G\rightarrow\mathbb{R}$ be a supersolution of problem (7). Assume that $\phi_{1}\leq\phi_{2}$ on $\partial G$. Then, for each $m\in\operatorname{cl}G$, $\phi_{1}(m)\leq\phi_{2}(m).$
Before we proceed with the proof of the comparison principle, let us formulate the uniqueness result.
Corollary 5.2.
There exists at most one solution of Dirichlet problem (7), (8).
This statement directly follows from Theorem 5.1 and Definition 4.8.
The proof of the comparison principle relies on the differentiability properties of the squared Wasserstein distance [2, Theorem 10.2.2] as well as the properties of a moment gauge function that will be used to regularize the difference $\phi_{1}-\phi_{2}$. We postpone the discussion of this function until Lemmas 5.3, 5.4 and now we recall [2, Theorem 10.2.2]. It states that, given $m^{\prime},m,m_{b}\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $\pi\in\Pi(m,m^{\prime})$, $\vartheta[m,m_{b}]\in\Pi(m,m_{b})$ that is an optimal plan between $m$ and $m_{b}$, and $\varpi\in\mathcal{P}((\mathbb{R}^{d})^{3})$ such that $\operatorname{p}^{1,2}\sharp\varpi=\pi$, while $\operatorname{p}^{1,3}\sharp\varpi=\vartheta$, one has that
$\begin{split}\frac{1}{2}W^{2}_{2}(m^{\prime},m_{b})-\frac{1}{2}W^{2}_{2}(m,m_{b})-\int_{(\mathbb{R}^{d})^{3}}\langle x-y,x^{\prime}-&x\rangle\varpi(d(x,x^{\prime},y))\\ &\leq\int_{(\mathbb{R}^{d})^{2}}\|x^{\prime}-x\|^{2}\pi(d(x,x^{\prime})).\end{split}$
Letting $m^{\prime}=(\operatorname{Id}+hF^{\prime})\sharp m$ in this inequality, where $F^{\prime}\in L^{2}(m)$, we obtain
$\frac{1}{2}W^{2}_{2}((\operatorname{Id}+hF^{\prime})\sharp m,m_{b})-\frac{1}{2}W^{2}_{2}(m,m_{b})-h\langle\hat{\vartheta}[m,m_{b}],F^{\prime}\rangle_{m}\leq h^{2}\|F^{\prime}\|^{2}_{L^{2}(m)}.$ (11)
Here, $\hat{\vartheta}[m,m_{b}]\in L^{2}(m)\subset\operatorname{LD}(m)$ is a barycenter of the plan $\vartheta$ defined by the rule:
$\hat{\vartheta}[m,m_{b}](x)\triangleq\int_{\mathbb{R}^{d}}(x-y)\vartheta[m,m_{b}](dy|x).$ (12)
Notice that
$\|\hat{\vartheta}[m,m_{b}]\|_{L^{2}(m)}=W_{2}(m,m_{b}).$ (13)
Since one can consider only $F^{\prime}\in\operatorname{SL}(m)$, while the convergence in $\operatorname{SL}(m)$ yields the convergence in $L^{2}(m)$, we have that $\hat{\vartheta}[m,m_{b}]$ is an element of the superdifferential of the function $m^{\prime}\mapsto\frac{1}{2}W_{2}^{2}(m^{\prime},m_{b})$ at $m$ in the sense of Definition 4.1.
Now let us discuss the moment gauge function, which plays a crucial role in the proof of the comparison principle. In what follows, the function $\mathscr{a}$ is defined by (3).
Lemma 5.3.
Let $m\in\mathcal{P}^{2}(\mathbb{R}^{d})$. Then, there exists a continuously differentiable function $g:(0,+\infty)\rightarrow(0,+\infty)$ such that
  1. (G1)
    $g$ is positive, monotonously increasing on $1,+\infty)$ with $g(r)\rightarrow+\infty$ as $r\rightarrow+\infty$;
  2. (G2)
    $\int_{\mathbb{R}^{d}}\mathscr{a}^{2}(x)g(\mathscr{a}(x))m(dx)<\infty$;
  3. (G3)
    $rg^{\prime}(r)\leq g(r)$, where $g^{\prime}$ stands for the derivative of the function $g$.
Lemma 5.4.
Let $m\in\mathcal{P}^{2}(\mathbb{R}^{d})$, and let $g:(0,+\infty)\rightarrow(0,+\infty)$ satisfy conditions (G1)(G3). Denote
$V_{g}(m)\triangleq\int_{\mathbb{R}^{d}}\mathscr{a}^{2}(x)g(\mathscr{a}(x))m(dx).$ (14)
Then, the function $\mathscr{j}_{g}(\cdot)$ defined by the rule:
$\mathscr{j}_{g}(x)\triangleq\big[2g(\mathscr{a}(x))+\mathscr{a}(x)g^{\prime}(\mathscr{a}(x))\big]x$
lies in the intersection of $\partial^{-}V_{g}(m)$ and $\partial^{+}V_{g}(m)$.
These auxiliary statements are proved in Appendix B.
Notice that the moment gauge function $g$ implies that the measure $m$ possesses a moment slightly higher than the second moment. Additionally, the sublevel sets of the potential energy function $V_{g}$ are compact (see [2, Lemma 5.1.7]). Furthermore, this functional is differentiable in the sense that $\mathscr{j}_{g}$ belongs to both the sub- and superdifferentials whenever $V_{g}$ is finite.
Proof of Theorem 5.1.
We modify the proof of [6, Theorem V.1.3]. As there, we will argue by contradiction. This means that there exist $\hat{m}\in G$ and $\varkappa>0$ such that
$\phi_{1}(\hat{m})-\phi_{2}(\hat{m})\geq 2\varkappa>0.$
Let $g$ be a function satisfying conditions (G1)(G3) for $m=\hat{m}$. For each $\varepsilon\in(0,1)$, we define the function $\Phi_{\varepsilon}:(\operatorname{cl}(G))^{2}\rightarrow\mathbb{R}$ as follows:
$\Phi_{\varepsilon}(m_{1},m_{2})\triangleq\phi_{1}(m_{1})-\phi_{2}(m_{2})-\frac{1}{2\varepsilon}W_{2}^{2}(m_{1},m_{2})-\beta\big[V_{g}(m_{1})\big]^{\kappa}-\beta\big[V_{g}(m_{2})\big]^{\kappa},$
where $V_{g}$ is defined by (14), $\kappa\triangleq\min\{1,1/(6C_{f})\}$, $C_{f}$ is a constant from Hypothesis 3.3, $\beta$ is such that
$\phi_{1}(\hat{m})-\phi_{2}(\hat{m})-2\beta\big[V_{g}(\hat{m})\big]^{\kappa}\geq\varkappa>0.$ (15)
This implies that, for every $\varepsilon>0$,
$\sup_{(m_{1},m_{2})\in(\operatorname{cl}G)^{2}}\Phi_{\varepsilon}(m_{1},m_{2})\geq\Phi_{\varepsilon}(\hat{m},\hat{m})\geq\varkappa.$
Notice that the function $\Phi_{\varepsilon}$ is u.s.c. Indeed, $\phi_{1}$ is u.s.c. by Definition 4.6, while $\phi_{2}$ is l.s.c. by Definition 4.5. Furthermore, the Wasserstein distance is continuous and the function $V_{g}$ is l.s.c. according to [2, Lemma 5.1.7]. Moreover, the function $\Phi_{\varepsilon}$ is bounded above.
Let $\hat{c}$ be a common bound for the values $|\phi_{1}(m)|$, $|\phi_{2}(m)|$, $m\in\operatorname{cl}(G)$, and let
$C_{1}\triangleq\big[2\hat{c}\beta^{-1}\big]^{1/\kappa}.$
Notice that, if $V_{g}(m_{1})\vee V_{g}(m_{2})>C_{1}$, $\varepsilon>0$, then
$\Phi_{\varepsilon}(m_{1},m_{2})<0.$
Set
$\mathcal{K}\triangleq\{m\in\operatorname{cl}(G):\,V_{g}(m)\leq C_{1}\}.$
Direct calculations show that the set $\mathcal{K}$ is tight and $2$-uniformly integrable. By [2, Proposition 7.1.5] and the lower semicontinuity of $V_{g}$, the set $\mathcal{K}$ is compact. Moreover, $(\hat{m},\hat{m})\in\mathcal{K}$.
Since $\Phi_{\varepsilon}$ is u.s.c. and $\mathcal{K}$ is compact, the function $\Phi_{\varepsilon}$ admits a maximizer $(m_{1}^{\varepsilon},m_{2}^{\varepsilon})$ on $\mathcal{K}^{2}$. Moreover, on $(\operatorname{cl}(G))^{2}\setminus(\mathcal{K})^{2}$, the function $\Phi_{\varepsilon}$ is negative, while at $\Phi_{\varepsilon}(\hat{m},\hat{m})\geq\varkappa$. Thus, the maximum of $\Phi_{\varepsilon}$ on $(\operatorname{cl}G)^{2}$ is attained at $(m_{1}^{\varepsilon},m_{2}^{\varepsilon})$. Moreover,
$\Phi_{\varepsilon}(m_{1}^{\varepsilon},m_{2}^{\varepsilon})\geq\Phi_{\varepsilon}(\hat{m},\hat{m})\geq\varkappa.$ (16)
From the latter inequality, it follows that $W_{2}(m_{1}^{\varepsilon},m_{2}^{\varepsilon})\rightarrow 0$ as $\varepsilon\rightarrow 0$. We wish to prove that
$\frac{W_{2}^{2}(m_{1}^{\varepsilon},m_{2}^{\varepsilon})}{2\varepsilon}\rightarrow 0\text{ as }\varepsilon\rightarrow 0.$ (17)
To this end, we let
$A\triangleq\sup_{m\in\mathcal{K}}\big[\phi_{1}(m)-\phi_{2}(m)-2\beta[V_{g}(m)]^{\kappa}\big].$
We use the fact that the maximum of $\Phi_{\varepsilon}$ is attained at $(m_{1}^{\varepsilon},m_{2}^{\varepsilon})$ and observe that
$\begin{split}A\leq\Phi_{\varepsilon}(m_{1}^{\varepsilon},m_{2}^{\varepsilon})\leq\phi_{1}(m_{1}^{\varepsilon})-\phi_{2}(m_{1}^{\varepsilon})-\beta\big[V_{g}(m_{1}^{\varepsilon})\big]^{\kappa}-\beta\big[V_{g}(m_{2}^{\varepsilon})\big]^{\kappa}.\end{split}$ (18)
Since $\phi_{1}$ is u.s.c., $\phi_{2}$ and $V_{g}$ are l.s.c., while $W_{2}(m_{1}^{\varepsilon},m_{2}^{\varepsilon})\rightarrow 0$ as $\varepsilon\rightarrow 0$, we have that
$\limsup_{\varepsilon\rightarrow 0}\Bigg[\phi_{1}(m_{1}^{\varepsilon})-\phi_{2}(m_{1}^{\varepsilon})-\beta\big[V_{g}(m_{1}^{\varepsilon})\big]^{\kappa}-\beta\big[V_{g}(m_{2}^{\varepsilon})\big]^{\kappa}\Bigg]\leq A.$
From this and (18), we obtain (17).
Now we consider two cases.
  1. (i)
    There exists a sequence $\{\varepsilon_{k}\}_{k=1}^{\infty}$ converging to zero such that $(m_{1}^{0},m_{2}^{0})=\lim_{k\rightarrow\infty}(m_{1}^{\varepsilon_{k}},m_{2}^{\varepsilon_{k}})\in\partial(G\times G)$;
  2. (ii)
    For every sequence $\{\varepsilon_{k}\}_{k=1}^{\infty}$ converging to zero such that $(m_{1}^{\varepsilon_{k}},m_{2}^{\varepsilon_{k}})\rightarrow(m_{1}^{0},m_{2}^{0})$ as $k\rightarrow\infty$ one has that $(m_{1}^{0},m_{2}^{0})\in\operatorname{int}(G\times G)$.
In case (i), we pass to the limit in inequality (16). Thanks to (17), we deduce that $m_{1}^{0}=m_{2}^{0}\in\partial G$, while
$\begin{split}\limsup_{k\rightarrow\infty}\Phi_{\varepsilon_{k}}(m_{1}^{\varepsilon_{k}},m_{2}^{\varepsilon_{k}})\leq\phi_{1}(m_{1}^{0})-\phi_{2}(m_{1}^{0}).\end{split}$
Here we used additionally the upper semicontinuity of the function $\phi_{1}$ and the lower semicontinuity of the function $\phi_{2}$. Thanks to the boundary conditions of these functions and (15), we obtain
$\varkappa\leq\limsup_{k\rightarrow\infty}\Phi_{\varepsilon_{k}}(m_{1}^{\varepsilon_{k}},m_{2}^{\varepsilon_{k}})\leq 0.$
This contradicts the choice of $\varkappa$.
In case (ii), we consider only those $\varepsilon$ for which $(m_{1}^{\varepsilon},m_{2}^{\varepsilon})\in\operatorname{int}(G\times G)$ and set
$\begin{split}\psi_{1,\varepsilon}(m_{1})&\triangleq\Phi_{\varepsilon}(m_{1},m_{2}^{\varepsilon})\\ &=\phi_{1}(m_{1})-\phi_{2}(m_{2}^{\varepsilon})-\frac{1}{2\varepsilon}W_{2}^{2}(m_{1},m_{2}^{\varepsilon})-\beta\big[V_{g}(m_{1})\big]^{\kappa}-\beta\big[V_{g}(m_{2}^{\varepsilon})\big]^{\kappa};\end{split}$
$\begin{split}\psi_{2,\varepsilon}(m_{2})&\triangleq-\Phi_{\varepsilon}(m_{1}^{\varepsilon},m_{2})\\ &=-\phi_{1}(m_{1}^{\varepsilon})+\phi_{2}(m_{2})+\frac{1}{2\varepsilon}W_{2}^{2}(m_{1}^{\varepsilon},m_{2})+\beta\big[V_{g}(m_{1}^{\varepsilon})\big]^{\kappa}+\beta\big[V_{g}(m_{2})\big]^{\kappa}.\end{split}$
Since the pair $(m_{1}^{\varepsilon},m_{2}^{\varepsilon})$ provides the maximum of $\Phi_{\varepsilon}$ on $(\operatorname{cl}(G))^{2}$, the function $\psi_{1,\varepsilon}$ attains a maximum at $m_{1}^{\varepsilon}$. Thus (see Remark 4.4), we have that $\mathbf{0}\in\partial^{+}\psi_{1,\varepsilon}(m_{1}^{\varepsilon})$. Let $\vartheta[m_{1}^{\varepsilon},m_{2}^{\varepsilon}]$ be an optimal plan between $m_{1}^{\varepsilon}$ and $m_{2}^{\varepsilon}$. Recall (see (12)) that
$\hat{\vartheta}[m_{1}^{\varepsilon},m_{2}^{\varepsilon}](x_{1})\triangleq\int_{\mathbb{R}^{d}}(x_{1}-x_{2})\vartheta[m_{1}^{\varepsilon},m_{2}^{\varepsilon}](dx_{2}|x_{1}).$
Thanks to (13), $\hat{\vartheta}[m_{1}^{\varepsilon},m_{2}^{\varepsilon}]$ lies in the superdifferential of the mapping $m_{1}\mapsto\frac{1}{2}W_{2}(m_{1},m_{2}^{\varepsilon})$ at $m_{1}^{\varepsilon}$.
Using this, Remark 4.3 and Lemma 5.4, we deduce that
$s_{1}^{\varepsilon}\triangleq\varepsilon^{-1}\hat{\vartheta}[m_{1}^{\varepsilon},m_{2}^{\varepsilon}]+\kappa\beta\big[V_{g}(m_{1}^{\varepsilon})\big]^{\kappa-1}\cdot j_{g}\in\partial^{+}\phi_{1}(m_{1}^{\varepsilon}).$
Since $\phi_{1}$ is a viscosity subsolution of (7), we have that
$H(m_{1}^{\varepsilon},s_{1}^{\varepsilon})+1-\phi_{1}(m_{1}^{\varepsilon})\geq 0.$ (19)
Analogously, if $\hat{\vartheta}[m_{2}^{\varepsilon},m_{1}^{\varepsilon}]$ is defined by the rule
$\hat{\vartheta}[m_{2}^{\varepsilon},m_{1}^{\varepsilon}](x_{2})\triangleq\int_{\mathbb{R}^{d}}(x_{2}-x_{1})\vartheta[m_{1}^{\varepsilon},m_{2}^{\varepsilon}](dx_{1}|x_{2}),$
then
$s_{2}^{\varepsilon}\triangleq-\varepsilon^{-1}\hat{\vartheta}[m_{2}^{\varepsilon},m_{1}^{\varepsilon}]-\kappa\beta\big[V_{g}(m_{2}^{\varepsilon})\big]^{\kappa-1}\cdot j_{g}\in\partial^{-}\phi_{2}(m_{2}^{\varepsilon}).$
Here once again used Remark 4.3, (11) and Lemma 5.4.
Recall that $\phi_{2}$ is a viscosity supersolution of (7). Hence,
$H(m_{2}^{\varepsilon},s_{2}^{\varepsilon})+1-\phi_{2}(m_{2}^{\varepsilon})\leq 0.$
This and (19) imply the following inequality
$\phi_{1}(m_{1}^{\varepsilon})-\phi_{2}(m_{2}^{\varepsilon})\leq H(m_{1}^{\varepsilon},s_{1}^{\varepsilon})-H(m_{2}^{\varepsilon},s_{2}^{\varepsilon}).$ (20)
Furthermore, let $u_{\varepsilon}$ be such that
$H(m_{2}^{\varepsilon},s_{2}^{\varepsilon})=\min_{u}\langle s_{2}^{\varepsilon},f(\cdot,m_{2}^{\varepsilon},u)\rangle_{m_{2}^{\varepsilon}}=\langle s_{2}^{\varepsilon},f(\cdot,m_{2}^{\varepsilon},u_{\varepsilon})\rangle_{m_{2}^{\varepsilon}}$
Therefore,
$\begin{split}H(m_{1}^{\varepsilon},s_{1}^{\varepsilon}&)-H(m_{2}^{\varepsilon},s_{2}^{\varepsilon})\\ \leq\varepsilon^{-1}\langle&\hat{\vartheta}[m_{1}^{\varepsilon},m_{2}^{\varepsilon}],f(\cdot,m_{1}^{\varepsilon},u_{\varepsilon})\rangle_{m_{1}^{\varepsilon}}-\varepsilon^{-1}\langle\hat{\vartheta}[m_{2}^{\varepsilon},m_{1}^{\varepsilon}],f(\cdot,m_{2}^{\varepsilon},u_{\varepsilon})\rangle_{m_{2}^{\varepsilon}}\\ +\kappa&\beta\big[V_{g}(m_{1}^{\varepsilon})\big]^{\kappa-1}\langle j_{g},f(\cdot,m_{1}^{\varepsilon},u_{\varepsilon})\rangle_{m_{1}^{\varepsilon}}-\kappa\beta\big[V_{g}(m_{2}^{\varepsilon})\big]^{\kappa-1}\langle j_{g},f(\cdot,m_{2}^{\varepsilon},u_{\varepsilon})\rangle_{m_{2}^{\varepsilon}}.\end{split}$ (21)
Now let us evaluate the right-hand side of this inequality. First, we use Hypothesis 3.2 and obtain
$\begin{split}\varepsilon^{-1}\langle&\hat{\vartheta}[m_{1}^{\varepsilon},m_{2}^{\varepsilon}],f(\cdot,m_{1}^{\varepsilon},u_{\varepsilon})\rangle_{m_{1}^{\varepsilon}}-\varepsilon^{-1}\langle\hat{\vartheta}[m_{2}^{\varepsilon},m_{1}^{\varepsilon}],f(\cdot,m_{2}^{\varepsilon},u_{\varepsilon})\rangle_{m_{2}^{\varepsilon}}\\ &=\varepsilon^{-1}\int_{(\mathbb{R}^{d})^{2}}\langle x_{1}-x_{2},f(x_{1},m_{1},u_{\varepsilon})-f(x_{2},m_{2},u_{\varepsilon})\rangle\vartheta[m_{1}^{\varepsilon},m_{2}^{\varepsilon}](d(x_{1},x_{2}))\\ &\leq 2L\varepsilon^{-1}W_{2}^{2}(m_{1},m_{2}).\end{split}$ (22)
Furthermore, using the definition of the function $j_{g}$, condition (G3) and Hypothesis 3.3, we conclude that
$\big|\langle j_{g},f(\cdot,m_{1}^{\varepsilon},u_{\varepsilon})\rangle_{m_{1}^{\varepsilon}}\big|\leq 3C_{f}\int_{\mathbb{R}^{d}}\mathscr{a}(x)g(\mathscr{a}(x))\Bigg[\mathscr{a}(x)+\int_{\mathbb{R}^{d}}\mathscr{a}(y)m_{1}^{\varepsilon}(dy)\Bigg]m_{1}^{\varepsilon}(dx).$
Inequality [30, (5)] and the monotonicity of the function $g$ yield that
$\big|\langle j_{g},f(\cdot,m_{1}^{\varepsilon},u_{\varepsilon})\rangle_{m_{1}^{\varepsilon}}\big|\leq 6C_{f}V_{g}(m_{1}^{\varepsilon}).$ (23)
Similarly,
$\big|\langle j_{g},f(\cdot,m_{2}^{\varepsilon},u_{\varepsilon})\rangle_{m_{2}^{\varepsilon}}\big|\leq 6C_{f}V_{g}(m_{2}^{\varepsilon}).$ (24)
Estimating the right-hand side of (21) via inequalities (22)–(24), we obtain that
$\begin{split}H(m_{1}^{\varepsilon}&,s_{1}^{\varepsilon})-H(m_{2}^{\varepsilon},s_{2}^{\varepsilon})\\ &\leq 2L\varepsilon^{-1}W_{2}^{2}(m_{1},m_{2})+6C_{f}\kappa\beta\big[V_{g}(m_{1}^{\varepsilon})]^{\kappa}+6C_{f}\kappa\beta\big[V_{g}(m_{2}^{\varepsilon})]^{\kappa}.\end{split}$
From this, (20) we obtain that
$\Phi_{\varepsilon}(m_{1}^{\varepsilon},m_{2}^{\varepsilon})\leq(2L+2^{-1})\varepsilon^{-1}W_{2}^{2}(m_{1}^{\varepsilon},m_{2}^{\varepsilon}).$
Here, we additionally used the choice of $\kappa=1\wedge(6C_{f})^{-1}$. Hence,
$\Phi(\hat{m},\hat{m})\leq(2L+2^{-1})\varepsilon^{-1}W_{2}^{2}(m_{1}^{\varepsilon},m_{2}^{\varepsilon}).$
Taking into account the fact that $\varepsilon^{-1}W_{2}^{2}(m_{1}^{\varepsilon},m_{2}^{\varepsilon})\rightarrow 0$ as $\varepsilon\rightarrow 0$ (see (17)), we conclude that $\Phi(\hat{m},\hat{m})\leq 0$, which contradicts (15). ∎
6 Characterization of the value function
We aim to show that the Kruzhkov transform of the value function (6) satisfies (7)–(8) in the viscosity sense. This result, in particular, ensures the existence of a viscosity solution for the system (7), (8).
Our proof will use the relaxations of the time-optimal problem in the space of measures. To define them, for every $\varepsilon>0$, we put
$M^{\varepsilon}\triangleq\operatorname{cl}\Big(\big\{m\in\mathcal{P}^{2}(\mathbb{R}^{d}):\,\operatorname{dist}(m,M)\leq\varepsilon\big\}\Big).$
Here,
$\operatorname{dist}(m,M)\triangleq\inf\big\{W_{2}(m,\mu):\,\mu\in M\big\}.$
Notice that
[…]
and
$M=\bigcap_{\varepsilon>0}M^{\varepsilon}.$
Now we consider the time-optimal problem for the target set $M^{\varepsilon}$. As above, given $m(\cdot)\in C([0,+\infty);\mathcal{P}^{2}(\mathbb{R}^{d}))$, we set
$\tau^{\varepsilon}(m(\cdot))\triangleq\inf\big\{t\in[0,+\infty):\,m(t)\in M^{\varepsilon}\big\}.$
By (25), we have that, if $\varepsilon_{1}\leq\varepsilon_{2}$,
$\tau^{\varepsilon_{1}}(m(\cdot))\geq\tau^{\varepsilon_{2}}(m(\cdot)).$ (26)
We define the value function for the time-optimal problem with the target set $M^{\varepsilon}$ by the rule:
$\operatorname{Val}^{\varepsilon}(\mu)\triangleq\inf\Big\{\tau^{\varepsilon}(m(\cdot;\mu,\xi)):\,\xi\in\mathcal{U}\Big\}.$
It is convenient to consider the function $\operatorname{Val}^{\varepsilon}$ on $G$. However, on $G\cap M^{\varepsilon}$ it is obviously equal to 0. As for the original value function we have the following properties.
Proposition 6.1.
The function $\operatorname{Val}^{\varepsilon}$ is lower semicontinuous. Moreover, $\operatorname{Val}^{\varepsilon}(\mu)=0$ whenever $\mu\in M^{\varepsilon}$.
Proposition 6.2.
If $\varepsilon>0$, $\mu\in G$ is such that $\operatorname{Val}^{\varepsilon}(\mu)<+\infty$. Then, there exists $\xi_{\varepsilon,\mu}$ satisfying
$\operatorname{Val}^{\varepsilon}(\mu)=\tau^{\varepsilon}(m(\cdot;\mu,\xi_{\varepsilon,\mu})).$
Proposition 6.3.
For every $\mu\in\operatorname{cl}G$ and $h\in[0,\operatorname{Val}^{\varepsilon}(\mu)]$,
$\operatorname{Val}^{\varepsilon}(\mu)-h=\inf_{\xi\in\mathcal{U}_{h}}\operatorname{Val}^{\varepsilon}(m(h;\mu,\xi)).$
Proposition 6.4.
If $\varepsilon_{1}\leq\varepsilon_{2}$, then
$\operatorname{Val}^{\varepsilon_{1}}(\mu)\geq\operatorname{Val}^{\varepsilon_{2}}(\mu).$
Доказательство.
If $\operatorname{Val}^{\varepsilon_{2}}(\mu)=+\infty$, there is nothing to prove. We will consider the case where $\operatorname{Val}^{\varepsilon_{2}}(\mu)$ is finite. Let $\xi_{\varepsilon_{1},\mu}\in\mathcal{U}$ be such that $\operatorname{Val}^{\varepsilon_{1}}(\mu)=\tau^{\varepsilon_{1}}(m(\cdot;\mu,\xi_{\varepsilon_{1},\mu})).$ By (26), we have that
$\tau^{\varepsilon_{1}}(m(\cdot;\mu,\xi_{\varepsilon_{1},\mu}))\geq\tau^{\varepsilon_{2}}(m(\cdot;\mu,\xi_{\varepsilon_{1},\mu})).$
This, together with the definition of the value function yield the statement of the proposition. ∎
Recall that above we introduced the function $\phi$ that is the Kruzhkov transform of the value function $\operatorname{Val}$ (see (6)).
Theorem 6.5.
The function $\phi$ is a viscosity solution to the Dirichlet problem (7), (8).
Доказательство.
We split the proof into four steps.
Step 1. Here, we prove that the function $\phi$ is a supersolution to (7), (8). First, notice that $\phi$ is bounded and lower semicontinuous, while $\phi(\mu)=0$ when $\mu\in\partial G$.
Furthermore, we choose a probability $\mu\in G$ and a function $s\in\partial^{-}\phi(\mu)$. By Remark 4.2, we have that, for every $F\in\operatorname{SL}(\mu)$,
$\left\langle s,F\right\rangle_{\mu}\leq d^{-}_{H}\phi(\mu;F).$ (27)
The dynamic programming principle (see Theorem 3.10) can be rewritten as follows: for each $h\in[0,-\ln(1-\phi(x)]$,
$(1-\phi(\mu))e^{h}=\sup_{\xi\in\mathcal{U}_{h}}(1-\phi(m(h;\mu,\xi))).$
Due to Theorem 3.7, there exists an optimal control $\xi_{h}\in\mathcal{U}_{h}$ such that
$(1-\phi(\mu))e^{h}=(1-\phi(m(h;\mu,\xi_{h}))).$
Let
$F_{h}(y)\triangleq\frac{1}{h}\int_{0}^{h}\int_{U}f(X(t;y,\mu,\xi_{h}),m(t;\mu,\xi_{h}),u)\xi_{h}(du|t)dt,$
where $X(\cdot;y,\mu,\xi_{h})$ is a trajectory such that
$\frac{d}{dt}X(t;y,\mu,\xi_{h})=\int_{U}f(X(t;y,\mu,\xi_{h}),m(t;\mu,\xi_{h}),u)\xi_{h}(du|t),\ \ X(0;y,\mu,\xi_{h})=y.$
Due to Proposition A.1 and Hypothesis 3.2, we have that $F_{h}\in\operatorname{SL}(\mu)$. Notice that $m(h;\mu,\xi_{h})=(\operatorname{Id}+hF_{h})\sharp\mu$. Thus, we have that
$\phi((\operatorname{Id}+hF_{h})\sharp\mu)-\phi(\mu)=(\phi(\mu)-1)\cdot(e^{h}-1).$ (28)
Furthermore, Proposition A.3 gives the existence of a sequence $\{h_{k}\}_{k=1}^{\infty}$ and a probability $\zeta\in\mathcal{P}(U)$ such that $h_{k}\rightarrow 0$, $F_{h_{k}}\rightarrow F$ as $k\rightarrow\infty$ for $F\in\operatorname{SL}(\mu)$ defined by the rule: $F(y)\triangleq\int_{U}f(y,\mu,u)\zeta(du)$. Therefore, from (28), we conclude that
$d^{-}_{H}\phi(\mu;F)\leq\phi(\mu)-1.$
This and (27) imply that
$\int_{U}\int_{\mathbb{R}^{d}}\left\langle s(x),f(x,\mu,u)\right\rangle\mu(dx)\zeta(du)\leq\phi(\mu)-1.$
Using the definition of the Hamiltonian, we obtain the following inequality, for every $\mu\in G$ and $s\in\partial^{-}\phi(\mu)$,
$H(\mu,s)+1-\phi(\mu)\leq 0.$
Thus, $\phi$ is a supersolution of problem (7), (8).
Step 2. Let $\psi^{\varepsilon}$ be the Kruzhkov transform of the function $\operatorname{Val}^{\varepsilon}$, i.e.,
$\psi^{\varepsilon}(\mu)\triangleq 1-e^{-\operatorname{Val}^{\varepsilon}(\mu)}.$
Moreover, we define $\phi^{\varepsilon}$ to be an upper envelope of the function $\psi^{\varepsilon}$. This means that, for every $\mu\in G$,
$\phi^{\varepsilon}(\mu)\triangleq\limsup_{\mu^{\prime}\rightarrow\mu,\ \ \mu^{\prime}\in\operatorname{cl}G}\psi^{\varepsilon}(\mu^{\prime})=\lim_{\delta\downarrow 0}\sup\big\{\psi^{\varepsilon}(\mu^{\prime}):\,W_{2}(\mu,\mu^{\prime})\leq\delta,\ \ \mu^{\prime}\in G\setminus M\big\}.$
Notice that, if $\mu\in G\cap M^{\varepsilon}$, then $\psi^{\varepsilon}(\mu)=0$. In particular, $\phi^{\varepsilon}(\mu)=0$ whenever $\mu\in\partial G$.
Step 3. Here we prove that each function $\phi^{\varepsilon}$ is a subsolution of (7), (8).
The fact that $\phi^{\varepsilon}$ satisfies the boundary condition was proved above. Moreover, the definition of $\phi^{\varepsilon}$ implies that it is upper semicontinuous and takes values in $[0,1]$.
Furthermore, let us show that the function $\phi^{\varepsilon}$ satisfies the following condition: for every $\mu\in G$ and $s\in\partial^{+}\phi^{\varepsilon}(\mu)$,
$H(\mu,s)+1-\phi^{\varepsilon}(\mu)\geq 0$ (29)
First, we consider the case when $\operatorname{Val}^{\varepsilon}(\mu)>0$.
By the construction of $\phi^{\varepsilon}$, there exists a sequence $\{\mu_{k}\}_{k=1}^{\infty}\subset G\setminus M$ converging to $\mu$, such that
$\lim_{k\rightarrow\infty}\psi^{\varepsilon}(\mu_{k})=\phi^{\varepsilon}(\mu).$
Without loss of generality, we assume that $\psi^{\varepsilon}(\mu_{k})\geq\hat{h}\triangleq\phi^{\varepsilon}(\mu)/2$. Since $\operatorname{Val}^{\varepsilon}(\mu)>0$, we have that $\hat{h}>0$. By the definition of the superdifferential, for every $F\in\operatorname{SL}(\mu)$, we have
$\left\langle s,F\right\rangle_{\mu}\geq d^{+}_{H}\phi^{\varepsilon}(\mu;F).$
We choose $f^{0}(x)\triangleq f(x,\mu,u_{0})$, where $u_{0}$ is such that
$H(\mu,s)=\left\langle s,f(\cdot,\mu,u_{0})\right\rangle_{\mu}.$ (30)
Hence,
$H(\mu,s)=\left\langle s,f^{0}\right\rangle_{\mu}\geq d^{+}_{H}\phi^{\varepsilon}(\mu;f^{0}).$ (31)
Furthermore, we evaluate $d^{+}_{H}\phi^{\varepsilon}(\mu;f^{0})$ in the following way.
Since $\psi^{\varepsilon}$ is the Kruzhkov transform of the function $\operatorname{Val}^{\varepsilon}$, the dynamic programming principle for problem with the target $M^{\varepsilon}$ (see Proposition 6.3) implies that, for $h\in[0,-\ln(1-\psi^{\varepsilon}(\mu_{k}))]$,
$(1-\psi^{\varepsilon}(\mu_{k}))e^{h}=\sup_{\xi\in\mathcal{U}_{h}}(1-\psi^{\varepsilon}(m(h;\mu_{k},\xi))).$ (32)
[…]
$(1-\psi^{\varepsilon}(\mu_{k}))e^{h}\geq 1-\psi^{\varepsilon}(m(h;\mu_{k},\xi^{0})).$
Using the definition of the function $\phi^{\varepsilon}$, the choice of the sequence $\{\mu_{k}\}_{k=1}^{\infty}$ and Proposition 3.5, we conclude that
$(1-\phi^{\varepsilon}(\mu))e^{h}\geq 1-\phi^{\varepsilon}(m(h;\mu,\xi^{0})).$
Now, as above, let $X(\cdot;y,\mu,\xi^{0})$ satisfy
$\frac{d}{dt}X(t;y,\mu,\xi^{0})=f(X(t;y,\mu,\xi^{0}),m(t;\mu,\xi^{0}),u^{0}),\ \ X(0;y,\mu,\xi^{0})=y.$
Set
$F_{h}(y)\triangleq\frac{1}{h}\int_{0}^{h}f(X(t;y,\mu,\xi^{0}),m(t),u^{0})dt.$
Proposition A.3 yields that $F_{h}\rightarrow f^{0}$ as $h\rightarrow 0$ in $\operatorname{SL}(\mu)$. Additionally, notice that $X(h;y,\mu,\xi^{0})=y+F_{h}(y)\cdot h.$ Thus,
$m(h;\mu,\xi^{0})=(\operatorname{Id}+hF_{h})\sharp\mu.$
Therefore,
$\phi^{\varepsilon}((\operatorname{Id}+hF_{h})\sharp\mu)-\phi^{\varepsilon}(\mu)\geq(\phi^{\varepsilon}(\mu)-1)\cdot(e^{h}-1).$
Dividing both parts on $h$ and using the definition of the Hadamard upper derivative, we conclude that
$d^{+}_{H}\phi^{\varepsilon}(\mu,f^{0})\geq\phi^{\varepsilon}(\mu)-1.$
This and (31) imply (29) for each $\mu\in G$ satisfying $\phi^{\varepsilon}(\mu)>0$ and every $s\in\partial^{+}\phi^{\varepsilon}(\mu)$.
Now, let $\phi^{\varepsilon}(\mu)=0$. As above, we choose $s\in\partial^{+}\phi^{\varepsilon}(\mu)$. Since $\phi^{\varepsilon}$ is nonnegative on $\operatorname{cl}(G)$, we have that $d^{+}_{H}\phi^{\varepsilon}(\mu;f^{0})\geq 0$. Here, $f^{0}(\cdot)=f(\cdot,\mu,u^{0})$ with $u^{0}$ satisfying (30). Thus (see (31)),
$H(\mu,s)\geq d^{+}_{H}\phi^{\varepsilon}(\mu;f^{0})\geq 0.$
This implies (29) for each $\mu$ satisfying $\phi^{\varepsilon}(\mu)=0$ and every $s\in\partial^{+}\phi^{\varepsilon}(\mu)$.
Step 4. We claim that, for each $\mu\in G$,
$\lim_{\varepsilon\downarrow 0}\phi^{\varepsilon}(\mu)=\phi(\mu).$ (33)
This equality directly follows from the inequality
$\lim_{\varepsilon\downarrow 0}\operatorname{Val}^{\varepsilon}(\mu)\geq\operatorname{Val}(\mu),$ (34)
the monotonicity of the Kruzhkov transform, Proposition 6.4 and the comparison principle (see Theorem 5.1). Equality (33) and Steps 1–3 give that the function $\phi$ is a viscosity solution of (7), (8).
So, it remains to prove (34). First, recall that Proposition 6.4 means that $\varepsilon\mapsto\operatorname{Val}^{\varepsilon}(\mu)$ decreases. Therefore, there exists a limit $\lim_{\varepsilon\downarrow 0}\operatorname{Val}^{\varepsilon}(\mu)$.
Now, we consider the case where
$T\triangleq\lim_{\varepsilon\downarrow 0}\operatorname{Val}^{\varepsilon}(\mu)<\infty.$ (35)
By Proposition 6.2, for each $\varepsilon>0$, there exists a relaxed control $\xi_{\varepsilon,\mu}$ such that
$\operatorname{Val}^{\varepsilon}(\mu)=\tau^{\varepsilon}(m(\cdot;\mu,\xi_{\varepsilon,\mu}))$
and $m(\operatorname{Val}^{\varepsilon}(\mu);\mu,\xi_{\varepsilon,\mu})\in M^{\varepsilon}$. Furthermore, the fact that $\mathcal{U}$ is compact implies that there exist a sequence $\{\varepsilon_{k}\}_{k=1}^{\infty}$ converging to zero and a relaxed control $\xi$ such that $\xi_{\varepsilon_{k},\mu}$ tend to $\xi$. Thus, by Proposition 3.5, we have that $\{m(\cdot;\mu,\xi_{\varepsilon_{k},\mu})\}_{k=1}^{\infty}$ converges to $m(\cdot;\mu,\xi)$. Using the definition of the sets $M^{\varepsilon}$, the functions $\operatorname{Val}^{\varepsilon}$ and the convergence (35), we conclude that $m(T;\mu,\xi)\in M.$ Hence,
$\operatorname{Val}(\mu)\leq T=\lim_{\varepsilon\downarrow 0}\operatorname{Val}^{\varepsilon}(\mu).$
The case where
$\lim_{\varepsilon\downarrow 0}\operatorname{Val}^{\varepsilon}(\mu)=\infty$
is obvious. ∎
7 Topological properties of the value function
First, we provide a sufficient condition for the continuity of the value function.
Theorem 7.1.
Assume that the viscosity solution of Dirichlet problem (7), (8) is continuous on $\partial G$. Then, it is continuous at every point of $\operatorname{cl}G$.
Remark 7.2.
Theorem 7.1 means that, if the value function is continuous on $\partial G$, then it is continuous everywhere on $\operatorname{cl}G$. Moreover, the continuity of the value function on $\partial G$ directly follows from the small time local attainability property introduced in [35]. For the examined case, it takes the following form: given $\varepsilon>0$, one can find $\delta>0$ such that $\operatorname{Val}(\mu)\leq\varepsilon$ whenever $\mu\in G$ satisfies $\operatorname{dist}(\mu,M)<\delta$. Obviously, under the small time local attainability assumption, the value function is continuous at every point of $\partial G$. If, additionally, $M$ is compact, the small time local attainability property is equivalent to the continuity of the value function on $\partial G$. For the space of probability measures, this property was considered in [17, 20]. In these studies, the regularity of the value function for the time-optimal problem in the space of probability measures was investigated under the assumption that the problem satisfies the small-time local attainability property, with the target being a hyperplane in the space of measures.
Proof of Theorem 7.1.
Let $\phi$ denote the aforementioned viscosity solution. For every $\mu\in\operatorname{cl}G$, we put
$\phi^{*}(\mu)\triangleq\limsup_{\mu^{\prime}\rightarrow\mu}\phi(\mu^{\prime}).$
Notice that the function $\phi^{*}:\operatorname{cl}G\rightarrow\mathbb{R}$ is upper semicontinuous. By following the same reasoning as in Step 3 of the proof of Theorem 6.5, one can show that
$H(\mu,s)+1-\phi^{*}(\mu)\geq 0$
for every $\mu\in G$ and $s\in\partial^{+}\phi^{*}(\mu)$. Moreover, the continuity of the function $\phi$ on $\partial G$ gives that $\phi^{*}(\mu)=0$ for every $\mu\in\partial G$. Thus, $\phi^{*}$ is a subsolution of (7), (8). The comparison principle (see Theorem 5.1) gives that $\phi^{*}(\mu)\leq\phi(\mu)$ on $\operatorname{cl}G$. Since the opposite inequality directly follows from the definition of the function $\phi^{*}$, we conclude that $\phi=\phi^{*}$ and the function $\phi$ is continuous. ∎
Now, let us consider the perturbation of the dynamics of the time-optimal problem. Since the value function is generally only lower semicontinuous, we require a suitable concept of convergence. In our opinion, the most natural convergence here is the $\Gamma$-convergence defined as follows (see [12, Definition 1.5]).
Definition 7.3.
For each natural $n$, let $F_{n}$ be a functional defined on a topological space $X$ with values in $\mathbb{R}\cup\{+\infty\}$. The sequence $\{F_{n}\}_{n=1}^{\infty}$ is said to $\Gamma$-converge to a functional $F:X\rightarrow\mathbb{R}\cup\{+\infty\}$ if the following two conditions hold:
  1. 1.
    for every sequence $\{x_{n}\}_{n=1}^{\infty}\in X$ satisfying $\lim\limits_{n\rightarrow\infty}x_{n}=x$,
    $F(x)\leq\liminf\limits_{n\rightarrow\infty}F_{n}(x_{n});$
  2. 2.
    for each $x\in X$, there is a sequence $\{x_{n}\}_{n=1}^{\infty}\in X$ such that
    $\lim\limits_{n\rightarrow\infty}x_{n}=x\text{ and }F(x)\geq\limsup\limits_{n\rightarrow\infty}F_{n}(x_{n}).$
If $\{F_{n}\}_{n=1}^{\infty}$ $\Gamma$-converges to $F$, then we write
$F=\Gamma\mbox{-}\lim\limits_{n\rightarrow\infty}F_{n}.$
Now we consider a perturbation of the dynamics of (2). For each natural $n$, let a function $f_{n}:\mathbb{R}^{d}\times\mathcal{P}^{2}(\mathbb{R}^{d})\times U\rightarrow\mathbb{R}^{d}$, $n\in\mathbb{N}$ satisfy Hypothesis 3.13.3. If $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $\xi\in\mathcal{U}$, then there exists a unique distributional solution to
$\partial_{t}m(t)+\operatorname{div}\Bigg(\int_{U}f_{n}(x,m(t),u)m(t)\xi(du|t)\Bigg)=0,\ \ m(0)=\mu.$
We denote this solution by $m_{n}(\cdot;\mu,\xi)$. The value function of the perturbed time-optimal is defined by the rule:
$\operatorname{Val}_{n}(\mu)\triangleq\inf\big\{\tau(m_{n}(\cdot;\mu,\xi):\,\xi\in\mathcal{U}\big\}.$
Theorem 7.4.
Assume the following convergence of the dynamics $f_{n}$ to $f$: for each $c>0$,
$\begin{split}\sup\Big\{\|f_{n}(x,&m,u)-f(x,m,u)\|:\\ &x\in\mathbb{R}^{d},\,m\in\mathcal{P}^{2}(\mathbb{R}^{d}),\,\varsigma(m)\leq c,\,u\in U\Big\}\rightarrow 0\text{ as }n\rightarrow\infty.\end{split}$
Then,
$\Gamma\mbox{-}\lim\limits_{n\rightarrow\infty}\operatorname{Val}_{n}=\operatorname{Val}.$
Доказательство.
Let us prove the first property in Definition 7.3. Let $\{\mu_{n}\}_{n=1}^{\infty}\in\mathcal{P}^{2}(\mathbb{R}^{d})$ be a sequence converging to $\mu$. If $\operatorname{Val}_{n}(\mu_{n})\rightarrow+\infty$ as $n\rightarrow\infty$, then, obviously,
$\operatorname{Val}(\mu)\leq\lim_{n\rightarrow\infty}\operatorname{Val}_{n}(\mu_{n}).$
If the limit is finite, without loss of generality, we can assume that the whole sequence $\{\operatorname{Val}_{n}(\mu_{n})\}_{n=1}^{\infty}$ converges to some number $\theta$. By Theorem 3.7, for each natural $n$, there exists a control $\xi_{n}\in\mathcal{U}$ such that
$\theta_{n}\triangleq\tau(m_{n}(\cdot;\mu_{n},\xi_{n}))=\operatorname{Val}_{n}(\mu_{n}).$
Additionally, for sufficiently large $n$ one has that $\tau(m_{n}(\cdot;\mu_{n},\xi_{n}))<2\theta$. Since $\mathcal{U}$ is compact, passing, if necessary, to a subsequence, we have that $\xi_{n}\rightarrow\xi$ as $n\rightarrow\infty$. Thus, Proposition 3.5 gives that $\{m_{n}(\cdot;\mu_{n},\xi_{n})\}_{n=1}^{\infty}$ converges to $m(\cdot;\mu,\xi)$ on $C([0,2\theta],\mathcal{P}^{2}(\mathbb{R}^{d}))$. Since, $m(\theta;\mu_{n},\xi_{n})\in M,$ whereas $M$ is closed, we have that $m(\theta;\mu,\xi)\in M$. Hence,
$\operatorname{Val}(\mu)\leq\tau(m(\cdot;\mu,\xi))\leq\theta=\liminf_{n\rightarrow\infty}\operatorname{Val}_{n}(\mu_{n}).$
This proves the first property of the $\Gamma$-convergence.
To prove the second property, we choose a measure $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$. First, we assume that $\operatorname{Val}(\mu)<\infty$. As above, by Theorem 3.7, one can find a control $\xi\in\mathcal{U}$, such that $\operatorname{Val}(\mu)=\tau(m(\cdot;\mu,\xi))$. For shortness, we put $\theta\triangleq\operatorname{Val}(\mu)$, $\nu\triangleq m(\theta,\mu,\xi)$. Notice that $\nu\in M$. Now, let $m^{\prime}_{n}(\cdot):[0,\theta]\rightarrow\mathcal{P}^{2}(\mathbb{R}^{d})$ satisfy
$\partial_{t}m_{n}^{\prime}(t)+\operatorname{div}\Bigg(\int_{U}f_{n}(\cdot,m_{n}^{\prime}(t),u)\xi(du|t)m_{n}^{\prime}(t)\Bigg)=0,\ \ m_{n}^{\prime}(\theta)=\nu.$
One can use time reversal to prove the existence of such a motion. Additionally, Proposition 3.5 applied in the reverse time gives that
$\sup_{t\in[0,\theta]}W_{2}(m_{n}^{\prime}(t),m(t,\mu,\xi))\rightarrow 0\text{ as }n\rightarrow\infty.$
We put $\mu_{n}\triangleq m_{n}^{\prime}(0)$. By construction, $m_{n}(\cdot,\mu_{n},\xi)=m_{n}^{\prime}(\cdot)$. Moreover, $W_{2}(\mu_{n},\mu)$ tends to 0 when $n\rightarrow\infty$. In particular, $\mu_{n}\in G$ for sufficiently large $n$. Thus, due to the fact that $m_{n}^{\prime}(\theta)=\nu\in M$, one has that $\tau(m_{n}(\cdot,\mu_{n},\xi))\leq\theta.$ Hence,
$\operatorname{Val}_{n}(\mu_{n})\leq\tau(m_{n}(\cdot;\mu_{n},\xi))\leq\tau(m(\cdot;\mu,\xi))=\operatorname{Val}(\mu).$
This means that $\limsup\limits_{n\rightarrow\infty}\operatorname{Val}_{n}(\mu_{n})\leq\operatorname{Val}(\mu).$
If $\operatorname{Val}(\mu)=+\infty$, then, for an arbitrary sequence $\{\mu_{n}\}_{n=1}^{\infty}\in G$ converging to $\mu\in G$, one has that $\limsup\limits_{n\rightarrow\infty}\operatorname{Val}_{n}(\mu_{n})\leq\operatorname{Val}(\mu).$
8 An example
We consider a time-optimal problem for the nonlocal continuity equation assuming that $d=1$, U=[1,0]U=[-1,0],
$f(x,m,u)=u-\int_{\mathbb{R}}ym(dy),$
while the target set is
$M=\Bigg\{m\in\mathcal{P}^{2}(\mathbb{R}):\ \int_{\mathbb{R}}ym(dy)=0\Bigg\}.$
Notice that this system satisfies Hypotheses 3.13.3.
In the examined case, nonlocal continuity equation (2) takes the form:
$\partial_{t}m(t)+\mbox{div}\Bigg(\bigg(u(t)-\int_{\mathbb{R}}ym(t,dy)\bigg)m(t)\Bigg)=0,\ m(0)=\mu\in\mathcal{P}^{2}(\mathbb{R}).$
The mean of the measure $m(t)$ defined by the rule:
$\bar{m}(t)\triangleq\int_{\mathbb{R}}ym(t,dy)$
obeys the following equation:
$\frac{d}{dt}\bar{m}(t)=u-\bar{m}(t),\ \ \bar{m}(0)=\bar{\mu}.$
Here, given a measure $\mu$, we put $\bar{\mu}\triangleq\int_{\mathbb{R}}y\mu(dy)$. Notice that the target set also is determined by the mean, i.e, the aim of the control is to find the first time $\tau$ such that $\bar{m}(\tau)=0$. Given an open-loop control $u(\cdot)$, we have that the corresponding evolution of the mean is
$\bar{m}(t)=\Bigg(\bar{\mu}+\int_{0}^{t}u(t^{\prime})e^{t^{\prime}}dt^{\prime}\Bigg)e^{-t}.$
Thus, the optimal control is as follows:
  1. 1.
    $u(\cdot)\equiv-1$ if $\bar{\mu}\geq 0$;
  2. 2.
    no optimal control if $\bar{\mu}<0$;
whereas the value function is
$\operatorname{Val}(\mu)=\left\{\begin{array}[]{ll}\ln(\bar{\mu}+1),&\bar{\mu}\geq 0,\\ +\infty,&\text{otherwise.}\end{array}\right$.
Notice that the value function is only lower semicontinuous.
Now, let us write down the Hamiltonian corresponding to the examined problem. It has the form:
$\begin{split}H(m,s)&=\inf\limits_{u\in[-1;0]}\int_{\mathbb{R}}s(x)\Bigl(u-\int_{\mathbb{R}}ym(dy)\Bigr)m(dx)\\ &=\begin{cases}\Bigl(-1-\bar{m}\Bigr)\int_{\mathbb{R}}s(x)m(dx),&\int_{\mathbb{R}}s(x)m(dx)>0,\\ -\bar{m}\int_{\mathbb{R}}s(x)m(dx),&\int_{\mathbb{R}}s(x)m(dx)\leq 0.\end{cases}\end{split}$
Thus, the Dirichlet problem is
$H(m,\nabla\phi(m))+1-\phi(m)=0,\ \ \int_{\mathbb{R}}ym(dy)\neq 0;\ \ \phi(m)=0\text{ when }\int_{\mathbb{R}}ym(dy)=0.$
Due to the Kruzhkov transform of the function $\operatorname{Val}$, the viscosity solution is equal to
$\phi(\mu)=\begin{cases}1-e^{-\ln(1+\bar{\mu})},&\bar{\mu}\geq 0,\\ 1,&\bar{\mu}<0.\end{cases}$
Acknowledgment. The article was prepared within the framework of the HSE University Basic Research Program. We sincerely thank the anonymous reviewer for their valuable comments and suggestions.
Приложение A Properties of relaxed controls
Recall that, if $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$,
$\varsigma(\mu)\triangleq\Bigg[\int_{\mathbb{R}^{d}}\|x\|^{2}\mu(dx)\Bigg]^{1/2}.$
Furthermore, for given $T>0$, $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, $\xi\in\mathcal{U}_{T}$ and $m(\cdot)\triangleq m(\cdot;\mu,\xi)$, we denote by $X(\cdot;y,\mu,\xi)$ a solution of the following initial value problem:
$\frac{d}{dt}x(t)=\int_{U}f(x(t),m(t),u)\xi(du|t),\ \ x(0)=y.$ (36)
Proposition A.1.
The following estimates hold true for every $t\in[0,T]$, $y\in\mathbb{R}^{d}$ and some constants dependent only on $\varsigma(\mu)$ and $T$:
  1. 1.
    $\varsigma(m(t))\leq c_{0}$;
  2. 2.
    $W_{2}(m(t),\mu)\leq c_{1}t$;
  3. 3.
    $\|X(t;y,\mu,\xi)\|\leq c_{3}(1+\|y\|)$;
  4. 4.
    $\|X(t;y,\mu,\xi)-y\|\leq c_{4}(1+\|y\|)t$.
Доказательство.
The proof relies on the integral representation of (36)
$X(\cdot;y,\mu,\xi)=y+\int_{0}^{t}\int_{U}f(X(t^{\prime};y,\mu,\xi),m(t^{\prime}),u)\xi(du|t^{\prime})dt^{\prime}$ (37)
and the upper bound on $f$ that directly follows from Hypothesis 3.2
$\|f(x,m,u)\|\leq C^{\prime}(1+\|x\|+\varsigma(m)).$ (38)
Here, $C^{\prime}$ is a constant.
Using this, along with the fact that $\varsigma(m(t))=\|X(t;\cdot,\mu,\xi)\|_{L^{2}(\mu)}$ and the Minkowski inequalities, we conclude that
$\varsigma(m(t))\leq\varsigma(\mu)+C^{\prime}\int_{0}^{t}(1+2\varsigma(m(t^{\prime}))dt^{\prime}.$
This and the Gronwall’s inequality imply the first statement of the proposition.
To show the second statement, it suffices to notice that $W_{2}(m(t),\mu)\leq\|X(t;\cdot,\mu,\xi)-\operatorname{Id}\|_{L^{2}(\mu)}$, use estimate (38) and the first statement.
The third and the fourth statements directly follow from integral representation of solution (37), estimate (38), the first statement of this proposition and the Gronwall’s inequality. ∎
Remark A.2.
One can use Hypothesis 3.3 and prove that
$\int_{\mathbb{R}^{d}}\mathscr{a}(x)m(t,dx)\leq c_{0}^{\prime}$
for some constant $c_{0}^{\prime}$. Here, $\mathscr{a}(x)=\sqrt{1+\|x\|^{2}}$ (as defined in (3)).
Proposition A.3.
Let
  • $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$;
  • $T>0$;
  • for each $h\in[0,T]$ $\xi_{h}\in\mathcal{U}_{h}$;
  • $F_{h}(y)\triangleq h^{-1}\int_{[0,h]\times U}f(X(t;y,\mu,\xi_{h}),m(t;\mu,\xi_{h}),u)\xi_{h}(d(t,u))$.
Then, there exists a sequence $\{h_{k}\}_{k=1}^{\infty}$ and a probability $\zeta\in\mathcal{P}(U)$ such that $h_{k}\rightarrow 0$, and $F_{h_{k}}\rightarrow F$ as $k\rightarrow\infty$, where $F\in\operatorname{SL}(\mu)$ is defined by the rule: $F(y)\triangleq\int_{U}f(y,\mu,u)\zeta(du)$. If, additionally, one can find a probability $\zeta_{0}\in\mathcal{P}(U)$ such that $\xi_{h}(\cdot|t)=\zeta_{0}$, then $\zeta=\zeta_{0}$ and $F_{h}\rightarrow F$ as $h\rightarrow 0$.
Доказательство.
First, notice that due to Proposition A.1 and Hypothesis 3.2, there exists a constant $C$ such that, for each $h\in[0,T]$ and $y\in\mathbb{R}^{d}$,
$\|F_{h}(y)\|\leq C(1+\|y\|)$
whenever $y\in\mathbb{R}^{d}$. Furthermore, we define the measure $\zeta_{h}\in\mathcal{P}(U)$ by the rule: for every Borel set $\Gamma\subset U$,
$\zeta_{h}(\Gamma)\triangleq h^{-1}\xi_{h}([0,h]\times\Gamma).$
Since $U$ is compact, there exists a sequence $\{h_{k}\}_{k=1}^{\infty}$ converging to zero such that $\{\zeta_{h_{k}}\}_{k=1}^{\infty}$ narrowly converges to some $\zeta\in\mathcal{P}(U)$. Due to Hypothesis 3.2,
$\begin{split}\|F_{h_{k}}(y)-F(y)\|\leq h_{k}^{-1}L\int_{0}^{h_{k}}\big(&\|X(t,y,\mu,\xi_{h_{h}})-y\|dt+W_{2}(m(t,\mu,\xi_{h_{k}}),\mu)\big)dt\\ &+\Bigg\|\int_{U}f(y,\mu,u)\zeta_{h_{k}}(du)-\int_{U}f(y,\mu,u)\zeta_{0}(du)\Bigg\|.\end{split}$
The first term tends to zero thanks to Proposition A.1. Since $\zeta_{h_{k}}$ converges to $\zeta$ narrowly, we have that the second term also vanishes. Hence, $F_{h_{k}}\rightarrow F$ in $\operatorname{SL}(\mu)$.
In the case where $\xi_{h}(\cdot|t)=\zeta_{0}$, we, as above, by Hypothesis 3.2, have that
$\|F_{h}(y)-F(y)\|\leq h^{-1}L\int_{0}^{h}\big(\|X(t,y,\mu,\xi_{h})-y\|+W_{2}(n(t,\mu,\xi_{h}),\mu)\big)dt.$
The desired convergence directly follows from this and Proposition A.3. ∎
Приложение B Moment gauge function
The main aim of this section is to prove Lemmas 5.3, 5.4.
Proof of Lemma 5.3.
First, notice that since $m$ is uniformly integrable, then
$\lim_{R\rightarrow\infty}\int_{\mathscr{a}(x)\geq R}\mathscr{a}^{2}(x)m(dx)\rightarrow 0.$
Now we construct an increasing sequence $\{R_{n}\}_{n=0}^{\infty}$ inductively. First, we put $R_{0}\triangleq 0$, $R_{1}\triangleq 1$. If we already constructed numbers $R_{1}<\ldots<R_{n}$, then we choose $R_{n+1}\geq(R_{n}+1)$ such that
$\int_{\mathscr{a}(x)\geq R_{n+1}}\mathscr{a}(x)m(dx)\leq(n+1)^{-1}2^{-n}$ (39)
and
$nR_{n+1}\geq(n+2)R_{n}.$
Notice that the latter inequality can be rewritten in the equivalent form
$\frac{1}{R_{n+1}-R_{n}}\leq\frac{n+2}{2R_{n+1}}.$ (40)
Now we define $g_{0}(r)$ by the rule:
  • if $r\in[R_{0},R_{1}]$, then $g(r)\triangleq 3$;
  • if $r\in[R_{n},R_{n+1}]$ for some $n\geq 1$, then we put
    $g_{0}(r)\triangleq(n+2)+(r-R_{n})(R_{n+1}-R_{n})^{-1}.$
Notice that the function $g_{0}$ is positive, monotonously increasing on $1,+\infty)$ and divergent. Since $g_{0}(r)\leq n+3$ when $r\in[R_{n},R_{n+1}]$, we have that
$\begin{split}\int_{\mathbb{R}^{d}}&\mathscr{a}^{2}(x)g_{0}(\mathscr{a}(x))m(dx)\\ &=\sum_{n=1}^{\infty}\int_{R_{n}\leq\mathscr{a}(x)<R_{n+1}}\mathscr{a}^{2}(x)g_{0}(\mathscr{a}(x))m(dx)\\ &\leq\sum_{n=1}^{\infty}\int_{R_{n}\leq\mathscr{a}(x)<R_{n+1}}\mathscr{a}^{2}(x)(n+3)m(dx)\\ &\leq 4\int_{\mathbb{R}^{d}}\mathscr{a}^{2}(x)m(dx)+\sum_{n=2}^{\infty}\int_{R_{n}\leq\mathscr{a}(x)}\mathscr{a}^{2}(x)nm(dx).\end{split}$ (41)
From this, the fact that $\int_{\mathbb{R}^{d}}\mathscr{a}^{2}(x)m(dx)<\infty$ and (39), we conclude that $\int_{\mathbb{R}^{d}}\mathscr{a}^{2}(x)g_{0}(\mathscr{a}(x))m(dx)$ is finite.
The function $g_{0}$ is differentiable on each interval $(R_{n},R_{n+1})$. The derivative on each interval $(R_{n},R_{n+1})$ is constant and equal to $A_{n}$, where $A_{0}=0$ and $A_{n}\triangleq(R_{n+1}-R_{n})^{-1}$ for $n\geq 1$. Notice that $A_{n}\in[0,1]$. From (40), we have that, if $r\in[R_{n},R_{n+1}]$, $n\geq 0$, then
$A_{n}\leq\frac{n+2}{2R_{n+1}}\leq\frac{g_{0}(r)}{2r}.$ (42)
Unfortunately, the function $g_{0}$ is only piecewise smooth. Thus, we use the following trick borrowed from the calculus of variation theory. Let
$\alpha(r)\triangleq\begin{cases}(1-|r|^{2})/4,&|r|\leq 1\\ 0,&\text{otherwise}\end{cases}$
Given natural $n$, let $\epsilon_{n}\in(0,1/2)$ be such that, for each $r\in[R_{n}-\epsilon_{n},R_{n}+\epsilon_{n}]$,
$A_{n}\vee A_{n-1}\leq 3g_{0}(r)/(4r).$
The existence of such $\epsilon_{n}$ follows from the fact that $A_{n}\vee A_{n-1}\leq g_{0}(R_{n})/2R_{n}$ (see (42)). Notice that $|A_{n}-A_{n-1}|\leq 1\leq g_{0}(r)$. Moreover, the intervals $R_{n}-\epsilon_{n},R_{n}+\epsilon_{n}]$ do not intersect.
Now we set
$g(r)\triangleq g_{0}(r)+\sum_{n=1}^{\infty}\epsilon_{n}(A_{n}-A_{n-1})\alpha\Bigg(\frac{r-R_{n}}{\epsilon_{n}}\Bigg).$
Notice that, $g(r)=g_{0}(r)$ outside the intervals $R_{n}-\epsilon_{n},R_{n}+\epsilon_{n}]$; whereas, on each interval $R_{n}-\epsilon_{n},R_{n}+\epsilon_{n}]$, $g(r)=g_{0}(r)+\epsilon_{n}(A_{n}-A_{n-1})\alpha((r-R_{n})/\epsilon_{n})$. Since each $A_{n}\in[0,1]$, $\epsilon_{n}\in(0,1/2)$, form (41) we have that
$\begin{split}\int_{\mathbb{R}^{d}}\mathscr{a}^{2}&(x)g(\mathscr{a}(x))m(dx)\\ &\leq\int_{\mathbb{R}^{d}}\mathscr{a}^{2}(x)g_{0}(\mathscr{a}(x))m(dx)+\frac{1}{8}\int_{\mathbb{R}^{d}}\mathscr{a}^{2}(x)m(dx)<\infty.\end{split}$
This proves (G2).
The very definition of the function $g$ gives that the function $g$ is continuously differentiable on $(0,+\infty)\setminus\cup_{n=1}^{\infty}[R_{n}-\epsilon_{n},R_{n}+\epsilon_{n}]$. Moreover, for $r\in(R_{n-1}+\epsilon_{n-1},R_{n}-\epsilon_{n})$, $g^{\prime}(r)=g_{0}^{\prime}(r)=A_{n-1}$. If $r\in[R_{n}-\epsilon_{n},R_{n}+\epsilon_{n}]$, then direct calculations gives that
$g^{\prime}(r)=(A_{n}+A_{n-1})/2+\frac{r-R_{n}}{2\epsilon_{n}}(A_{n}-A_{n-1}).$ (43)
This gives that $g^{\prime}(\cdot)$ is continuous on $0,+\infty)$ and positive on $R_{1}-\epsilon_{1},+\infty)$. The fact that $g(r)\rightarrow\infty$ as $r\rightarrow\infty$ follows from this, and definitions of the functions $g$, $g_{0}$. Thus, we obtain (G1).
It remains to show that condition (G3) holds true. When $r\in(0,+\infty)\setminus\cup_{n=1}^{\infty}[R_{n}-\epsilon_{n},R_{n}+\epsilon_{n}]$, we have (see (42)) that $g^{\prime}(r)\leq g(r)/(2r)\leq g(r)/r$. Notice that, on $R_{n}-\epsilon_{n},R_{n}+\epsilon_{n}]$, $g^{\prime}(r)\leq A_{n}\vee A_{n-1}$ (see (43)). Taking into account the inequality $|g(r)-g_{0}(r)|\leq 1/4$, we conclude that
$g^{\prime}(r)\leq A_{n}\vee A_{n-1}\leq\frac{3g_{0}(r)}{4r}\leq\frac{g(r)}{r}.$
This completes the proof. ∎
Lemma B.1.
If a function $g:(0,+\infty)\rightarrow(0,+\infty)$ satisfies conditions (G1) and (G3), then, for every $r\in(0,+\infty)$,
$g(2r)\leq(1-\ln 2)^{-1}g(r).$
Доказательство.
Due to (G3), we have that
$g(2r)-g(r)=\int_{0}^{1}g^{\prime}(r+hr)rdh\leq\int_{0}^{1}\frac{g(r+hr)}{1+h}dh.$
Condition (G1) gives that $g(r+hr)\leq g(2r)$. The statement of the lemma directly follows from this. ∎
Proof of Lemma 5.4.
For simplicity, we denote
$\varphi(x)=\mathscr{a}^{2}(x)g(\mathscr{a}(x)).$
Obviously,
$\nabla\varphi(x)=\mathscr{j}_{g}(x).$
Due to the definition of the function $\mathscr{a}$ (see (3)) and condition (G3), we have that $\mathscr{j}_{g}\in\operatorname{LD}(m)$. Notice that due to (G2), $V_{g}(m)$ is finite. Furthermore, if $C>0$, $F\in\operatorname{SL}(\mu)$ is such that $\|F(x)\|\leq C(1+\|x\|)$ $m$-a.e., we have that
$\mathscr{a}^{2}(x+hF(x))\leq 1+4C^{2}h^{2}+(2+4C^{2}h^{2})\|x\|^{2}\ \ m\text{-a.e.}$
Hence, for sufficiently small $h$, we have that, for $m$-a.e. $x\in\mathbb{R}^{d}$,
$\mathscr{a}(x+hF(x))\leq 4\mathscr{a}(x).$
This estimate is uniform over the class of functions $F$ satisfying the sublinear growth condition for a constant $C$.
From this and Lemma B.1, we have that $V_{g}((\operatorname{Id}+hF)\sharp m)$ is uniformly bounded whenever $h$ is sufficiently small.
Now let $C>0$, let $\{F_{n}\}_{n=1}^{\infty}\in\operatorname{SL}(m)$ be a sequence converging to some function $F\subset\operatorname{SL}(m)$ and let $h_{n}\in(0,C^{-1})$ with $h_{n}\rightarrow 0$. For each $x\in\mathbb{R}^{d}$, we have that
$\varphi(x+h_{n}F_{n}(x))-\varphi(x)=\int_{0}^{1}\langle\mathscr{j}_{g}(x+rh_{n}F_{n}(x)),h_{n}F_{n}(x)\rangle dr.$
Thanks to the fact that $\|\mathscr{j}_{g}(x)\|\leq 3g(\mathscr{a}(x))\|x\|$ and the choice of the sequence $\{F_{n}\}_{n=1}^{\infty}$, we have that the function $x\mapsto\int_{0}^{1}\langle\mathscr{j}_{g}(x+rh_{n}F_{n}(x)),F_{n}(x)\rangle dr$ is bounded by some function that is integrable w.r.t. the measure $m$. Moreover, for each $x$, $\int_{0}^{1}\langle\mathscr{j}_{g}(x+rh_{n}F_{n}(x)),F_{n}(x)\rangle dr\rightarrow\langle\mathscr{j}_{g}(x),F(x)\rangle$ as $n\rightarrow\infty$. Thus,
$\int_{\mathbb{R}^{d}}\int_{0}^{1}\langle\mathscr{j}_{g}(x+rh_{n}F_{n}(x)),F_{n}(x)\rangle drm(dx)\rightarrow\int_{\mathbb{R}^{d}}\langle\mathscr{j}_{g}(x),F(x)\rangle m(dx)\text{ as }n\rightarrow\infty.$
This gives that
$\lim_{h\downarrow 0,F^{\prime}\in\operatorname{SL}(m),F^{\prime}\rightarrow F}\frac{\varphi(x+hF_{n}(x))-\varphi(x)-\langle\mathscr{j}_{g},F\rangle_{m}}{h}=0.$
Using the definition of the sub- and superdifferentials, we obtain the statement of the lemma. ∎
Список литературы
  • [1] L. Ambrosio and W. Gangbo. Hamiltonian ODEs in the Wasserstein space of probability measures. Comm. Pure Appl. Math., 61(1):18–53, 2008.
  • [2] L. Ambrosio, N. Gigli, and G. Savaré. Gradient flows in metric spaces and in the space of probability measures. Lectures in Mathematics ETH Zürich. Birkhäuser, 2. ed edition, 2008.
  • [3] Y. Averboukh. Krasovskii–Subbotin approach to mean field type differential games. Dyn. Games Appl., 9:573–593, 2019.
  • [4] Y. Averboukh. Stability analysis of mean field type control system with major agent. COT, 2023:1–22, 2023.
  • [5] Z. Badreddine and H. Frankowska. Hamilton–JJacobi inequalities on a metric space. J. Differ. Equ., 271:1058–1091, 2021.
  • [6] M. Bardi and I. Capuzzo-Dolcetta. Optimal control and viscosity solutions of Hamilton–Jacobi–Bellman equations. Birkhäuser Boston, Boston, 1997.
  • [7] V. I. Bogachev. Measure theory. Vol. I and II. Berlin: Springer, 2007.
  • [8] M. Bongini, M. Fornasier, F. Rossi, and F. Solombrino. Mean-field Pontryagin maximum principle. J. Optim. Theory Appl., 175(1):1–38, 2017.
  • [9] B. Bonnet. A Pontryagin maximum principle in Wasserstein spaces for constrained optimal control problems. ESAIM: COCV, 25:38, 2019. Id/No 52.
  • [10] B. Bonnet and H. Frankowska. Necessary optimality conditions for optimal control problems in Wasserstein spaces. Appl. Math. Optim., 84:1281–1330, 2021.
  • [11] B. Bonnet and F. Rossi. The Pontryagin maximum principle in the Wasserstein space. Calc. Var. Partial Differ. Equ., 58(1):36, 2019. Id/No 11.
  • [12] A. Braides. Gamma-convergence for beginners. Oxford University Press, 2002.
  • [13] R. W. Brockett. Notes on the control of the Liouville equation. In Control of Partial Differential Equations, pages 101–129, 2012.
  • [14] P. Cardaliaguet, F. Delarue, J.-M. Lasry, and P.-L. Lions. The master Equation and the convergence problem in mean field games. Princeton University Press, 2019.
  • [15] P. Cardaliaguet and M. Quincampoix. Deterministic differential games under probability knowledge of initial condition. Int. Game Theory Rev., 10(01):1–16, 2008.
  • [16] J. A. Carrillo, M. Fornasier, G. Toscani, and F. Vecil. Particle, kinetic, and hydrodynamic models of swarming. In Mathematical modeling of collective behavior in socio-economic and life sciences, pages 297–336. Boston, MA: Birkhäuser, 2010.
  • [17] G. Cavagnari. Regularity results for a time-optimal control problem in the space of probability measures. Math. Control Relat. Fields, 7(2):213–233, 2017.
  • [18] G. Cavagnari and A. Marigonda. Time-optimal control problem in the space of probability measures. In I. Lirkov, S. Margenov, and J. Waśniewski, editors, Large-Scale Scientific Computing (LSSC 2015), volume 9374 of Lecture Notes in Computer Science, pages 94–102. Springer, Cham, 2015.
  • [19] G. Cavagnari and A. Marigonda. Attainability property for a probabilistic target in Wasserstein spaces. Discrete Continuous Dyn. Syst. - A, 41(2):777–812, 2021.
  • [20] G. Cavagnari, A. Marigonda, and M. Quincampoix. Compatibility of state constraints and dynamics for multiagent control systems. J. Evol. Equ., 21(4):4491–4537, 2021.
  • [21] R. M. Colombo, M. Garavello, M. Lécureux-Mercier, and N. Pogodaev. Conservation laws in the modeling of moving crowds. Preprint, arXiv:1211.0408 [math.AP] (2012), 2012.
  • [22] R. M. Colombo, M. Herty, and M. Mercier. Control of the continuity equation with a non local flow. ESAIM Control Optim. Calc. Var., 17(2):353–379, 2011.
  • [23] A. Cosso, F. Gozzi, I. Kharroubi, H. Pham, and M. Rosestolato. Master Bellman equation in the Wasserstein space: Uniqueness of viscosity solutions. Trans. Am. Math. Soc., 377(01):31–83, 2024.
  • [24] M. G. Crandall and P.-L. Lions. Viscosity solutions of Hamilton–Jacobi equations. Trans. Amer. Math. Soc., 277(1):1, 1983.
  • [25] M. G. Crandall and P.-L. Lions. Hamilton–JJacobi equations in infinite dimensions i. uniqueness of viscosity solutions. J. Funct. Anal., 62(3):379–396, 1985.
  • [26] S. Daudin and B. Seeger. A comparison principle for semilinear Hamilton–Jacobi–Bellman equations in the Wasserstein space. Calc. Var. Partial Differ. Equ., 63(3):106, 2024.
  • [27] C. Dogbe. On the modelling of crowd dynamics by generalized kinetic models. J. Math. Anal. Appl., 387(2):512–532, 2012.
  • [28] R. C. Fetecau, Y. Huang, and T. Kolokolnikov. Swarm dynamics and equilibria for a nonlocal aggregation model. Nonlinearity, 24(10):2681–2716, 2011.
  • [29] W. Gangbo and A. Tudorascu. On differentiability in the Wasserstein space and well-posedness for Hamilton-Jacobi equations. J. Math. Pures Appl., 125:119–174, 2019.
  • [30] J. Gurland. The teacher’s corner: An inequality satisfied by the expectation of the reciprocal of a random variable. Am. Stat., 21(2):24–25, 1967.
  • [31] R. L. Hughes. A continuum theory for the flow of pedestrians. Transp. Res. Part B, 36(6):507–535, 2002.
  • [32] C. Jimenez. Equivalence between Strict Viscosity Solution and Viscosity Solution in the Wasserstein Space and Regular Extension of the Hamiltonian in $L^{2}_{\mathbb{P}}$. J. Convex Anal., 31(2):619–670, 2024.
  • [33] C. Jimenez, A. Marigonda, and M. Quincampoix. Optimal control of multiagent systems in the Wasserstein space. Calc. Var. Partial Differ. Equ., 59(2):45, 2020. Id/No 58.
  • [34] C. Jimenez, A. Marigonda, and M. Quincampoix. Dynamical systems and Hamilton–Jacobi–Bellman equations on the Wasserstein space and their ${L}^{2}$ representations. SIAM J. Math. Anal., 55(5):5919–5966, 2023.
  • [35] M. Krastanov and M. Quincampoix. Local small time controllability and attainability of a set for nonlinear control system. ESAIM Control Optim. Calc. Var., 6:499–516, 2001.
  • [36] A. Mogilner and L. Edelstein-Keshet. A non-local model for a swarm. J. Math. Biol., 38(6):534–570, 1999.
  • [37] G. Naldi, L. Pareschi, and G. Toscani, editors. Mathematical modeling of collective behavior in socio-economic and life sciences. Model. Simul. Sci. Eng. Technol. Boston, MA: Birkhäuser, 2010.
  • [38] H. Pham and X. Wei. Bellman equation and viscosity solutions for mean-field stochastic control problem. ESAIM Control Optim. Calc. Var., 24(1):437–461, 2018.
  • [39] B. Piccoli and F. Rossi. Measure-theoretic models for crowd dynamics. In Crowd dynamics, Volume 1. Theory, models, and safety problems, pages 137–165. Cham: Birkhäuser, 2018.
  • [40] N. Pogodaev. Optimal control of continuity equations. Nonlinear Differ. Equ. Appl., 23(2):24, 2016. Id/No 21.
  • [41] A. I. Subbotin. Generalized solutions of first-order PDEs: The dynamical optimization perspective. Birkhäuser Boston, Boston, 1995.
  • [42] B.-C. Wang and Y. Liang. Robust mean field social control problems with applications in analysis of opinion dynamics. Int. J. Control, 95(12):3309–3325, 2022.
  • [43] J. Warga. Optimal Control of Differential and Functional Equations. Elsevier Inc, Academic Press Inc, 1972.